### v2

# CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk

Abstract
Accurate uncertainty quantification is critical for reliable predictive modeling. Existing methods typically address either aleatoric uncertainty due to measurement noise or epistemic uncertainty resulting from limited data, but not both in a balanced manner. We propose CLEAR, a calibration method with two distinct parameters, $\gamma_{1}$ and $\gamma_{2}$, to combine the two uncertainty components and improve the conditional coverage of predictive intervals for regression tasks. CLEAR is compatible with any pair of aleatoric and epistemic estimators; we show how it can be used with (i) quantile regression for aleatoric uncertainty and (ii) ensembles drawn from the Predictability–Computability–Stability (PCS) framework for epistemic uncertainty. Across 17 diverse real-world datasets, CLEAR achieves an average improvement of 28.2% and 17.4% in the interval width compared to the two individually calibrated baselines while maintaining nominal coverage. Similar improvements are observed when applying CLEAR to Deep Ensembles (epistemic) and Simultaneous Quantile Regression (aleatoric). The benefits are especially evident in scenarios dominated by high aleatoric or epistemic uncertainty. [^1] Code: https://github.com/Unco3892/clear
1 Introduction
Uncertainty quantification (UQ) is essential for building reliable machine learning systems [1, 2]. Despite their impressive capabilities, modern machine learning methods can give a false sense of reliability; therefore, producing valid prediction intervals remains an open problem. Calibration [3] and conformal methods [4, 5, 6, 7, 8] adjust prediction intervals to obtain marginal coverage (that is, covering a certain percentage of the data on average). However, they may suffer from poor conditional coverage, meaning well-calibrated coverage at the individual or subgroup level [9]. In particular, under distribution shift or model misspecification, conditional coverage can degrade substantially, especially in extrapolation regions. Most conformal methods, such as conformalized quantile regression (CQR) [10], only capture aleatoric uncertainty while ignoring epistemic uncertainty.
Figure 1: Left: Blue represents aleatoric uncertainty, which reflects randomness inherent in the data such as measurement noise. Red represents epistemic uncertainty, which arises from limited sample size. Right: Estimated prediction sets using the CLEAR method, which combines both sources of uncertainty in a data-driven manner.
It is important to distinguish between the two main sources of uncertainty, namely epistemic and aleatoric. Epistemic uncertainty [11] arises from our limited understanding of the data generation process and the model, encompassing issues related to data collection, preprocessing, transformation, and model specification. Notably, this uncertainty is typically large in extrapolation regions where training data is sparse. In contrast, aleatoric uncertainty [12] reflects the inherent variability within the data (stemming from measurement errors, missing covariates, randomness, or intrinsic noise) that cannot be reduced simply by gathering more observations or refining the model unless the data acquisition process itself is improved where more features are measured. Separating the two sources can be beneficial for various applications [13, 14]. For instance, in active learning, epistemic uncertainty helps in selecting which samples to label, while aleatoric uncertainty is less relevant [15]. However, for prediction tasks, appropriately combining both epistemic and aleatoric parts is necessary to account for the overall uncertainty in the model’s predictions [16].
In this work, we contribute to the existing literature by combining aleatoric and epistemic uncertainties in a data-driven manner. We consider prediction intervals $C(x)$ of the form
$\begin{split}C(x)=\left[\hat{f}(x)\pm\big(\gamma_{1}\times\text{aleatoric}(x)+\gamma_{2}\times\text{epistemic}(x)\big)\right],\quad\text{with }\gamma_{2}=\lambda\gamma_{1},\end{split}$ (1)
where $\hat{f}$ is a point estimate, and $\gamma_{1},\gamma_{2}\in[0,\infty)$ are coefficients selected to 1) calibrate marginal coverage and 2) optimally balance the two types of uncertainty (optimality is defined in terms of a quantile loss metric introduced later). The parameter $\lambda$ controls the trade-off between the two uncertainty types: when $\lambda=0$, the interval reflects only aleatoric uncertainty, while as $\gamma_{1}\to 0,\lambda=\frac{\gamma_{2}}{\gamma_{1}}\to\infty$, it reflects only epistemic uncertainty. By allowing an adaptively chosen $\lambda$, we ensure a more flexible and data-driven trade-off to the two components, leading to prediction intervals that are both well-calibrated and more informative (Figure 1). Moreover, estimating $\lambda$ can help practitioners better understand which source of uncertainty is the dominant contributor to the overall uncertainty.
The combination of epistemic and aleatoric uncertainty is not new: Bayesian methods have long incorporated both components [17, 18], and several recent approaches have also explored this decomposition in the context of conformal prediction [19, 20, 21]. However, to the best of our knowledge, existing methods either do not explicitly distinguish between the two types of uncertainty or implicitly fix the combination ratio, for instance, setting $\lambda=1$ [22, 17, 18], or fixing $\gamma_{1}=1$ [19]. This fixed choice may be suboptimal, as the relative importance of each uncertainty type varies with the data distribution and prediction task (see Appendix A for more details on related work).
1.1 Contributions
  1. 1.
    We are the first to introduce two calibration parameters $\gamma_{1}$ and $\gamma_{2}$, to balance the scales of epistemic and aleatoric uncertainty on the validation dataset.
  2. 2.
    We demonstrate that fitting quantiles on the residuals provides much more sensible estimators of aleatoric uncertainty than fitting the quantiles directly on the targets.
  3. 3.
    We are the first to combine the ensemble perturbation intervals of the PCS framework [23] with the CQR aleatoric uncertainty estimator, and empirically show the strengths of this combination.
  4. 4.
    We conduct large-scale UQ benchmarking for several models on 17 regression datasets.
2 Method
2.1 Problem scenario
Consider a classical setting, where an i.i.d. sample $(X_{i},Y_{i}),i=1,\dots,n$ is drawn from distribution $P_{X}\times P_{Y\mid X}$. The goal of conformal inference is to construct a prediction set $C(X_{n+1})\subseteq supp(Y)$ for a new data-point $(X_{n+1},Y_{n+1})$ satisfying marginal coverage
$\mathbb{P}\big(Y_{n+1}\in C(X_{n+1})\big)\geq 1-\alpha,$ (2)
where $\alpha\in(0,1)$ is for instance $\alpha=0.05$. In order to construct $C$, data $\mathcal{D}=\{(X_{i},Y_{i}),i=1,\dots,n\}$ can be split into train and calibration subsets $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{cal}}$. On the training data, a first estimate of $C$ can be constructed, and then we can use data from $\mathcal{D}_{\text{cal}}$ to calibrate $C$ such that (2) is satisfied.
In case of CQR, we first estimate conditional quantiles $\hat{q}_{\alpha/2}(x),\hat{q}_{1-\alpha/2}(x)$ using $\mathcal{D}_{\text{train}}$, and then construct $C(x)=[\hat{q}_{\alpha/2}(x)-\gamma,\hat{q}_{1-\alpha/2}(x)+\gamma]$, where the calibration parameter $\gamma$ is chosen so that the prediction interval $C(X_{i})$ contains $Y_{i}$ for exactly $\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\rceil$ points in the calibration set $\mathcal{D}_{\text{cal}}$. While this procedure guarantees finite-sample distribution-free marginal coverage [8], conditional coverage
$\mathbb{P}\big(Y_{n+1}\in C(X_{n+1})\mid X_{n+1}=x\big)\geq 1-\alpha$
does not need to hold. As pointed out in [24, 25], any algorithm with finite-sample distribution-free conditional coverage guarantees for all $x$ must be trivial $C(x)=(-\infty,\infty)$. However, we aim to design estimators such that conditional coverage holds approximately under reasonable real-world scenarios, even if exact finite-sample guarantees are impossible in general.
2.2 Epistemic uncertainty
The traditional machine learning approach trains a predictive algorithm on a single version of the cleaned/preprocessed dataset and uses the best-performing algorithm (compared using the validation set) for future predictions. While theoretically sound in the infinite-sample limit, this approach ignores the uncertainty stemming from finite sample size and model choice (epistemic uncertainty). Various methods have been proposed to estimate this uncertainty, including Deep Ensembles [22], MC dropout in NN [26], Orthonormal Certificates [13], NOMU [27], BNNs [28] and Laplace Approximation [29], among others.
Estimating Epistemic Uncertainty via PCS: In practice, additional sources of uncertainty arise from subjective choices in data cleaning, imputation, and dataset construction, which we also consider as extended epistemic uncertainty. The Predictability, Computability, and Stability (PCS) framework [23] offers a holistic point of view on the data-science-life-cycle, without explicitly modeling aleatoric uncertainty. One can obtain an ensemble of $m$ estimators $\hat{f}_{1},\dots,\hat{f}_{m}$ as follows:
  1. 1.
    Split the data into $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}}$.
  2. 2.
    Create $N_{1}$ differently preprocessed versions of the data and define $N_{2}$ different models (e.g., linear regression, random forest, neural networks). Then, train models for all $N_{1}\times N_{2}$ combinations on $\mathcal{D}_{\text{train}}$ and pick the top-$k$ based on their performance on $\mathcal{D}_{\text{val}}$.
  3. 3.
    Refit each of the top-$k$ models on $b$ bootstrap samples $\mathcal{D}_{\text{train}}^{1},\dots,\mathcal{D}_{\text{train}}^{b}$ of $\mathcal{D}_{\text{train}}$ to obtain an ensemble of $m=k\times b$ estimators $\hat{f}_{1},\dots,\hat{f}_{m}$.
Taking the point-wise median of $\hat{f}_{1},\dots,\hat{f}_{m}$, yields the final PCS estimate $\hat{f}$, and the point-wise $\alpha/2$ and $1-\alpha/2$ quantiles (denoted $\hat{f}_{\alpha/2}$ and $\hat{f}_{1-\alpha/2}$, respectively) define the (uncalibrated) uncertainty band. The widths of this interval, denoted as $\hat{q}^{\text{epi}}_{1-\alpha/2}(x):=\hat{f}_{1-\alpha/2}(x)-\hat{f}(x)$ and $\hat{q}^{\text{epi}}_{\alpha/2}(x):=\hat{f}(x)-\hat{f}_{\alpha/2}(x)$, quantifies the pre-calibrated epistemic uncertainty. [30] extends this by using the combined data set $\mathcal{D}_{\text{train}}\cup\mathcal{D}_{\text{val}}$ for training and calibrating uncertainty based on out-of-bag data. The focus is only on the modeling step of the framework. We later demonstrate—through a case study in Section 4.3—that accounting for uncertainty from data cleaning and pre-processing is also important, and how the uncertainty contributions can be adapted and improved using our approach.
2.3 Aleatoric uncertainty
Aleatoric uncertainty can be estimated by modeling the conditional distribution of the outcome given the inputs. Common approaches include direct conditional quantile regression [31], either parametric or nonparametric, such as smooth quantile regression [32] and quantile random forests (QRF) [33]; heteroskedastic models that estimate input-dependent noise levels $\sigma(x)$ under Gaussian assumptions [34]; and distributional regression techniques such as simultaneous quantile regression [13]. More flexible alternatives include conditional density estimation and deep generative models such as conditional generative adversarial networks [35], conditional variational autoencoders [36] and diffusion models [37].
In this work, we estimate aleatoric uncertainty using quantile regression models $\hat{r}_{\alpha/2},\hat{r}_{0.5},\hat{r}_{1-\alpha/2}$ (selected in Line 2 from PCS) trained on the residuals $Y_{i}-\hat{f}(X_{i})$. This approach offers improved stability: underfitting quantile regression directly on the $y$-values can severely distort aleatoric uncertainty estimates. In contrast, extreme underfitting on residuals, at worst, corresponds to assuming homoskedastic noise, which can be an acceptable bias. This improves the stability with respect to hyperparameters. To further improve the stability, we apply a PCS-inspired bagging strategy, by taking the empirical median
$\hat{q}^{\text{ale}}_{1-\alpha/2}(x):=\text{Median}\left[\left(\hat{r}_{1-\alpha/2}(x)-\hat{r}_{0.5}(x)\right)^{+}\right]\text{ and }\hat{q}^{\text{ale}}_{\alpha/2}(x):=\text{Median}\left[\left(\hat{r}_{0.5}(x)-\hat{r}_{\alpha/2}(x)\right)^{+}\right]$
over the ensemble members. Overall, it is computationally efficient and straightforward to implement.
2.4 CLEAR: combining aleatoric & epistemic uncertainty
To combine both aleatoric and epistemic uncertainties, we use a weighted scheme as in Equation (1). Specifically, using a PCS-type estimator $\hat{q}_{\alpha}^{\text{epi}}$ and a quantile regression estimator $\hat{q}_{\alpha}^{\text{ale}}$ trained on the residuals $Y_{i}-\hat{f}(X_{i})$, we define the prediction interval:
$\begin{split}C=\left[\hat{f}-\gamma_{1}\hat{q}_{\alpha/2}^{\text{ale}}-\gamma_{2}\hat{q}_{\alpha/2}^{\text{epi}},\quad\hat{f}+\gamma_{1}\hat{q}_{1-\alpha/2}^{\text{ale}}+\gamma_{2}\hat{q}_{1-\alpha/2}^{\text{epi}}\right].\end{split}$ (3)
Given a fixed ratio $\lambda=\frac{\gamma_{2}}{\gamma_{1}}$, we compute $\gamma_{1}$ on a held-out calibration set using the standard split conformal prediction procedure. While the natural choice $\gamma_{1}=\gamma_{2}$ may seem appealing, it is often suboptimal. The relative contribution of aleatoric and epistemic uncertainty can vary across datasets, and the corresponding estimators may differ substantially in scale and precision when $\gamma_{1}=\gamma_{2}$.
To choose $\lambda$ from data, we evaluate a grid of positive values $\Lambda$. For each candidate $\lambda\in\Lambda$, we construct the (calibrated) interval $C_{\lambda}$. To ensure the best trade-off between uncertainty sources, we select $\lambda^{\star}$ such that $C_{\lambda^{\star}}$ performs best under the chosen metric on $\mathcal{D}_{\text{val}}$. We have chosen quantile loss [31] (defined in Algorithm 1, also known as pinball loss) as a simple metric to balance both coverage and width. However, any other metric can also be used. As a proper scoring rule, quantile loss incentivizes truthfulness from a theoretical perspective (see Section B.3). This procedure is summarized in Algorithm 1.
Parameter $\lambda^{\star}$ balances aleatoric and epistemic uncertainties: if one estimator fails, $\lambda^{\star}$ compensates by re-weighting the other. When both estimators $\hat{q}^{\text{epi}}$ and $\hat{q}^{\text{ale}}$ are reliable (up to scaling), $\lambda^{\star}$ is interpretable. A large ratio
$\lambda^{\star}\tfrac{\hat{q}_{1-\alpha/2}^{\text{epi}}(x)+\hat{q}_{\alpha/2}^{\text{epi}}(x)}{\hat{q}_{1-\alpha/2}^{\text{ale}}(x)+\hat{q}_{\alpha/2}^{\text{ale}}(x)}\gg 1$
indicates that epistemic uncertainty dominates at $x$ (reducible with more data or stronger assumptions), while a small ratio $\ll 1$ indicates aleatoric uncertainty dominates (irreducible, though sometimes mitigated by adding covariates).
Algorithm 1 CLEAR: Calibrated Learning for Epistemic and Aleatoric Risk
1:Input: Data $(X_{i},Y_{i})$ for $i=1,\dots,n$, split into training $\mathcal{D}_{\text{train}}$, calibration $\mathcal{D}_{\text{cal}}$, and validation $\mathcal{D}_{\text{val}}$ (we consider $\mathcal{D}_{\text{cal}}=\mathcal{D}_{\text{val}}$); grid of $\lambda$ values $\Lambda$; significance level $\alpha$.
2:Step 1: Estimate epistemic uncertainty on $\mathcal{D}_{\text{train}}$.
3: Example: Estimate stable point predictor $\hat{f}$ and epistemic quantiles $\hat{q}_{\alpha/2}^{\text{epi}},\hat{q}_{1-\alpha/2}^{\text{epi}}$ using PCS ensembles across data perturbations.
4:Step 2: Estimate aleatoric uncertainty on $\mathcal{D}_{\text{train}}$.
5: Example: train a quantile regression model on the residuals $Y_{i}-\hat{f}(X_{i})$ to estimate conditional quantiles $\hat{q}_{\alpha/2}^{\text{ale}},\hat{q}_{1-\alpha/2}^{\text{ale}}$.
6:Step 3: Define prediction intervals for each $\lambda\in\Lambda$.
7: Define $C_{\lambda}$ by selecting the smallest value $\gamma_{1}$ such that the prediction set
$C_{\lambda}=\left[\hat{f}-\gamma_{1}\hat{q}_{\alpha/2}^{\text{ale}}-\lambda\gamma_{1}\hat{q}_{\alpha/2}^{\text{epi}},\,\hat{f}+\gamma_{1}\hat{q}_{1-\alpha/2}^{\text{ale}}+\lambda\gamma_{1}\hat{q}_{1-\alpha/2}^{\text{epi}}\right]$
contains at least $\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\rceil$ of points in $\mathcal{D}_{\text{cal}}$. See Section B.5 for implementation.
8:Step 4: Select $\lambda^{\star}$ that minimizes a chosen evaluation metric.
9: For any metric in Section D.5 (e.g., quantile loss), evaluate $C_{\lambda}(x)$ on $\mathcal{D}_{\text{val}}$ and set
$\lambda^{\star}=\arg\min_{\lambda\in\Lambda}\text{QuantileLoss}(\mathcal{D}_{\text{val}},C_{\lambda}),$
where $\text{QuantileLoss}(\mathcal{D}_{\text{val}},C_{\lambda}):=\frac{1}{|\mathcal{D}_{\text{val}}|}\sum_{i\in\mathcal{D}_{\text{val}}}\left[QL_{\alpha/2}\big(Y_{i},l(X_{i})\big)+QL_{1-\alpha/2}\big(Y_{i},u(X_{i})\big)\right]/2$, with $l(x),u(x)$ denoting the bounds of $C_{\lambda}(x)$, and $QL_{\tau}(y,q)=(y-q)\big(\tau-\mathds{1}_{(-\infty,q]}(y)\big)$.
10:Output: $\lambda^{\star}$ and calibrated prediction interval $C_{\lambda^{\star}}(x)$.
Lemma 2.1.
Let $\Lambda$ be compact. Suppose that at least $k$ of the base models used in the PCS ensemble are consistent for the true function $f(x)$, and the quantile regression estimators $\hat{q}_{\tau}^{\text{ale}}$ are consistent for both $\tau\in\{\alpha/2,1-\alpha/2\}$. Then we obtain asymptotic conditional validity: for any fixed $x\in\mathcal{X}$, it holds that
$\liminf_{|\mathcal{D}_{\text{train}}|,|\mathcal{D}_{\text{cal}}|\to\infty}\mathbb{P}\big(Y_{n+1}\in C(X_{n+1})\mid X_{n+1}=x\big)\geq 1-\alpha.$
Proof.
The consistency of the predictors implies $\hat{q}_{\tau}^{\text{epi}}(x)\to 0$, and the consistency of the quantile estimators implies $\hat{q}_{\tau}^{\text{ale}}(x)\to q_{\tau}^{\text{ale}}(x)$ point-wise. Hence, the estimated pre-calibrated intervals converge to their population analogues. Since $\sup_{\lambda\in\Lambda}\lambda\hat{q}_{\tau}^{\text{epi}}(x)\to 0$ converges to 0 uniformly over $\lambda\in\Lambda$ (due to compactness of $\Lambda$), therefore $C(x)$ is asymptotically equal to true conditional quantiles ${r}_{\alpha/2}(x),{r}_{1-\alpha/2}(x)]$; and necessarily $\lim_{|\mathcal{D}_{\text{cal}}|\to\infty}\gamma_{1}=1$, as is the case in classical CQR (see e.g. [8, Section 5]). ∎
Note that these assumptions are satisfied by the finite grid $\Lambda$ used in our implementation and by many base models, including tree-based methods and neural networks. We elaborate on the theoretical implications in Appendix B, with formal guarantees for marginal coverage in Lemma B.2.
3 Experimental Setup
3.1 Data
We conduct experiments using both simulations and real-world data. The synthetic experiments assess the theoretical guarantees of our approach. We sample $X\sim\mathcal{N}(0_{d},I_{d})$ and compute the response $Y=\mu(X)+\sigma(X)\cdot\varepsilon$, where $\varepsilon\sim\mathcal{N}(0,1)$. The mean function $\mu(X)$ introduces non-linearity through transformations of input features (involving absolute values and fractional powers with random coefficients); its explicit form, alongside $\sigma(X)$, is detailed in Appendix C. The sample size is fixed at $n=5000$ and divided into 70-30% training and validation splits. In the univariate case, we generate 100 datasets for $d=1$, and in multivariate, 100 datasets randomly sampled for $d\in\{2,3,20\}$. We then assess the conditional performance as a function of distance from the training data, where test points are randomly generated on the surfaces of spheres with varying radii. For the real-world scenarios, we use 17 regression datasets curated by [30], forming one of the largest benchmarks for UQ (see Section D.1 for details). Categorical features are one-hot encoded, with no further preprocessing. To ensure robustness, each dataset is evaluated over 10 random train-validation-test splits (60%-20%-20%). We conduct experiments in two configurations: standard (using $\mathcal{D}_{\text{val}}$ also as $\mathcal{D}_{\text{cal}}$) and conformalized (splitting the 20% validation into 10% $\mathcal{D}_{\text{val}}$ + 10% $\mathcal{D}_{\text{cal}}$ for stronger theoretical guarantees). Note that while the body of the paper only focuses on the standard experiments, the conformalized results are provided in Appendix G.
3.2 Baselines
We compare CLEAR against its core components, CQR and PCS ensemble (for details, including PCS implementation and further baselines that are relevant only to the appendices, see Appendix D). Notably, the underlying uncertainty estimation models from CQR and PCS are reused within the CLEAR framework. Our CQR follows the classical implementation [10] with the difference that in our experiments, we primarily utilize an enhanced variant, termed ALEATORIC, which uses bootstrapping ($b=100$) on $Y_{i}$ and the model selection from PCS. We further improve ALEATORIC in a new baseline called ALEATORIC-R, which models the residuals $Y_{i}-\hat{f}(X_{i})$ instead, using $\hat{f}$ from the corresponding PCS ensemble. The CLEAR method presented uses the uncalibrated aleatoric uncertainty estimate from this ALEATORIC-R and the uncalibrated epistemic uncertainty from PCS. For fairness and cross-comparison, the model type for quantile estimation within the ALEATORIC-R model is determined by the best-chosen model for epistemic uncertainty.
We perform ablation studies by exploring three different variants of base models for PCS (epistemic) and CQR (aleatoric). Our main approach, which we refer to as variant (a), uses a quantile PCS for estimating epistemic uncertainty. We employ a diverse set of models designed to estimate conditional quantiles, namely quantile random forests (QRF) [33], quantile XGBoost (QXGB) [38], and Expectile GAM [39]. Then, the top-performing model ($k=1$) on the validation set is selected and bootstrapped ($b=100$) to generate the epistemic uncertainty estimate $\hat{q}^{\text{epi}}$ and the median $\hat{f}$. CLEAR then combines this bootstrapped $\hat{q}^{\text{epi}}$ with the bootstrapped aleatoric estimate from ALEATORIC-R, ensuring ALEATORIC-R also uses the same selected quantile model. The other two variants are explained in Section D.4, both using the same $b$ and $k$ as above and only modifying the models. Variant (b) restricts quantile PCS as well as CQR baselines to only QXGB to remove any impact on CLEAR’s evaluation due to the choice of the base models. In contrast, variant (c) uses the standard PCS models to estimate the conditional mean. In all cases, PCS intervals are calibrated using the multiplicative method.
To further validate the generalizability of CLEAR, we also apply and evaluate it to other uncertainty estimators. This is a separate setup, where for epistemic uncertainty, we employ Deep Ensembles (DE) [22]—an ensemble of neural networks trained with diverse initializations—and for aleatoric uncertainty, we use Simultaneous Quantile Regression (SQR) [13], which directly models multiple conditional quantiles (more detail in Section D.3). These represent state-of-the-art deep learning approaches for uncertainty quantification, complementing our primary PCS and CQR baselines.
In all cases, CLEAR parameters ($\lambda,\gamma_{1}$) are optimized via quantile loss on the validation set. $\lambda$ is chosen from a dense grid combining linearly spaced values from 0 to 0.09 and logarithmically spaced values from 0.1 to 100 (totaling over 4000 points), and $\gamma_{1}$ is determined via conformal calibration for the chosen $\lambda$. All intervals are evaluated at 95% nominal coverage. In our benchmarks, we only address the model perturbations of PCS, not the data processing part, where we simplify the process to a single dataset ($N_{1}=1$), as in [30]. However, we show an example in Section 4.3 of how CLEAR can be applied in the $N_{1}>1$ setting from [40, Chapter 13].
3.3 Metrics
To evaluate the quality of our prediction intervals, we employ interval coverage (PICP), normalized interval width (NIW), average interval score loss and the quantile loss that are common in the interval prediction literature [41, 42, 43]. We also evaluate our work on Normalized Calibrated Interval Width (NCIW), defined as:
$\text{NCIW}(\hat{f},l,u)=\text{NIW}\left(\hat{f}-c_{\textrm{test-cal}}l,\hat{f}+c_{\textrm{test-cal}}u\right),$
with a calibration constant
$c_{\textrm{test-cal}}:=\operatorname*{arg\,min}_{c\geq 0}\left\{\text{PICP}\left(\hat{f}-cl,\hat{f}+cu\right)\geq 1-\alpha\right\}.$
In the standard configuration, all methods are calibrated on the validation set; in the conformalized configuration, calibration uses a separate 10% calibration split. In both cases the methods are already calibrated before evaluation on the test set, resulting in $c_{\textrm{test-cal}}\approx 1$. We primarily discuss NCIW and quantile loss, but provide the results for all the available metrics.
4 Results
4.1 Simulations
Figure 2 shows conditional coverage and interval width for the univariate homoskedastic case. While ALEATORIC-R achieve good coverage in high-density regions but under-cover in low-density or extrapolation regions. When aleatoric and epistemic uncertainties are correlated, all methods perform similarly in the data-rich region, but only CLEAR maintains correct conditional coverage throughout, adapting interval width as needed. Additional results for heteroskedastic and multivariate settings in Section C.2.2 show similar findings.
Figure 2: Results for univariate homoskedastic case averaged over 100 simulations: On the left, conditional coverage, and on the right, mean width, for $X=x$. It compares CLEAR, PCS, and ALEATORIC-R (bootstrapped CQR trained on residuals $Y_{i}-\hat{f}(X_{i})$). The dashed horizontal line is the target coverage level of 0.9. CLEAR adapts to maintain target coverage across the input space.
4.2 Real-world data
Figure 3 shows the Normalized Calibrated Interval Width (NCIW) and Quantile Loss for 95% prediction intervals across all datasets for our main approach, CLEAR (variant a), compared to several baselines. CLEAR consistently demonstrates superior performance, achieving better or comparable interval width and loss metrics while rigorously maintaining nominal coverage. The inset boxplots show that CLEAR (a) compared to PCS, ALEATORIC, and ALEATORIC-R has an improved quantile loss of 15.8%, 34.5%, and 9.4%, respectively. Similar relative increases are observed for NCIW, with PCS, ALEATORIC, and ALEATORIC-R exhibiting increases of 17.4%, 28.2%, and 3%. Moreover, CLEAR (a) was, in fact, the top-performing method on 15 of the 17 datasets, while remaining the most stable compared to the baselines. These trends hold across our other model variants as well (Appendix F for standard and Appendix G for conformalized results): both CLEAR (b) and CLEAR (c) exhibit very similar relative improvements (Figures 7 and 8 in Sections F.2 and F.3), with variant (a) remaining the strongest and most robust configuration.
Figure 3: Results for real-world data: Quantile loss and NCIW performance of different methods over 10 seeds normalized relative to CLEAR (baseline = 1.0) with error bars are $\pm 1$ $\sigma$. Lower values are better. The inset boxplot shows the average (%) relative increase of the metric over CLEAR. EPISTEMIC is PCS-UQ, ALEATORIC is bootstrapped CQR, and ALEATORIC-R uses residuals.
While we observed that setting $\lambda=1$ or $\gamma_{1}=1$ could marginally improve results if one had prior knowledge of uncertainty components—an unlikely scenario in practice—fully optimizing both parameters offers greater robustness. A limited size of the validation dataset can lead to overfitting of the two parameters, and incorporating some prior on them could further improve the results. Importantly, the absolute value of $\lambda$ is relative to the pre-calibrated scales of the uncertainty estimators and dataset noise; its interpretation is best contextualized by observing its behavior when systematically varying dataset characteristics, such as the number of features or observations. While fixed parameters ($\lambda=1$ or $\gamma_{1}=1$) could marginally improve results with prior knowledge, fully optimizing both parameters offers greater robustness. CLEAR’s dual-parameter calibration enhances stability by adaptively re-weighting potentially unreliable uncertainty components, as evidenced in datasets like energy_efficiency, where baselines show markedly larger NCIW. The dataset-dependent variability in optimal $\lambda$ underscores the need for adaptive selection over fixed heuristics. The calibration runtime (grid-search) is also extremely negligible in practice (Section F.5).
Beyond PCS and CQR, CLEAR demonstrates substantial improvements when applied to DE and SQR (Table 1). When using DE for epistemic uncertainty and SQR for aleatoric uncertainty at 95% nominal coverage, CLEAR achieves average width reductions (NCIW) of 28.6% and 13.4%, respectively, with similar improvements in quantile loss (24.0% and 13.7%). The gains persist even after conformal calibration of the baselines, particularly relevant to the aleatoric component (SQR). The results underscore CLEAR’s ability to balance the two uncertainty sources well, regardless of the underlying estimators. This consistency across both PCS and neural approaches validates the generality of CLEAR (full results in Appendix E).
Table 1: Mean (%) improvement of CLEAR over DE & SQR across 17 datasets (higher is better).
Metric DE SQR DE-conformal SQR-conformal
PICP +0.05% -0.66% -0.09% -0.15%
NIW +28.81% +17.38% +29.55% +14.07%
NCIW +28.57% +13.36% +27.90% +13.23%
QuantileLoss +23.98% +13.66% +24.08% +10.12%
4.3 Case study: Accounting For Data Uncertainty
We consider a case study on the Ames Housing dataset, detailed in Appendix H, where we demonstrate the full PCS pipeline and vary the number of predictor variables used to model housing prices. Starting from approximately 80 features, we construct a reduced version using only the top two predictors; this setup naturally balances aleatoric uncertainty (increases due to limited information) and epistemic uncertainty (decreases due to reduced model complexity). Table 2 reveals that while PCS performs better with many predictors and CQR excels with fewer features, CLEAR adapts effectively to both scenarios through its calibrated uncertainty combination. CLEAR estimates $\lambda=0.6$ in the 2-variable case (prioritizing aleatoric uncertainty) versus $\lambda=14.5$ in the full-feature case (emphasizing epistemic uncertainty), with corresponding epistemic-to-aleatoric ratios of 0.03 and 7.72. This adaptive weighting enables CLEAR to maintain sharp intervals across both scenarios, with nearly halved average interval width when using all features.
Table 2: Ames Housing results (90% coverage target).
Experiment Method Coverage Average Width (\$) Quantile Loss NCIW
2 features PCS 0.87 107,880 3,818 0.213
CQR 0.89 104,741 3,448 0.186
CLEAR 0.89 98,571 3,191 0.179
All features PCS 0.89 57,594 1,922 0.105
CQR 0.87 62,398 2,194 0.117
CLEAR 0.89 56,745 1,873 0.102
5 Conclusion, Limitations & Future Work
This paper introduces CLEAR, a novel framework for constructing prediction intervals by adaptively balancing epistemic and aleatoric uncertainty. Through a calibration process involving two distinct parameters, $\gamma_{1}$ and $\gamma_{2}$, CLEAR offers an improvement over classical methods that often address these uncertainty types in isolation or rely on their fixed, non-adaptive combination. Our evaluations using CQR (and SQR) for aleatoric uncertainty and PCS (and DE) for epistemic uncertainty show that interval width and quantile loss improve on both simulated and the 17 real-world datasets. CLEAR consistently achieves improved conditional coverage, notably by adapting interval widths appropriately in extrapolation regions, while yielding narrower interval widths.
Limitations remain despite CLEAR’s significant advantages, and these warrant further discussion. When scaling CLEAR to significantly larger datasets and models, specific hyperparameters (e.g., number of bootstraps) can be adjusted to save computational costs. The accuracy of CLEAR is intrinsically linked to the quality of the base estimators for epistemic and aleatoric uncertainty. All our calibration datasets contained at least 150 data points, but for smaller calibration datasets, overfitting $\gamma_{1}$ and $\lambda$ could be a problem. Epistemic uncertainty, as demonstrated in the case study, must be addressed through careful judgment calls, in line with the principles of the PCS framework. The additive combination of scaled uncertainties, while powerful, is also a specific structural choice.
Future work could explore extension to classification tasks [44], alternative $\lambda$ selection techniques, integration with active learning, and more scalable epistemic UQ approaches [45]. Extending CLEAR to time series settings could also be valuable, particularly for capturing the dynamics of temporal uncertainty. Finally, a deeper study of the interpretability of learned parameters could provide insights into the dominance of the two uncertainty sources.
Acknowledgments and Disclosure of Funding
We are grateful to Abhineet Agarwal and Michael Xiao for their valuable feedback, helpful discussions, and for sharing early previews and detailed explanations of their code. We also thank Natasa Tagasovska, João Vitor Romano, Erez Buchweitz, Omer Ronen, Jakob Weissteiner, Hanna Wutte, Florian Krach, Markus Chardonnet, Josef Teichmann, Bertrand Charpentier, Janis Postels, Alexander Immer, James-A. Goulet, and Hans Bühler for insightful and stimulating discussions. We are also grateful to Valérie Chavez-Demoulin for her continuous support.
Juraj Bodik acknowledges support from the Hasler Foundation [grant number 24009]. Part of this work was conducted during his research visit to the UC Berkeley, supported by the Hasler Foundation. He thanks the Department of Statistics at UC Berkeley for their kind hospitality.
Jakob Heiss acknowledges support from the Swiss National Science Foundation (SNSF) Postdoc.Mobility fellowship [grant number P500PT_225356]. Early ideas for this work began to take shape while he was affiliated with the Department of Mathematics, ETH Zürich, Switzerland.
Ilia Azizi acknowledges support from the HEC Research Fund grant that enabled and facilitated computations on DCSR clusters at the University of Lausanne.
We gratefully acknowledge partial support from NSF grant DMS-2413265, NSF grant DMS 2209975, NSF grant 2023505 on Collaborative Research: Foundations of Data Science Institute (FODSI), the NSF and the Simons Foundation for the Collaboration on the Theoretical Foundations of Deep Learning through awards DMS-2031883 and 814639, NSF grant MC2378 to the Institute for Artificial CyberThreat Intelligence and OperatioN (ACTION), and NIH (DMS/NIGMS) grant R01GM152718.
Appendix A Related Literature
In the main paper, Section 1 provides a high-level overview of uncertainty quantification in ML. Section 2.2 contains an overview of epistemic uncertainty, and Section 2.3 discusses the aleatoric uncertainty. In this appendix, we discuss these concepts in more depth. For an introduction to epistemic and aleatoric uncertainty, see [11], which provides a comprehensive overview.
A.1 Literature that implicitly assumes $\gamma_{1}=1$
UACQR, introduced by [19], is conceptually closest to CLEAR. Conceptually on a high level, this method corresponds to a variant of CLEAR where $\gamma_{1}=1$ is fixed, and only $\lambda$ is tuned. This is justified when aleatoric uncertainty is well-calibrated, which may often hold asymptotically. When validation dataset is very small, setting $\gamma_{1}=1$ can even bring small advantages compared to tuning $\gamma_{1}$. In practice, however, aleatoric uncertainty is often miscalibrated due to over- or under-regularization. Tuning $\gamma_{1}$ helps correct its scale. Furthermore, it can happen in practice that the estimator of the aleatoric uncertainty has a much lower or much higher quality than the epistemic uncertainty. In this case, optimizing both parameters $\gamma_{1}$ and $\lambda$ allows us to compensate for the failure of one of the two uncertainties to some extent by putting more weight on the other type of uncertainty without changing the marginal coverage. This results in a higher robustness and stability of CLEAR.
Empirical Comparison
When comparing against UACQR, we see in Tables 24, 25 and 26 that the performance of CLEAR is much more reliable across the considered datasets and metrics. For example, on the airfoil, energy efficiency and naval propulsion CLEAR is approximately 2 to 3 times better in metrics such as Quantile Loss, NCIW and Average Interval Score Loss, while UACQR never outperforms CLEAR by more than 27% in any of the datasets in any of the considered metrics. Sometimes UACQR-P can also output infinitely wide predictive intervals [19, p. 5], which we observed for one dataset in our experiments. CLEAR performs better than both versions of UACQR in 14 out of 17 datasets. These large difference in performance in Section F.3 can partially be explained by our approach to fit the aleatoric uncertainty on the residuals. However, in Sections F.1 and F.2 every method uses exactly the same module for estimating the point prediction $\hat{f}$, the uncalibrated epistemic uncertainty $q^{\text{epi}}$, and the uncalibrated $q^{\text{ale}}$. This allows us to only compare the high-level concepts of fixing $\gamma_{1}=1$ versus optimizing both $\gamma_{1}$ and $\lambda$. In this case the gap narrows down substantially, but still CLEAR is more robust. Fixing $\gamma_{1}=1$ can never outperform CLEAR by more than 10% on any dataset with respect to any metric. We hypothesize that CLEAR is more stable and robust since it can compensate for shortcomings of the base models more easily: If the aleatoric uncertainty is overestimated or underestimated, CLEAR can correct its scale by adjusting $\gamma_{1}$.
A.2 Literature that implicitly assumes $\lambda=1$
Most of the literature that combines epistemic and aleatoric uncertainty implicitly assumes that $\lambda=1$, when they simply combine epistemic and aleatoric uncertainty in the ratio 1:1. On top of the resulting uncertainty, one can use (conformal) calibration, which corresponds to CLEAR’s calibration of $\gamma_{1}$. However, in contrast to CLEAR, they do not rebalance the ratio of epistemic and aleatoric uncertainty with $\lambda$.
In practice, it is possible to underestimate the aleatoric uncertainty and overestimate epistemic uncertainty. This results in a one-dimensional calibration via $\gamma_{1}$, and consequently, narrow intervals in regions of dominating aleatoric uncertainty or in wide intervals in regions of dominating epistemic uncertainty. $\gamma_{1}$ alone cannot solve this. CLEAR can deal well with such situations by compensating for such an imbalance via $\lambda$.
A.2.1 Deep Ensembles
While it is quite common to refer to deep ensembles (DE), whenever one uses an ensemble of neural networks (NNs) for uncertainty estimation, it is important to note that [22] introduced DE as a method which both estimates epistemic and aleatoric uncertainty. For regression, they train each NN with 2 outputs estimating $\mu$ and $\sigma$ via a Gaussian Maximum-Likelihood-loss (as in [34]), where $\sigma$ is responsible for the aleatoric uncertainty, which they refer to as “ambiguity in targets y for a given x”. Moreover, they use the ensemble diversity to estimate epistemic uncertainty, which they refer to as “model uncertainty”. Although DE is mainly known for its ensembling approach, [22, Table 2 in Appendix A.1] clearly shows that both the aleatoric and epistemic parts are crucial in terms of empirical performance. In their paper, the authors do not apply any calibration on top, i.e., as it is implicitly assumed that $\gamma_{2}=\gamma_{1}=1=\lambda$. However, it is common to apply (conformal) calibration on top. The commonly used calibration techniques only calibrate $\gamma_{1}$, while keeping $\lambda=1$ fixed, in contrast to CLEAR. In particular, for DE, both the diversity of the ensemble and the bias on aleatoric are very sensitive to various hyperparameters. One type of uncertainty may be strongly under-estimated. In contrast, the other type is strongly over-estimated, which motivates the need to explicitly calibrate $\lambda$ in a data-driven way.
Technical Details on Adversarial Attacks for Deep Ensembles.
The original paper [22] is written as if applying adversarial attacks is an integral part of DEs and one of the paper’s main contributions. However, to the best of our knowledge, many practitioners refer to “DEs” without implying adversarial attacks during training, and adversarial attacks during training are seen as an optional add-on to DEs, but not an important part of DEs. Furthermore, the regression results in [22, Table 2 in Appendix A.1] do not suggest that adversarial attacks during training are particularly beneficial to the performance.
Technical Details on Measuring Deep Ensemble’s diversity.
Intuitively, one should estimate high epistemic uncertainty if there is a significant disagreement among the ensemble’s predictions and small epistemic uncertainty if they agree. While [40, Chapter 13] and [30] suggest using quantiles, [22] suggest using the empirical standard deviation to estimate DE’s disagreement among ensemble members’ predictions. It seems plausible that in [22]’s case of 5 ensemble members, using the standard deviation and some Gaussian assumptions can be more appropriate while in [40, Chapter 13]’s and [30]’s case of 100 or more ensemble members, quantiles can be more accurate.
Variations of Deep Ensembles.
There exist several modifications of DE that, for example, promote the ensemble’s diversity on the function space via an additional loss term during training [46], ensemble over multiple different hyperparameters [47], or reduce the computational training cost [17, 26, 48, 49, 19, 50, 30].
A.2.2 Monte Carlo Dropout
[26] originally studied Monte Carlo Dropout (MC Dropout) without explicitly modeling aleatoric uncertainty, which was then extended by [17] to also explicitly model aleatoric uncertainty. While [26] see MC Dropout as an approximation of a Bayesian neural network, one can also see it as another ensemble method. The main difference to DE is that MC Dropout only trains one NN with dropout and obtains an ensemble after training by randomly setting weights of the model to zero at inference time. Analogously to DE, MC Dropout also adds epistemic and aleatoric uncertainty in the ratio 1:1 with the same disadvantages as described before in Section A.2.1.
A.2.3 “A Deeper Look into Aleatoric and Epistemic Uncertainty Disentanglement”
[51] empirically concludes that ensembles have the best uncertainty and disentangling behavior of epistemic and aleatoric uncertainty. In their paper, the authors do not use any form of calibration. This would correspond to $\gamma_{2}=\gamma_{1}=1=\lambda$ in our notation. Their paper suggests a different loss function, which they call $\beta$-NLL, for training to mitigate the underestimation of aleatoric uncertainty to some extent in their experimental setting without comparing this approach to calibration. In [51, Figure 6], one can observe that even without the $\beta$-NLL, both epistemic and aleatoric uncertainty already have a good shape (that is, good relative uncertainty, see Appendix I) for deep ensembles (DE). The main problem of DE in [51, Figure 6] is that the epistemic uncertainty is too small by a very large factor (that is, poor absolute scale of uncertainty, see Appendix I), while aleatoric uncertainty has already almost the correct scaling. From our perspective, applying CLEAR on their DE would probably largely fix their problem of DE if CLEAR chooses $\gamma_{1}\approx 1$ and $\lambda\gg 1$. However, they show that for this specific experiment, $\beta$-NLL also fixes the problem. In general, where one suspects that the aleatoric or the epistemic uncertainty might be too small or too large across the domain, we strongly recommend simply applying CLEAR on top of the already trained ensemble instead of retraining all the models with a new training pipeline. The concept of CLEAR can be implemented in a few minutes and calibrates an already trained ensemble in a few seconds. We considered an interesting open problem to study if $\beta$-NLL can improve the relative epistemic and aleatoric uncertainty (see Appendix I).
A.2.4 “Recalibration of Aleatoric and Epistemic Regression Uncertainty in Medical Imaging”
[14] also combines epistemic and aleatoric uncertainty in the ratio 1:1 and applies a single constant (corresponding to $\gamma_{1}$ in our notation) to scale the total predictive uncertainty, which corresponds to fixing $\lambda=1$, resulting in the same potential for problems as mentioned before.
A.3 Bayesian Models
If one had access to a perfect prior, perfectly computed Bayesian inference, it would provide well-calibrated epistemic and aleatoric uncertainty, at least in theory. However, in practice, the ratio of estimated epistemic and aleatoric uncertainty can be very wrong, and the uncertainty can be miscalibrated.
A.3.1 Gaussian Process Regression (GPR)
Gaussian Process regression [52] provides a closed form for exact posterior epistemic uncertainty for a given Gaussian process as a prior and for a given known noise Gaussian noise distribution. If the (scale of the) prior or the scale (of the noise) is misspecified, the ratio of estimated epistemic and aleatoric uncertainty can be arbitrarily bad, and the uncertainty can be miscalibrated. Therefore, it is common to optimize hyperparameters like the noise scale and the prior scale (typically in a non-Bayesian way). The main differences to CLEAR are that 1) for GPR, one has to refit the model for every considered possibility of hyperparameters, while CLEAR optimizes $\gamma_{1}$ and $\lambda$ after fitting the model, resulting in much lower computational costs; 2) CLEAR can be applied to other base models as well such as tree-based models which are more popular in many applications; 3) standard-implementations of GPR fit the hyperparameter on the training data rather than on the validation data.
A.3.2 Bayesian Neural Networks (BNNs)
Bayesian neural networks [28, 53] offer a principled Bayesian framework for quantifying both epistemic and aleatoric uncertainty through the placement of a prior distribution on network weights. As for GPR, the ratio of estimated epistemic and aleatoric uncertainty in BNNs is highly sensitive to the choice of prior. Consequently, we advocate applying CLEAR to an already trained BNN, calibrating both uncertainty types via scaling factors $\gamma_{1}$ and $\gamma_{2}$ with negligible additional computational overhead. While exact Bayesian inference in large BNNs is computationally intractable, numerous approximation techniques have been proposed [54, 55, 56, 26, 22, 29, 57, 27, 58, 59, 60, 61, 62]. Interestingly, theoretical [63, 64] and empirical [58] studies suggest that some of these approximations can actually provide superior estimates compared to their exact counterparts, due to poor choices of priors, such as i.i.d. Gaussian priors, in certain settings.
A.3.3 EPISCORE
The recent work by [21] introduces EPICSCORE. Similar to our work, this method addresses the limitations of standard conformal prediction in capturing epistemic uncertainty. EPICSCORE focuses on enhancing existing conformal scores with Bayesian epistemic uncertainty. They also compute the average interval score for 95% predictive intervals on the datasets airfoil (where their best method out of 6 varaints performs more than 2 times worse than the worst of the 3 variants of CLEAR), concrete (where their best method performs more than 1.5 times worse than the worst variant of CLEAR) and Superconductivty (where their best method performs approximately 1.3 times worse than the worst variant of CLEAR).
A.3.4 TabPFN
TabPFN [65, 66] is a transformer trained to emulate Bayesian inference over a diverse prior of realistic tabular problems, achieving strong predictive uncertainty estimates with a single forward pass. Its prior spans diverse noise structures and function classes, yielding uncertainty estimates that inherently mix aleatoric and epistemic components. Although recent variants such as TabPFN-TS [67] extend its applicability to time series, the method is limited to small tabular datasets and does not disentangle uncertainty types. In contrast, our approach explicitly separates and calibrates epistemic and aleatoric uncertainty and scales to arbitrary modalities and data sizes.
A.4 Conformal Literature that does not account for Epistemic and Aleatoric Uncertainty
The classical conformal prediction literature [4, 6, 24, 10, 8] primarily focuses on achieving marginal coverage, often with attention to asymptotic conditional coverage. However, it often overlooks epistemic uncertainty, which is especially critical in finite-sample settings. While [25] established the impossibility of achieving exact distribution-free conditional coverage in finite samples, several recent works attempt to improve coverage guarantees in more restricted settings under certain assumptions. For instance, [9] propose coverage guarantees over a subclass of distribution shifts, effectively interpolating between marginal and conditional coverage. Others, such as [68] and [69], provide guarantees over a finite set of prespecified subgroups. We argue that to obtain reliable conditional coverage, one must model the uncertainty arising from each stage of the data science life cycle [40], and appropriately integrate these uncertainties to achieve meaningful coverage guarantees.
A.5 Other Calibration Methods
Predictive uncertainty can be expressed either as a predictive set for a given level $\alpha$, as a predictive distribution, or as a numerical value quantifying the level of uncertainty (e.g., the entropy of the predictive distribution). Our implementation of CLEAR yields predictive intervals, making it compatible with both base models that output intervals (such as QR) and with base models that output predictive distributions from which one can easily obtain predictive intervals. Conceptually, CLEAR could also be extended to predictive distributions or numerical values.
In the applied literature on predictive intervals, it is quite common to calibrate a constant factor to rescale the width of predictive intervals [14, 27, 40, 30]. Independently, the conformal community came up with almost identical methods based on theoretical considerations, as already discussed in Section A.4.
In the literature on predictive distributions, early works suggested more sophisticated non-linear transformations for the predictive distributions [3, 70, 71]. However, [72] demonstrated that simply calibrating a constant scaling factor performs on par with these more sophisticated calibration methods.
To summarize, there are multiple (slightly) different methods to calibrate uncertainty with similar empirical performance. These methods only monotonically transform the uncertainty without changing the ranking of the uncertainties (i.e., if one was more uncertain on $x$ than on $\tilde{x}$ before the calibration step, one will also be more uncertain on $x$ than on $\tilde{x}$ afterwards; see Appendix I). In contrast, CLEAR and UACQR can also change the ranking of the uncertainties (e.g., in the situation of Figure 1, a small value of $\lambda$ would assign more uncertainty to the center around $x=0$ compared to the region around x=6x=-6, whereas a large value of $\lambda$ would assign more uncertainty to x=6x=-6 than to $x=0$).
A.6 Methods that Mainly Focus on Distrbutional Aleatoric Uncertainty
In this section, we provide more details on the literature overview on aleatoric uncertainty from Section 2.3.
One can directly estimate conditional quantiles for a given level $\alpha$ via quantile regression (QR) [31], either parametric or nonparametric, such as smooth quantile regression [32] and quantile random forests (QRF) [33]. This is computationally cheap if you a priori know which level $\alpha$ is of interest for your predictive intervals.
Alternatively, one can also try to estimate the conditional distribution. This can initially be computationally more expensive, but once the conditional distributions are estimated, one can easily obtain conditional quantiles for multiple different levels $\alpha$. One of the computationally cheapest ways to estimate conditional distributions is to estimate parameters of a specific distribution (e.g., $\mu(x)$ and $\sigma(x)$ of a Gaussian distribution) [34]. There are semi-parametric extensions of this with universal approximation properties [73]. There are many similar approaches to estimate conditional densities [74]. Another way to obtain a conditional distribution is to estimate the conditional quantiles for all levels $\alpha\in[0,1]$ simultaneously, as in simultaneous quantile regression (SQR) [13] and its extension MAQR [75]. Especially for applications where the model output is high-dimensional (e.g., models that output images), more modern deep generative models have become popular, e.g., conditional generative adversarial networks [35], conditional variational autoencoders [36], and diffusion models [37, 50].
Limitations and Empirical Comparison
The majority of these works fail to properly account for epistemic uncertainty222[50] include epistemic uncertainty in the ratio 1:1; [75] conduct limited experiments on including epistemic uncertainty in their appendix; [13] proposes Orthonormal Certificates for estimating epistemic uncertainty, but explicitly leave the combination for future work., and don’t include a post-hoc calibration step. Each of these methods could be incorporated into CLEAR as an alternative to our recommended ALEATORIC-R module. For example, in some preliminary experiments combining SQR [13] with ensemble-diversity [76] via CLEAR’s calibration step results in significant improvements over the (calibrated versions) of the individual models. This shows the applicability of CLEAR’s calibration step on pure deep-learning pipelines. Furthermore, when we compare the results of our entire CLEAR pipeline reported in our paper with the results of the entire SQR-pipeline and MAQR-pipeline reported in [13, 75], we can see that CLEAR massively outperforms their pipelines. This shows that our PCS-based approach and our novel variant of QR are already highly recommendable choices, resulting in state-of-the-art performance of the CLEAR-pipeline directly out of the box.
A.7 Further Related Literature
A.7.1 “Beyond Pinball Loss: Quantile Methods for Calibrated Uncertainty Quantification”
[75] explores the limitations of the pinball loss (which we call Quantile Loss), criticizing that its direct minimization does not guarantee correct marginal coverage. In CLEAR, we address this by framing the optimization of $\gamma_{1}$ and $\lambda$ as a constrained optimization problem (see Equation 4 in Section B.3): we minimize the pinball loss on a calibration dataset $\mathcal{D}_{\text{cal}}$ (see 9 in Algorithm 1) while constraining the marginal coverage on $\mathcal{D}_{\text{cal}}$ (see 7 in Algorithm 1). This constrained approach directly mitigates the criticism of [75].
A.7.2 Uncertainty Quantification for Conditional Image Generation
[50] suggests combining a diffusion model for aleatoric uncertainty with a hyper-network-generated ensemble for epistemic uncertainty to quantify the uncertainty in conditional image generation. Diffusion models are particularly well-suited for conditional image generation, and using a hyper-network can be a computationally more efficient way to obtain an ensemble. However, they combine these two sources of uncertainty simply in the ratio 1:1 without any form of calibration (corresponding to hard-coding $\lambda=\gamma_{1}=\gamma_{2}=1$ in our notation). It would be interesting future work to apply calibration in the spirit of CLEAR on top of [50] to extend the idea of CLEAR to conditional image generation.
A.7.3 Uncertainty Quantification for Pretrained Models
[77] estimates the epistemic uncertainty for pre-trained classification models, while CLEAR focuses on both epistemic and aleatoric components for regression. It would be interesting future work to extend CLEAR to also be applicable to already pre-trained models using techniques presented in [77].
A.7.4 Uncertainty Quantification as a Binary Classification Problem
[78] do not provide predictive intervals or distributions. Instead, they partition test data into “good” (more certain) and “bad” (less certain) points. As future work, one could derive a binary classifier from CLEAR by thresholding the predictive interval width, and then evaluate it under their proposed metrics.
A.7.5 Credal Sets
[79] describe epistemic uncertainty via credal sets on the space of distributions. This approach is quite natural for classification, where the space of distributions over $K$ classes is (K1)(K-1)-dimensional. However, they do not provide any concrete algorithms for regression, where the space of all distributions over a continuous set is infinite-dimensional. [79] suggest multiple ideas on how one can translate credal sets into two real/valued numbers describing the magnitude of the epistemic and the aleatoric uncertainty. [80] proposes a method that can provide conformal predictive sets for classification with guaranteed input-conditional coverage under the assumption that one already has access to credal sets that guaranteeably cover the true input/conditional distribution. They also briefly discuss that for real-world applications, this assumption is not satisfied.
A.7.6 In-sample calibration yields conformal calibration guarantees
For the ditributional regression, [81] suggest using conformal binning or conformal isotonic distributional regression and prove theoretical guarantees under the exchangability assumption.
A.7.7 Engression
[82] uses generative methods for uncertainty estimation that, in particular, take advantage of pre-additive noise (i.e., noise that is directly added to $x$).
Appendix B Theory: Coverage Guarantees and Theoretical Justifications
B.1 Finite-sample marginal coverage for CLEAR
While conformal methods typically offer finite-sample marginal coverage guarantees, our implementation of CLEAR does not strictly adhere to these guarantees for two reasons:
  1. 1.
    We reuse the validation dataset $\mathcal{D}_{\text{val}}$ as the calibration dataset, i.e., $\mathcal{D}_{\text{cal}}=\mathcal{D}_{\text{val}}$.
  2. 2.
    We optimize two parameters $\gamma_{1}$ and $\lambda$ on the calibration dataset $\mathcal{D}_{\text{cal}}$.
However, our empirical results in Tables 12, 17 and 22 show that for reasonably sized datasets, this theoretical discrepancy does not visibly impact our practical marginal coverage. From a theoretical perspective, CLEAR still achieves asymptotic marginal coverage (as a consequence of asymptotic conditional coverage, see Lemma 2.1).
By slightly modifying the CLEAR procedure, we can obtain a theoretical finite-sample guarantee for marginal coverage under the standard exchangeability assumption, as in conformal inference, without sacrificing asymptotic conditional coverage. The key idea is to optimize the parameter $\lambda$ using only a validation dataset $\mathcal{D}_{\text{val}}$, and then calibrate $\gamma_{1}$ using a separate, previously unseen calibration dataset $\mathcal{D}_{\text{cal}}$.
Definition B.1 (Conformalized CLEAR).
Let the available data be split into a training set $\mathcal{D}_{\text{train}}$ and an old validation set $\mathcal{D}_{\text{val}}^{\text{old}}$. We first split $\mathcal{D}_{\text{val}}^{\text{old}}$ into two disjoint sets: a new validation set $\mathcal{D}_{\text{val}}$ and a calibration set $\mathcal{D}_{\text{cal}}$. The procedure is as follows:
  1. 1.
    Train and Optimize $\lambda$ by running Algorithm 1 on $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{val}}$: The model $\hat{f}$ and the uncertainty estimators are trained on $\mathcal{D}_{\text{train}}$ (Step 2 and 3 of Algorithm 1). The optimal value $\lambda^{*}$ is found by optimizing the QuantileLoss on $\mathcal{D}_{\text{val}}$, without using any data from $\mathcal{D}_{\text{cal}}$ (Step 4 and 5 of Algorithm 1).
  2. 2.
    Compute Conformity Scores: For each data point $(X_{i},Y_{i})\in\mathcal{D}_{\text{cal}}$, compute the conformity score $S_{i}^{\lambda^{*}}$:
    $S_{i}^{\lambda^{*}}=\max\left\{\frac{\tilde{l}_{\lambda^{*}}(X_{i})-Y_{i}}{\hat{f}(X_{i})-\tilde{l}_{\lambda^{*}}(X_{i})},\frac{Y_{i}-\tilde{u}_{\lambda^{*}}(X_{i})}{\tilde{u}_{\lambda^{*}}(X_{i})-\hat{f}(X_{i})}\right\}$
    where $\tilde{l}_{\lambda^{*}}:=\hat{f}-\hat{q}^{\text{ale}}_{\alpha/2}-\lambda^{*}\hat{q}^{\text{epi}}_{\alpha/2}$ and $\tilde{u}_{\lambda^{*}}:=\hat{f}+\hat{q}^{\text{ale}}_{1-\alpha/2}+\lambda^{*}\hat{q}^{\text{epi}}_{1-\alpha/2}$.
  3. 3.
    Calibrate the Prediction Interval: Set the calibration parameter $\gamma_{1}^{*}$ to be the $\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\rceil$-th smallest value among the conformity scores $S_{i}^{\lambda^{*}}$ from $\mathcal{D}_{\text{cal}}$. If $\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\rceil>|\mathcal{D}_{\text{cal}}|$, set $\gamma_{1}^{*}=\infty$.
  4. 4.
    Form the Final Prediction Interval: For a new test point $x_{\text{new}}$, the final $(1-\alpha)$-prediction interval is given by:
    $C(x_{\text{new}})=\left[\hat{f}(x_{\text{new}})-\gamma_{1}^{*}\left(\hat{q}^{\text{ale}}_{\alpha/2}+\lambda^{*}\hat{q}^{\text{epi}}_{\alpha/2}\right),\quad\hat{f}(x_{\text{new}})+\gamma_{1}^{*}\left(\hat{q}^{\text{ale}}_{1-\alpha/2}+\lambda^{*}\hat{q}^{\text{epi}}_{1-\alpha/2}\right)\right]$
This modified version of CLEAR, which we call Conformalized CLEAR, satisfies the standard finite-sample marginal coverage guarantee of conformal prediction.
Lemma B.2.
Under the assumption that the data points in the calibration set $\mathcal{D}_{\text{cal}}$ and the test point $(X_{\text{new}},Y_{\text{new}})$ are exchangeable, the prediction interval $C(X_{\text{new}})$ generated by the Conformalized CLEAR procedure satisfies:
$\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}}))\geq 1-\alpha.$
Proof.
The proof is a direct application of the standard theoretical guarantees for split conformal prediction. Since $\lambda^{*}$ is determined using only data from $\mathcal{D}_{\text{val}}$, it is fixed with respect to the calibration set $\mathcal{D}_{\text{cal}}$. The conformity scores $S_{i}^{\lambda^{*}}$ are therefore exchangeable for all $(X_{i},Y_{i})\in\mathcal{D}_{\text{cal}}\cup\{(X_{\text{new}},Y_{\text{new}})\}$. The choice of $\gamma_{1}^{*}$ as the empirical $(1-\alpha)(1+1/|\mathcal{D}_{\text{cal}}|)$-quantile of the calibration scores ensures that the resulting prediction interval achieves at least $1-\alpha$ marginal coverage, as established by [8, Theorem 1.4]. ∎
However, in practice, we recommend using our default implementation of CLEAR, where the validation data is reused for calibration ($\mathcal{D}_{\text{cal}}=\mathcal{D}_{\text{val}}$). This improves data efficiency while still achieving strong approximate marginal coverage in our experiments (see Tables 12, 17 and 22). Avoiding a split allows more data for both validation and calibration, which improves model selection and stabilizes marginal and conditional coverage. Even when marginal coverage $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}}))\geq 1-\alpha$ is satisfied, the actual coverage $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{cal}})$ can be significantly lower than the target coverage $1-\alpha$ for a fixed calibration dataset $\mathcal{D}_{\text{cal}}$, due to the inherent variation of $\mathcal{D}_{\text{cal}}$. Only with sufficiently large calibration datasets $\mathcal{D}_{\text{cal}}$ can we expect $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{cal}})$ to be reliably close to $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}}))$. [^3] Conformal theory provides theoretical guarantees for $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}}))\geq 1-\alpha$ and $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}})\geq 1-\alpha$ under the standard exchangeability assumption. However, it does not provide finite-sample guarantees for $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{val}},\mathcal{D}_{\text{cal}})$ and $\mathbb{P}(Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{cal}})$ for a fixed calibration dataset, even under the standard exchangeability assumption.
B.1.1 Limitations of Conformal Marginal Coverage Guarantees
The conformal theory heavily relies on the assumptions of exchangeability. Exchangability means that the joint distribution of calibration and test observations is invariant to permutations (e.g., iid observations satisfy this assumption). While exchangeability is theoretically convenient, it is unrealistic in many real-world settings. Models are typically trained on past data and deployed in the future, where the distribution of $X_{\text{new}}$ usually shifts, i.e., $\mathbb{P}[X_{\text{new}}]\neq\mathbb{P}[X]$. Even if $\mathbb{P}[Y_{\text{new}}|X_{\text{new}}]=\mathbb{P}[Y|X]$ remains fixed, such marginal shifts in $X_{\text{new}}$ can cause conformal methods to catastrophically fail to provide valid marginal coverage. In Sections 4.1 and C (Figures 2, 4, 5 and 6), CLEAR empirically remains robust, while CQR and Naive-Conformal fail under distribution shifts in $X_{\text{new}}$. E.g., Figure 4, suggests $\mathbb{P}[Y_{\text{new}}\in C_{\text{Naive}}(X_{\text{new}})\mid|X_{\text{new}}|\geq 2]\leq 70\%\ll 90\%=1-\alpha$, thus a marginal distribution shift of $X_{\text{new}}$ that strongly increases the probability of $|X_{\text{new}}|>2$, would lead to a large drop of marginal coverage for $(X_{\text{new}},Y_{\text{new}})$. CLEAR likewise lacks formal guarantees under extreme shifts, but consistently performs more reliably across our experiments. In the case study (Section 4.3), a realistic temporal split (see [40, Section 8.4.3]) also violates exchangeability, and CLEAR outperforms conformal baselines in marginal coverage. Caution is required when trusting conformal guarantees, as the assumption of exchangeability is often not met in practice, and some conformal methods catastrophically fail for slight deviations from the exchangeability assumption.
Even under the assumption of exchangeability, conformal guarantees have further weaknesses:
  1. 1.
    The conformal marginal coverage guarantee
    $\mathbb{P}[Y_{\text{new}}\in C(X_{\text{new}})]=\mathbb{E}_{\mathcal{D}_{\text{train}},\mathcal{D}_{\text{cal}}}[\mathbb{P}[Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{cal}}]]\geq 1-\alpha$
    does not imply that $\mathbb{P}[Y_{\text{new}}\in C(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{cal}}]\geq 1-\alpha$ for a fixed realization of the calibration set $\mathcal{D}_{\text{cal}}$, as already discussed before. If the calibration residuals are small by chance, conformal intervals may be too narrow, especially with tiny calibration datasets. Reliable calibration is generally unattainable with tiny calibration datasets: Even if the exchangeability assumption is satisfied, even methods with conformal guarantees often strongly undercover, i.e., $\mathbb{P}_{\mathcal{D}_{\text{train}},\mathcal{D}_{\text{cal}}}\left[\mathbb{P}[Y_{\text{new}}\in C_{\text{conformal}}(X_{\text{new}})|\mathcal{D}_{\text{train}},\mathcal{D}_{\text{cal}}]\ll 1-\alpha\right]\gg 0$. Therefore, caution is required when communicating conformal guarantees to practitioners.
  2. 2.
    Beyond marginal coverage, CLEAR is designed to improve conditional calibration: $\mathbb{P}[Y_{\text{new}}\in C(X_{\text{new}})|X_{\text{new}}]\approx 1-\alpha$. This is crucial in human-in-the-loop settings, where interventions are prioritized based on an accurate ranking of predictive uncertainty across data points (see Appendix I). Marginal coverage guarantees offer no guarantees for such rankings nor for conditional coverage. A method could have perfect marginal coverage but rank uncertainties arbitrarily. In other words, marginal coverage guarantees only address one specific metric (marginal coverage), while ignoring many other metrics that are often more important in practice.
To summarize, conformal marginal coverage guarantees (such as Lemma B.2) say very little about the overall quality of an uncertainty quantification method. Conformal marginal coverage guarantees only shed light on a very specific aspect of uncertainty quantification and only under the quite unrealistic assumption of exchangability.
B.2 Asymptotic conditional coverage for CLEAR: Lemma 2.1 assumptions
The assumptions underlying Lemma 2.1 are generally mild and often satisfied in practice. The assumption that $\Lambda$ is compact is particularly mild. In our experiments, we use a finite grid; for example, a combination of linearly spaced values from 0 to 0.09 and logarithmically spaced values from 0.1 to 100, resulting in approximately 4010 points. Intuitively, we believe that even if $\Lambda$ were unbounded, the result of Lemma 2.1 would still hold, since QuantileLoss is a strictly proper scoring rule and thus would not lead to excessively large values of $\lambda^{*}$ in the limit.
Lemma 2.1 also assumes consistency of the estimators. Many estimators satisfy this condition for both regression and quantile regression tasks:
  1. 1.
    Quantile Random Forests have been shown to consistently estimate quantiles under reasonable conditions, such as by regularizing the minimum number of samples per leaf [33]. Similarly, classical Random Forests are known to be consistent under standard assumptions [83]. Both QRF and Random Forests are available in our implementation of CLEAR.
  2. 2.
    Boosting methods trained on general loss functions can be made consistent under certain conditions, especially when regularized through early stopping [84]. This suggests that XGBoost and its quantile version, QXGB, can be consistent for suitable choices of hyperparameters. Both are included in our implementation of CLEAR.
  3. 3.
    More generally, regularized minimization of QuantileLoss over a sufficiently expressive function class on $\mathcal{D}_{\text{train}}$ yields consistent quantile estimators under broad assumptions [85].
In our experiments, we set $k=1$, so Lemma 2.1 requires only one of the base models to be consistent.
Finally, note that Lemma 2.1 implicitly assumes i.i.d. data, as described in Section 2.1. For real-world datasets, this i.i.d. assumption is often the most difficult to justify in practice and may pose a greater challenge than the other assumptions in the lemma.
B.3 Properties of the QuantileLoss
We define the $\text{QuantileLoss}(\mathcal{D},C):=\frac{1}{\left|\mathcal{D}\right|}\sum_{(x,y)\in\mathcal{D}}\left[QL_{\alpha/2}\big(y,l(x)\big)+QL_{1-\alpha/2}\big(y,u(x)\big)\right]/2$ (also denoted as “pinball loss”, “quantile loss”, “asymmetric piecewise linear loss”, “linlin loss”, “hinge loss”, “tick loss”, or “newsvendor loss” [86]), with $l(x),u(x)$ denoting the bounds of $C(x)$, and $QL_{\tau}(y,q)=(y-q)\big(\tau-\mathds{1}_{(-\infty,q]}(y)\big)$. Note that the majority of the literature defines the quantile loss as $QL_{\tau}$, whereas our QuantileLoss already aggregates $QL_{\tau}$ over both the upper and the lower bound of the intervals $C$ and over the data points in $\mathcal{D}$.
We use the QuantileLoss for three different purposes in this paper:
  1. 1.
    Some QR-methods in Step 2 of Algorithm 1 use the QuantileLoss to train their models.
  2. 2.
    In Step 4 of Algorithm 1 we minimize the QuantileLoss to determine $\lambda^{*}$. In other words, Step 3 and 4 of Algorithm 1 together (approximately) solve
    $(\gamma_{1}^{\star},\lambda^{\star})=\!\operatorname*{arg\,min}_{\begin{subarray}{c}(\gamma_{1},\lambda)\in(0,\infty)\times\Lambda\\ \text{s.t. }\left|\left\{(x,y)\in\mathcal{D}_{\text{cal}}:y\in C_{\gamma_{1},\lambda}(x)\right\}\right|\geq\left\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\right\rceil\end{subarray}}\!\text{QuantileLoss}(\mathcal{D}_{\text{val}},C_{\gamma_{1},\lambda}),$ (4)
    where $C_{\gamma_{1},\lambda}=\left[\hat{f}-\gamma_{1}\hat{q}_{\alpha/2}^{\text{ale}}-\lambda\gamma_{1}\hat{q}_{\alpha/2}^{\text{epi}},\,\hat{f}+\gamma_{1}\hat{q}_{1-\alpha/2}^{\text{ale}}+\lambda\gamma_{1}\hat{q}_{1-\alpha/2}^{\text{epi}}\right]$.
  3. 3.
    We use the QuantileLoss as an evaluation metric on the test dataset $\mathcal{D}_{\text{test}}$.
In the following, we will discuss multiple favorable properties of the QuantileLoss (many of these properties also hold for the Interval Score Loss).
B.3.1 Intuition Behind the QuantileLoss
In many real-world applications, the severity of a prediction error often depends on the magnitude of the deviation from the predictive interval. For example, in financial portfolio management, a massive drop below the predicted lower bound can be significantly more damaging than a slight deviation. Similarly, for flood protection systems where dam heights are based on an upper predictive bound, a large amount of overflow causes substantially more damage than a small overflow. Traditional metrics like PICP treat coverage as a binary outcome, distinguishing only between points that are covered and those that are not. The QuantileLoss, however, offers a more nuanced evaluation by penalizing the magnitude of a data point’s distance from the predictive interval when it is not covered.
Intuitively, the QuantileLoss strongly penalizes the distance to the predictive intervals for data points outside the predictive interval, and gently penalizes the width of the predictive intervals that do not cover the data points. This incentivizes the predictive intervals to widen in regions of large uncertainty and to adaptively narrow down in regions of low uncertainty, incentivizing good relative uncertainty (see Appendix I). Simultaneously, the QuantileLoss incentivizes good absolute uncertainty (i.e., marginal calibration, see Appendix I): If the proportion of data points below the upper bounds is less than $1-\frac{\alpha}{2}$, then increasing the upper bounds by a constant improves the QuantileLoss until the proportion reaches $1-\frac{\alpha}{2}$, i.e., for all $c>0$,
$\left|\left\{(x,y)\in\mathcal{D}_{\text{cal}}:y\leq u(x)+c\right\}\right|<(1-\frac{\alpha}{2})|\mathcal{D}_{\text{cal}}|$
$\implies\text{QuantileLoss}(\mathcal{D},[l,u])>\text{QuantileLoss}(\mathcal{D},[l,u+c]),$
and vice versa if the proportion of data points below the upper bounds is more than $1-\frac{\alpha}{2}$, i.e., for all $c>0$,
$\left|\left\{(x,y)\in\mathcal{D}_{\text{cal}}:y\leq u(x)\right\}\right|>(1-\frac{\alpha}{2})|\mathcal{D}_{\text{cal}}|$
$\implies\text{QuantileLoss}(\mathcal{D},[l,u])<\text{QuantileLoss}(\mathcal{D},[l,u+c]),$
and analogously, the QuantileLoss incentivizes the lower bound to be above $\frac{\alpha}{2}|\mathcal{D}_{\text{cal}}|$ data points and below $(1-\frac{\alpha}{2})|\mathcal{D}_{\text{cal}}|$ data points. The QuantileLoss simultaneously evaluates multiple different properties of predictive intervals and cannot be easily tricked, in contrast to other metrics: PCIP can be easily maximized by infinitely wide intervals, which are completely useless in practice; NIW can be minimized by zero-width intervals, which are useless in practice; NCIW only measures the relative uncertainty while completely ignoring the marginal coverage.
B.3.2 The QuantileLoss Measures Input-Conditional Calibration
In the following, we will mathematically argue why the QuantileLoss measures input-conditional calibration. Within this paper, we denote input conditional calibration $\mathbb{P}[Y_{\text{new}}\in C(X_{\text{new}})|X_{\text{new}}]=1-\alpha$ simply as conditional calibration.
The true input conditional quantiles minimize the expected QuantileLoss, i.e.,
$(q_{\alpha/2},q_{1-\alpha/2})\in\operatorname*{arg\,min}_{(l,u)\in\mathcal{Y}^{\mathcal{X}}\times\mathcal{Y}^{\mathcal{X}}}\mathbb{E}_{(X_{\text{new}},Y_{\text{new}})}\left[\text{QuantileLoss}\big(\{(X_{\text{new}},Y_{\text{new}})\},[l,u]\big)\right].$ (5)
Any minimizer of the QuantileLoss satisfies input-conditional coverage almost surely (a.s.) [^4] A statement holds almost surely if it holds a probability of $100\%$. E.g., a standard normally distributed random variable $X\sim\mathcal{N}(0,1)$ is a.s. not exactly equal to $\sqrt{2}$, i.e., $X\overset{\text{a.s.}}{\neq}\sqrt{2}$., i.e., any solution $(l^{*},u^{*})$ of the minimization problem (5) satisfies that $l^{*}(X)$ is an input-conditional $\alpha/2$-quantile a.s. and $u^{*}(X)$ is an input-conditional $1-\alpha/2$-quantile a.s., thus, $\mathbb{P}[Y_{\text{new}}\in[l(X_{\text{new}}),u(X_{\text{new}})]|X_{\text{new}}]\overset{\text{a.s.}}{\geq}1-\alpha$. In other words, any deviation from the true quantiles gets penalized by the expected QuantileLoss, as it is a strictly proper scoring rule [87]. If the true conditional CDF is a.s. continuous, then any solution $(l^{*},u^{*})$ of (5) satisfies input-conditional calibration $\mathbb{P}[Y_{\text{new}}\in[l(X_{\text{new}}),u(X_{\text{new}})]|X_{\text{new}}]\overset{\text{a.s.}}{=}1-\alpha$. In other words, any deviation from input-conditional coverage gets penalized. Applying the QuantileLoss on a finite dataset $\mathcal{D}$ can be seen as a Monte-Carlo approximation of the expected QuantileLoss.
Another common evaluation method is to compare the average interval width NIW (or NCIW) among methods that approximately obtain the targeted marginal coverage. However, this evaluation method can be exploited: Even if you perfectly know the true distribution, reporting intervals that do not satisfy input-conditional coverage would be optimal. This evaluation method prefers intervals that over-cover in regions with low uncertainty and under-cover in regions of high uncertainty, as demonstrated in the following examples (Examples B.3 and B.4 are easier to derive, but Examples B.5 and B.6 are slightly more insightful).
Example B.3.
Let $X\in\{1,2\}$, with $\mathbb{P}[X=1]=0.5$ and $\mathbb{P}[X=2]=0.5$. Let $Y|X=1\sim\mathcal{U}(-1,1)$ and $Y|X=2\sim\mathcal{U}(-2,2)$. For a target coverage of $1-\alpha=0.9$, the true conditional intervals are:
  • [0.9,0.9][-0.9,0.9] when $X=1$ (width=1.8, coverage=0.9)
  • [1.8,1.8][-1.8,1.8] when $X=2$ (width=3.6, coverage=0.9)
The average width of the true intervals is $\mathbb{E}[\text{width}]=0.5\times 2\cdot 0.9+0.5\times 2\cdot 1.8=2.7$. The marginal coverage is exactly 0.9. Now, consider an alternative method that sacrifices conditional calibration to minimize average width. This method could report the following intervals:
  • [1,1][-1,1] when $X=1$ (width=2, coverage=1)
  • [1.6,1.6][-1.6,1.6] when $X=2$ (width=3.2, coverage=0.8)
The marginal coverage of this method is $\mathbb{P}[\text{covered}]=0.5\times 1+0.5\times 0.8=0.9$. It still achieves the target marginal coverage of $90\%$, but its average width is $\mathbb{E}[\text{width}]=0.5\times 2+0.5\times 2\cdot 1.6=2.6$. Since $2.6<2.7$, this method would be preferred, demonstrating that NIW can incentivize deviations from input-conditional calibration. ∎
Example B.4.
Under the setting of Example B.3, the situation gets even more extreme if we change to $\alpha=0.5$: For a target coverage of $1-\alpha=0.5$, the true conditional intervals are:
  • [0.5,0.5][-0.5,0.5] when $X=1$ (width=1, coverage=0.5)
  • [1,1][-1,1] when $X=2$ (width=2, coverage=0.5)
The average width of the true intervals is $\mathbb{E}[\text{width}]=0.5\times 1+0.5\times 2=1.5$. The marginal coverage is exactly $0.5$. Now, consider an alternative method that sacrifices conditional calibration to minimize average width. This method could report the following intervals:
  • [1,1][-1,1] when $X=1$ (width=2, coverage=1)
  • $[127,127]$ when $X=2$ (width=0, coverage=0)
The marginal coverage of this untruthful method is $\mathbb{P}[\text{covered}]=0.5\times 1+0.5\times 0=0.5$. It still achieves the target marginal coverage, but its average width is $\mathbb{E}[\text{width}]=0.5\times 2+0.5\times 0=1$. Since $15<1.5$, this untruthful method would be preferred, demonstrating that NIW can incentivize strong deviations from input-conditional calibration. Here, the interval that minimizes NIW (and NCIW) under the constraint of maintaining marginal coverage, outputs a much wider interval for $X=1$ than for $X=2$, while there is obviously more uncertainty for $X=2$ than for $X=1$. ∎
Example B.5.
Let $\mathbb{P}[X=1]=0.5=\mathbb{P}[X=2]$, and $Y|X=x\sim\mathcal{N}(\mu=0,\sigma=x)$. Then for $\alpha=5\%$, the true conditional quintiles would be $q_{0.975}(x)=\Phi^{-1}(0.975)x\approx 1.96x$ and $q_{0.025}(x)=\Phi^{-1}(0.025)x\approx-1.96x$, thus $q_{0.025},q_{0.975}]$ exactly satisfy input-conditional calibration $\mathbb{P}\left[Y_{\text{new}}\in[q_{0.025}(X_{\text{new}}),q_{0.975}(X_{\text{new}})]\mid X_{\text{new}}\right]\overset{\text{a.s.}}{=}95\%$. However, the on average narrowest intervals satisfying marginal coverage would be approximately
  • [2.16357,2.16357][-2.16357,2.16357] when $X=1$ ($\text{width}=2\cdot 2.16357$, $\text{coverage}\approx 0.969502$)
  • $1.81514\cdot 2,1.81514\cdot 2]$ when $X=2$ ($\text{width}=2\cdot 1.81514\cdot 2$, $\text{coverage}\approx 0.930498$).
These intervals fail input-conditional coverage while maintaining marginal coverage and result in a narrower average width $5.79385<5.88$ than the true intervals $q_{0.025},q_{0.975}]$. ∎
Example B.6.
Let $\mathbb{P}[X=1]=\frac{9}{19}=\mathbb{P}[X=2]$, $\mathbb{P}[X=11]=\frac{1}{19}$, and $Y|X=x\sim\mathcal{N}(\mu=0,\sigma=x)$. Then for $\alpha=10\%$, the true conditional quintiles would be $q_{0.9}(x)=\Phi^{-1}(0.95)x\approx 1.645x$ and $q_{0.05}(x)=\Phi^{-1}(0.05)x\approx-1.645x$, thus $q_{0.5},q_{0.95}]$ exactly satisfy input-conditional calibration $\mathbb{P}\left[Y_{\text{new}}\in[q_{0.05}(X_{\text{new}}),q_{0.95}(X_{\text{new}})]\mid X_{\text{new}}\right]\overset{\text{a.s.}}{=}90\%$. However, the on average narrowest intervals satisfying marginal coverage would be approximately
  • [2.16357,2.16357][-2.16357,2.16357] when $X=1$ ($\text{width}=2\cdot 2.16357$, $\text{coverage}\approx 0.969502$)
  • $1.81514\cdot 2,1.81514\cdot 2]$ when $X=2$ ($\text{width}=2\cdot 1.81514\cdot 2$, $\text{coverage}\approx 0.930498$).
  • $[12345,12345]$ when $X=11$ ($\text{width}=0$, $\text{coverage}=0$).
These intervals fail input-conditional coverage while maintaining marginal coverage and result in a narrower average width $5.48891<6.705263$ than the true intervals $q_{0.5},q_{0.95}]$. ∎
Interpretation of Examples B.3, B.4, B.5 and B.6.
While QuantileLoss is only minimized for intervals consisting of the true quantiles $q_{\frac{\alpha}{2}},q_{1-\frac{\alpha}{2}}]$, the Examples B.3, B.4, B.5 and B.6 showed that metrics like PCIP, NIW, NCIW, and their combinations can be misaligned with input-conditional coverage. This misalignment can be very severe in Examples B.4 and B.6, whereas it is less severe in Examples B.3 and B.5. Intuitively, the malalignment can be particularly extreme for large values of $\alpha$ (as in Example B.3) or for inputs that are much more uncertain than the average uncertainty (as for $x=11$ in Example B.5). In our experiments on real-world data in Sections 4.2, 4.3 and F, compared to these extreme examples, the different metrics are more aligned, in the sense that CLEAR shows superior performance across all considered metrics. Reasons for this empirically observed alignment could be that 1st none of the compared methods actively tries to exploit the weaknesses of PCIP, NIW, NCIW; 2nd in our experiments, we use only small values of $\alpha=5\%$ (and $\alpha=10\%$); 3rd there might not be many regions of extremely high uncertainty in these dataset.
Now we have concluded that under the QuantileLoss, it is optimal always to report the true quantiles if the true underlying distributions are known. In the next subsection, we will discuss how this translates to situations where the true distributions are not known.
B.3.3 The QuantileLoss Incentivizes Truthfulness Even Under Incomplete Information (From a Bayesian Point of View)
Proper scoring rules are mathematically guaranteed to incentivize reporting the true distribution, but their application to uncertainty quantification requires care in interpreting what kind of uncertainty they incentivize. [88] show that some proper scoring rules impose asymmetric penalties for over- versus under-estimation. While this may appear to induce bias, we argue that such asymmetry faithfully captures epistemic uncertainty rather than distorting the estimate of total predictive uncertainty.
In our case, the relevant scoring rule is the QuantileLoss. For a given quantile, e.g., $q_{0.975}$, the loss penalizes underestimation more heavily than overestimation. This asymmetry is not a flaw; rather, it incentivizes appropriately wider predictive intervals in the presence of epistemic uncertainty. Specifically, this asymmetry encourages intervals to widen more in regions of high epistemic uncertainty than in regions of low epistemic uncertainty.
This phenomenon can be described more quantitatively from a Bayesian perspective. The posterior predictive distribution,
$\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\mathcal{D},\pi]=\mathbb{E}\left[\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\theta]\mid\mathcal{D},\pi\vphantom{\big|}\right]$
optimally reflects both aleatoric uncertainty $\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\theta]$ (e.g., noise) and epistemic uncertainty $\mathbb{P}[\theta\mid\mathcal{D},\pi]$ (e.g., over parameters) given a prior $\pi$. [^5] Mathematicians (with a background in measure-theory) should read $\mathbb{P}[Y\mid X]$ as $\mathbb{P}[Y\in(\cdot)\mid X]$. Marginalizing over the posterior naturally yields a wider distribution than relying on a single point estimate, with a greater widening effect in regions of high epistemic uncertainty. Let $q_{\tau,P}$ denote the $\tau$-quantile of a distribution $P$. The posterior predictive quantiles $(q_{\alpha/2,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\mathcal{D},\pi]},q_{1-\alpha/2,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\mathcal{D},\pi]})$ minimize the expected quantile loss:
$\mathbb{E}\left[\text{QuantileLoss}\big(\{(X_{\text{new}},Y_{\text{new}})\},[l,u]\big)\mid\mathcal{D},\pi\right].$
Since the posterior predictive distribution includes epistemic uncertainty, its quantiles also appropriately include this component. This stands in contrast to other estimators, such as $q_{1-\alpha/2,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\hat{\theta}]}$, $\mathbb{E}\left[q_{1-\alpha/2,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\theta]}\mid\mathcal{D},\pi\right]$, or $\text{Median}\left[q_{1-\alpha/2,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\theta]}\mid\mathcal{D},\pi\right]$ (similar to how we estimate the pure aleatoric uncertainty in Section 2.3), which ignore epistemic uncertainty.
Example B.7.
Let $Y|X=x\sim\mathcal{N}(\theta(x),\sigma=1)$ with an unknown mean $\theta(x)\in\mathbb{R}$ and a known standard deviation of 1. Further, assume that the posterior distribution of $\theta(x)$ is $\mathcal{N}(0,\sigma=10)$ due to large epistemic uncertainty. Then, the predictive posterior distribution of $Y|X=x$ is $\mathcal{N}\left(0,\sigma=\sqrt{10^{2}+1}\right)$, which includes both aleatoric and epistemic uncertainty.
  • $q_{0.975,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\mathcal{D},\pi]})=q_{0.975,\mathcal{N}\left(0,\sigma=\sqrt{10^{2}+1}\right)}\approx 1.96\cdot\sqrt{10^{2}+1}$ appropriately takes epistemic uncertainty into account.
  • $q_{0.975,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\hat{\theta}]}=q_{0.975,\mathcal{N}\left(\hat{\theta}(x),\sigma=1\right)}\approx\hat{\theta}(x)+1.96$ ignores the large epistemic uncertainty (with an interval width of $2\cdot 1.96$ as $q_{0.025,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\hat{\theta}]}\approx\hat{\theta}(x)-1.96$).
  • $\mathbb{E}\left[q_{0.975,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\theta]}\mid\mathcal{D},\pi\right]=\mathbb{E}\left[q_{0.975,\mathcal{N}\left(\theta(x),\sigma=1\right)}\mid\mathcal{D},\pi\right]\approx\mathbb{E}\left[\theta(x)+1.96\mid\mathcal{D},\pi\right]=1.96$ ignores the large epistemic uncertainty.
  • $\text{Median}\left[q_{0.975,\mathbb{P}[Y_{\text{new}}\mid X_{\text{new}},\theta]}\mid\mathcal{D},\pi\right]=\text{Median}\left[q_{0.975,\mathcal{N}\left(\theta(x),\sigma=1\right)}\mid\mathcal{D},\pi\right]\approx\text{Median}\left[\theta(x)+1.96\mid\mathcal{D},\pi\right]=1.96$ ignores the large epistemic uncertainty.
We fully agree with [88], that minimizing the QuantileLoss leads to a biased estimator of aleatoric uncertainty alone (as in Example B.7, the aleatoric uncertainty corresponds to an interval width of $2\cdot 1.96$). However, minimizing the QuantileLoss yields a principled estimator for total predictive uncertainty, as it is optimal from a Bayesian point of view, taking epistemic uncertainty into account. An unbiased estimator for total predictive uncertainty must also incorporate the epistemic uncertainty and is therefore naturally biased for estimating aleatoric uncertainty.
Thus, the asymmetry of the QuantileLoss is precisely what makes it appropriate for evaluating models that estimate total predictive uncertainty. This justifies its use both in Step 4 of our Algorithm 1 and as an evaluation metric on the unseen test dataset $\mathcal{D}_{\text{test}}$.
B.4 Intuitive Theoretical Motivation of CLEAR
PCS achieves empirically good results and nicely captures common sense, with limited theoretical guarantees, while CQR satisfies theoretical guarantees, but misses crucial common sense. In our work, we combine practical experience, theoretical insights, common sense, and empirical results, always with the goal in mind to obtain a useful, verdicial, reliable, stable, high-performing method for practical real-world applications.
Note that the original version of PCS did not have any theoretical guarantees for uncertainty quantification [40], and [30] mentioned a modified version of PCS-UQ that satisfies conformal marginal coverage guarantees. However, no version of PCS offered a theoretical guarantee for asymptotic input-conditional coverage.
Via CLEAR, we are the first to propose an extension of PCS-UQ that provably satisfies asymptotic input-conditional coverage guarantees (see Lemma 2.1). All previous versions of PCS, did not only lack a proof for input-conditional coverage guarantees, but actually do not satisfy input-conditional coverage guarantees. Pure PCS-UQ has the systematic bias of under-coverage in regions with many observations and large noise. This was a central motivation for CLEAR. While this theoretical bias is quite obvious from a theoretical common sense point of view, we conducted our large-scale experimental evaluation to see that this bias can actually have a significant impact on the performance, which can be substantially mitigated by CLEAR.
On the other hand, CQR (or QR) has major problems based on common sense. If one looks deep into the proof of the asymptotic input-conditional coverage of CQR (or QR), one can see that the main idea of the proof is that, asymptotically, you will have infinitely many training data points around every point $x$, resulting in accurate quantile predictions. However, for finitely many data points, there will always be regions in the input space with too few data points. In this region (C)QR cannot accurately estimate the quantiles. However, (C)QR will sometimes output very narrow intervals in these regions, which are absolutely not trustworthy and will undercover substantially in these regions. We can see this phenomenon across all different synthetic data-generating processes that we’ve tried (see Sections 3.1, 4.1 and C). When we sample $X\sim\mathcal{N}(0,I_{d})$ standard Gaussian, we can see that even for $n=5000$ training data points $\mathbb{P}\left[Y\in C_{\text{CQR}}\mid\|X\|_{2}>4\right]\ll 1-\alpha$. Based on common sense it is strongly expected that for $\forall n\in\mathbb{N}:\exists r(n):\mathbb{P}\left[Y\in C_{\text{CQR}}\mid\|X\|_{2}>r(n)\right]\ll 1-\alpha$, as the intervals of (C)QR do not have any reason to become wider out-of-sample while their predictions become more unreliable the further you move away from the training data.
Hypothesis B.8.
We hypothesize that under some fairly general and mild assumptions, for every “non-trivial” data-generating process,
$\exists c>0:\forall n\in\mathbb{N}:\exists R(n)\in(0,\infty):\forall r>R(n)\mathbb{P}\left[Y\in C_{\text{CQR}}\mid\|X\|_{2}>r\right]<1-\alpha-c$
holds. We think that this claim strongly agrees with basic intuition, and all our empirical results strongly support this hypothesis. However, we leave the proof for future work. The main technical challenge is probably to define “non-trivial” appropriately.
This does not contradict asymptotic input-conditional coverage guarantees as $\lim_{n\to\infty}r(n)=\infty$. However, for any finite sample size $n$, this is a huge problem both from a common-sense perspective and from what you can see from our empirical results. In other words, even in regions where the ensemble members heavily disagree on the predictions (usually in regions with few training data), (C)QR would be way too over-confident. CQR tries to compensate for this overconfidence in the calibration step, which results in CQR over-covering in regions with many training data points and under-covering in regions with few training data points (see Figures 1, 2, 4, 5 and 6).
This is the reason why PCS-UQ and CQR form a wonderful symbiosis. In regions with many training data points, the ensemble predictions coming from different bootstraps of the training data will be close to each other. PCS-UQ has a systematic basis to under-cover in regions with many training data points and to overcover in regions with few training data points, whereas for CQR, it is exactly the other way around. If you apply any monotonic transformation of the uncertainty of PCS (multiplicative (conformal) calibration, additive (conformal) calibration, or the calibration methods suggested by [3, 70, 71, 72]), you will always have the systematic bias that any monotonically calibrated version of PCS-UQ will under-cover in regions with many training data points and will over-cover in regions with few training data points, whereas it is exactly the other way around for any monotonically calibrated version of CQR. By combining the two of them, we can mitigate this bias (see Figures 1, 2, 4, 5 and 6). This aligns very well with classical statistical results, telling us that for predictive intervals, one needs to combine the estimated noise structure with epistemic uncertainty on the parameters (confidence intervals).
B.5 CLEAR Algorithm: Details
Algorithm 2 Fast Conformal Implementation of Step 3 from Algorithm 1
1:Input: Data $(X_{i},Y_{i})$ for $i=1,\dots,n$, split into training $\mathcal{D}_{\text{train}}$, calibration $\mathcal{D}_{\text{cal}}$, and validation $\mathcal{D}_{\text{val}}$ (we consider $\mathcal{D}_{\text{cal}}=\mathcal{D}_{\text{val}}$); grid of $\lambda$ values $\Lambda$; significance level $\alpha$.
2:Step 3: Define prediction intervals for each $\lambda\in\Lambda$.
3: First, define a preliminary (non-calibrated) interval:
$\tilde{C}_{\lambda}=\left[\hat{f}-\hat{q}_{\alpha/2}^{\text{ale}}-\lambda\hat{q}_{\alpha/2}^{\text{epi}},\,\hat{f}+\hat{q}_{1-\alpha/2}^{\text{ale}}+\lambda\hat{q}_{1-\alpha/2}^{\text{epi}}\right]$
Then, compute conformity scores $S_{i}^{\lambda}=\max\left\{\frac{\tilde{l}_{\lambda}(X_{i})-Y_{i}}{\hat{f}(X_{i})-\tilde{l}_{\lambda}(X_{i})},\frac{Y_{i}-\tilde{u}_{\lambda}(X_{i})}{\tilde{u}_{\lambda}(x_{i})-\hat{f}(X_{i})}\right\}$, where $\tilde{l}_{\lambda}(x),\tilde{u}_{\lambda}(x)$ are the lower and upper bounds of $\tilde{C}_{\lambda}(x)$. Let $\gamma_{1}$ be the $\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\rceil$-th smallest score among $\{S_{i}^{\lambda}\}$ (if $\lceil(1-\alpha)(|\mathcal{D}_{\text{cal}}|+1)\rceil>|\mathcal{D}_{\text{cal}}|$, take the largest score). Define the calibrated interval:
$C_{\lambda}=\left[\hat{f}-\gamma_{1}\hat{q}_{\alpha/2}^{\text{ale}}-\lambda\gamma_{1}\hat{q}_{\alpha/2}^{\text{epi}},\,\hat{f}+\gamma_{1}\hat{q}_{1-\alpha/2}^{\text{ale}}+\lambda\gamma_{1}\hat{q}_{1-\alpha/2}^{\text{epi}}\right]$
4:Output: calibrated prediction intervals $C_{\lambda}$ for each $\lambda$ from the grid $\Lambda$.
Appendix C Simulations: Details
For each simulation run, the coefficients $\beta_{1},\dots,\beta_{d}$ are drawn independently from a Gaussian distribution $\mathcal{N}(1,0.5^{2})$. The mean function $\mu(X)$ is then defined as $\mu(X)=5.0+\sum_{i=1}^{d}(-1)^{i+1}\beta_{i}|X_{i}|^{e_{i}}$, where the exponent $e_{i}=1.5$ if $i$ is odd, and $e_{i}=1.25$ if $i$ is even.
C.1 Heteroskedastic Case
In  Section 3.1, we focused on the univariate homoskedastic setting ($d=1$ and $\sigma(x)=1$). Here, we briefly report results for heteroskedastic data, which show similar patterns. Figures 4 and 5 illustrate the conditional coverage and interval width of the algorithms under two heteroskedastic noise structures: $\sigma_{2}(x)=1+|x|$ and $\sigma_{3}(x)=1+\frac{1}{1+x^{2}}$. As mentioned, the results are analogous to the homoskedastic setting in  Section 3.1.
C.2 Multivariate Case
C.2.1 Test point generation
Following setup in [89], let $r_{1}<r_{2}<\dots<r_{K}$ denote a set of predetermined distances. For each radius $r_{k}$, we sample points uniformly from the surface of the unit $d$-dimensional sphere as follows. First, we draw a vector $\mathbf{v}\in\mathbb{R}^{d}$ whose entries are independent standard normal random variables. We then normalize $\mathbf{v}$ to obtain a unit vector $\mathbf{u}=\mathbf{v}/\|\mathbf{v}\|_{2}$. Finally, we scale $\mathbf{u}$ by $r_{k}$ to obtain the test point $\mathbf{x}=r_{k}\cdot\mathbf{u}$. In the one-dimensional case, this procedure reduces to selecting $x=r_{k}$ or $x=-r_{k}$ with equal probability. This mechanism ensures that, for each $r_{k}$, the generated points are uniformly distributed on the surface of the sphere of radius $r_{k}$, thereby allowing a precise evaluation of prediction intervals as a function of distance from the origin. Finally, we generate $Y$ with the same data-generating mechanism as in the train set.
C.2.2 Results
Figure 6 illustrates the conditional coverage and interval width of the evaluated algorithms in the multivariate setting, where $X$ is drawn from an independent multivariate Gaussian distribution and $Y=\mu(X)+\varepsilon$, with $\varepsilon\sim\mathcal{N}(0,1)$. The regression function $\mu(X)$, as previously defined, involves a sum of randomly weighted power transformations of the absolute values of the input features.
Consistent with the univariate homoskedastic results in Section 3.1, Figure 6 shows that both CQR and naive conformal prediction achieve reliable coverage in high-density regions but tend to under-cover in low-density or extrapolation areas. In contrast, CLEAR maintains valid conditional coverage across the entire input space by appropriately adjusting interval widths.
Figure 4: Univariate conditional coverage and average width of the prediction intervals for a heteroskedastic case where $\sigma_{2}(x)=1+|x|$.
Figure 5: Univariate conditional coverage and average width of the prediction intervals for a heteroskedastic case where $\sigma_{3}(x)=1+\frac{1}{1+x^{2}}$.
Figure 6: Multivariate conditional coverage and average width (distance from the origin is shown) of the prediction intervals for a homoskedastic case where $d\in\{1,2,3,20\}$ was randomly selected.
Appendix D Experimental Setup: Details
D.1 Real-World Data
To evaluate our method in real-world scenarios, we apply it to 17 publicly available regression datasets curated by [30] that includes various domains such as housing, energy, materials science, and healthcare [90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108]. Description of the datasets can be found in Table 11.
D.2 Baselines for Variants (a), (b), (c)
The following baselines provide further context and comparisons, omitted from the main paper for brevity. All compared methods (except UACQR) use the same median $\hat{f}(x)=\hat{f}_{\text{PCS}}(x)$ obtained from PCS. The results from Section F.1 use PCS variant (a), Section F.2 uses PCS variant (b), and Section F.3 uses PCS variant (c). This allows us to isolate the effect of different UQ methods.
  1. 1.
    Conformalized PCS median (Naive): For the PCS ensemble, we compute the median prediction $\hat{f}(x)=\hat{f}_{\text{PCS}}(x)$ for each $x$. On the calibration set, absolute residuals $a_{i}=|y_{i}-\hat{f}(x_{i})|$ are computed. The $(1-\alpha)$-th quantile $\gamma_{\text{naive}}$ of these absolute residuals is then used to define the interval:
    $C_{\text{Naive}}(x)=[\hat{f}(x)-\gamma_{\text{naive}},\;\hat{f}(x)+\gamma_{\text{naive}}]$
    This method applies a constant, symmetric width adjustment to the point predictions. The result of this baseline has been presented for the simulations. We omit it from the real-world data for brevity. However, all the baselines can be found in the supplementary code.
  2. 2.
    CLEAR with fixed $\lambda=1$: This is a variant of the main CLEAR methodology where the ratio $\lambda=\gamma_{2}/\gamma_{1}$ is fixed to 1. This implies $\gamma_{1}=\gamma_{2}$. The prediction interval, based on Equation (3), becomes:
    $C_{\lambda=1}(x)=\left[\hat{f}(x)-\gamma_{1}\left(\hat{q}_{\alpha/2}^{\text{ale}}(x)+\hat{q}_{\alpha/2}^{\text{epi}}(x)\right),\quad\hat{f}(x)+\gamma_{1}\left(\hat{q}_{1-\alpha/2}^{\text{ale}}(x)+\hat{q}_{1-\alpha/2}^{\text{epi}}(x)\right)\right].$
    In this configuration, the single parameter $\gamma_{1}$ is calibrated using the standard split conformal procedure on the validation set. It effectively learns a single scaling factor for the sum of the pre-calibrated aleatoric and epistemic uncertainty widths, without adjusting their ratio.
  3. 3.
    CLEAR with fixed $\gamma_{1}=1$: Another variant of CLEAR where $\gamma_{1}$ is fixed to 1. With $\gamma_{1}=1$, then $\gamma_{2}=\lambda$ and the prediction interval from Equation (3) reads as:
    $C_{\gamma_{1}=1}(x)=\left[\hat{f}(x)-\hat{q}_{\alpha/2}^{\text{ale}}(x)-\lambda\hat{q}_{\alpha/2}^{\text{epi}}(x),\quad\hat{f}(x)+\hat{q}_{1-\alpha/2}^{\text{ale}}(x)+\lambda\hat{q}_{1-\alpha/2}^{\text{epi}}(x)\right].$
    Here, $\lambda$ (or equivalently $\gamma_{2}$) is the parameter calibrated on the calibration set through a coverage-based adjustment. This approach fixes the contribution of the (pre-calibrated) aleatoric uncertainty component and adaptively scales the epistemic uncertainty component.
  4. 4.
    UACQR-S and UACQR-P For variant (c), instead of computing $C_{\gamma_{1}=1}(x)$ and $C_{\lambda=1}(x)$ with CLEAR, we directly use the implementation from [19] to assess our performance against this alternative. Since this is only relevant for variant (c), we use the exact same configuration as the aleatoric part for QRF (see Section D.4 and Table 4).
D.3 Deep Ensembles and Simultaneous Quantile Regression Implementation
To demonstrate CLEAR’s versatility beyond tree-based methods, we implement state-of-the-art deep learning approaches for uncertainty quantification. These neural network-based methods offer complementary strengths to our primary PCS framework, particularly in capturing complex nonlinear relationships and heteroskedastic noise patterns.
D.3.1 Deep Ensembles for Epistemic Uncertainty
Our deep ensemble implementation follows [22] with several enhancements for improved diversity and calibration. Each ensemble consists of $M=5$ neural networks (adaptive based on dataset size), with the following specifications:
  • Architecture: Each network employs a fully-connected architecture with hidden layers of sizes (256, 128), ReLU activations, batch normalization, and dropout (rate=0.1). Skip connections are incorporated to stabilize training.
  • Diversity strategies: To promote ensemble diversity, we employ: (i) different random initializations, (ii) bootstrap sampling where each member trains on different subsamples, (iii) varied learning rates, (iv) different learning rate schedules, and (v) small input noise augmentation.
  • Training: Each network is trained for up to 1500 epochs using Adam optimizer with learning rate $10^{-3}$, early stopping (patience=50), weight decay ($10^{-5}$), and batch size 64.
  • Calibration: Following training, we apply a multiplicative calibration factor computed on the validation set using a grid search over c-values to ensure the ensemble achieves target coverage.
The epistemic uncertainty is quantified through the empirical quantiles of the ensemble predictions: $\hat{q}^{\text{epi}}_{\alpha/2}(x)=\hat{f}(x)-Q_{\alpha/2}(\{\hat{f}_{m}(x)\}_{m=1}^{M})$ and $\hat{q}^{\text{epi}}_{1-\alpha/2}(x)=Q_{1-\alpha/2}(\{\hat{f}_{m}(x)\}_{m=1}^{M})-\hat{f}(x)$, where $\hat{f}(x)$ is the ensemble median and $Q_{\tau}$ denotes the empirical quantile function.
D.3.2 Simultaneous Quantile Regression for Aleatoric Uncertainty
For aleatoric uncertainty, we implement simultaneous quantile regression (SQR) following [13], which directly models multiple conditional quantiles through a single neural network with specialized architecture:
  • Architecture: The network uses hidden layers of sizes (256, 256, 128) with LeakyReLU activations (negative slope=0.01), layer normalization, and dropout (rate=0.2). The output layer produces three values corresponding to the $\alpha/2$, 0.5, and $1-\alpha/2$ quantiles.
  • Loss function: We minimize a combined loss consisting of: (i) the pinball loss for each quantile, and (ii) a crossing penalty term that encourages monotonicity of quantiles: $\mathcal{L}_{\text{crossing}}=\text{ReLU}(q_{\alpha/2}-q_{0.5})+\text{ReLU}(q_{0.5}-q_{1-\alpha/2})$.
  • Training: The model is trained for up to 3000 epochs using Adam optimizer with learning rate $5\times 10^{-4}$, cosine annealing schedule, and early stopping (patience=200). Gradient clipping (max norm=1.0) prevents training instabilities.
  • Ensemble averaging: To improve stability, we train an ensemble of SQR models with different random seeds and average their quantile predictions.
The aleatoric uncertainty estimates are computed as $\hat{q}^{\text{ale}}_{\alpha/2}(x)=q_{0.5}(x)-q_{\alpha/2}(x)$ and $\hat{q}^{\text{ale}}_{1-\alpha/2}(x)=q_{1-\alpha/2}(x)-q_{0.5}(x)$, ensuring consistency with the median prediction.
D.3.3 Integration with CLEAR
The DE and SQR components are integrated into CLEAR using the same calibration procedure as our PCS-UQ (model selection and tree-based methods). The epistemic estimates from DE and aleatoric estimates from SQR are combined using the optimized parameters $\lambda$ and $\gamma_{1}$ according to Equation (3). The conformal calibration step ensures marginal coverage while the adaptive $\lambda$ selection balances the relative contributions of epistemic and aleatoric uncertainties, accounting for potential scale differences between neural network and tree-based estimators.
D.4 PCS Implementation: Details
This section outlines the specific implementation details for generating the PCS ensembles, which provide the point predictor $\hat{f}$ and the raw epistemic uncertainty estimates $\hat{q}^{\text{epi}}$ used as input for the CLEAR method (Section 2.4), and also form the basis for the standalone calibrated PCS baseline intervals. We explicitly opted for experimenting with three variants of CLEAR to assess the framework’s performance and robustness across different modeling choices for epistemic and aleatoric uncertainty components, as discussed in the main text. The core methodology, involving model selection and model perturbations via bootstrapping to capture epistemic uncertainty, follows the principles described in Section 2.
The process begins with data partitioning and bootstrapping. For each dataset and unique random seed, the data is divided into training (60%), validation (20%), and test (20%) sets. Subsequently, $n_{boot}=100$ bootstrap resamples are drawn from this designated training set to construct the PCS ensemble. Then, for our variants, two distinct pools of base models were developed to generate these PCS ensembles, catering to the different CLEAR variants. Table 4 details the quantile estimators used for CLEAR variant (a). Variant (b) only uses QXGB from Table 4. Table 5 describes the mean estimators utilized for CLEAR variant (c).
Table 4: Base models and key hyperparameters for the quantile models used in CLEAR variants (a). Variant (b) uses only QXGB from this table. All models target the conditional median ($\tau=0.5$).
Model Key Hyperparameters Ref.
QRF 100 trees, min. leaf size: 10 [33]
QXGB 100 trees, tree method: histogram, min. child weight: 10 [38]
Expectile GAM 10 P-splines (order 3), smoothing parameter optimized via CV (5 min. timeout) [39]
Included only if hyperparameter optimization converged successfully, otherwise using the default values.
Table 5: Base models and key hyperparameters for the mean estimators (used in CLEAR variant c). All models target the conditional mean and are from Scikit-learn [109], except XGBoost, which is from [38]. All unspecified hyperparameters use the Scikit-learn defaults.
Category Model Key Hyperparameters
Linear Models Ordinary Least Squares Default
Ridge Default alphas (CV)
Lasso 3-fold CV
ElasticNet 3-fold CV
Tree Ensembles Random Forest 100 trees, min. leaf size: 5, max. features: 0.33
Extra Trees Same hyperparameters as random forest
AdaBoost Default
XGBoost Default
Neural Network MLP Single hidden layer (64 neurons)
Ensemble Construction and Model Selection:
For each of the $n_{boot}=100$ bootstrap samples, all models within the relevant pool (median estimators for variants a/b, mean estimators for variant c) were trained. The single best-performing model type ($k=1$) was then identified based on the lowest RMSE achieved on the held-out validation set. Consequently, the final PCS ensemble for each random seed comprised 100 instances of this selected top-performing model type. Specifically, CLEAR variant (a) considered all models from Table 4 for this selection process, variant (b) was restricted to selecting always QXGB from this pool, and variant (c) considered all the models from Table 4 instead. All models used default parameters from their respective libraries unless otherwise specified in the tables, and random states were fixed to ensure reproducibility.
Derivation of $\hat{f}$ and $\hat{q}^{\text{epi}}$ for CLEAR:
The point predictor $\hat{f}(x)$ and the raw epistemic uncertainty contributions $\hat{q}^{\text{epi}}_{\alpha/2}(x)$ and $\hat{q}^{\text{epi}}_{1-\alpha/2}(x)$ supplied to the CLEAR method are derived from this final ensemble of 100 model instances. As detailed in Section 2.2, $\hat{f}(x)$ is the pointwise empirical median of the ensemble’s predictions. The epistemic uncertainty terms represent the pointwise distances from this median to the ensemble’s empirical $\alpha/2$ and $1-\alpha/2$ quantiles, respectively.
Calibration of the Standalone PCS Baseline:
The standalone PCS baseline method involves a distinct calibration process. From the ensemble’s raw pointwise $\alpha/2$ and $1-\alpha/2$ quantile predictions ($\hat{f}_{\alpha/2}(x)$ and $\hat{f}_{1-\alpha/2}(x)$), a single, global multiplicative calibration factor, $\gamma_{\text{PCS}}$, is computed. This $\gamma_{\text{PCS}}$ is the smallest value ensuring that prediction intervals, formed by scaling the raw epistemic uncertainty around $\hat{f}(x)$ (i.e., $\hat{f}(x)-\gamma_{\text{PCS}}(\hat{f}(x)-\hat{f}_{\alpha/2}(x)),\,\hat{f}(x)+\gamma_{\text{PCS}}(\hat{f}_{1-\alpha/2}(x)-\hat{f}(x))]$), achieve the target $1-\alpha$ coverage on the validation set, incorporating the standard finite-sample correction. This $\gamma_{\text{PCS}}$ is then applied to generate the PCS baseline intervals on the test set. It is important to reiterate that this $\gamma_{\text{PCS}}$ is separate and computed independently from $\gamma_{1}$ and $\lambda$ parameters optimized within the CLEAR framework.
D.5 Metrics
This appendix provides detailed definitions for the evaluation metrics used in the main paper. We consider test data ${(X_{i},Y_{i})}_{i=1}^{N}$ and prediction intervals $L_{i},U_{i}]$.
  • Prediction Interval Coverage Probability (PICP): Measures the proportion of true values falling within the predicted intervals, calculated as:
    $\text{PICP}(L,U)=\frac{1}{N}\sum_{i=1}^{N}\mathds{1}_{[L_{i},U_{i}]}(Y_{i})$
    where $\mathds{1}_{[L_{i},U_{i}]}(Y_{i})$ is the indicator function; it equals 1 if $Y_{i}\in[L_{i},U_{i}]$ and 0 otherwise.
  • Normalized Interval Width (NIW): Quantifies the average width of prediction intervals normalized by the range of the target variable:
    $\text{NIW}(L,U)=\frac{\frac{1}{N}\sum_{i=1}^{N}(U_{i}-L_{i})}{\max(Y)-\min(Y)}$
  • Quantile Loss (also known as pinball loss): Evaluates the accuracy of predicted quantiles by penalizing both under- and overestimation. It reflects a trade-off between coverage (PICP) and interval width (NIW), rewarding narrow intervals that still maintain proper coverage and penalizing data points that are far outside the intervals (see Section B.3 for more details).
    For a given quantile level $\tau$, the quantile loss function is:
    $QL_{\tau}(y,q)=(y-q)\big(\tau-\mathds{1}_{(-\infty,q]}(y)\big),$
    where $q$ is the predicted $\tau$-quantile. For prediction intervals at level $1-\alpha$, we evaluate this at both $\tau=\alpha/2$ and $\tau=1-\alpha/2$ using
    $\text{QuantileLoss}(L,U)=\sum_{i=1}^{N}\left[QL_{\alpha/2}(Y_{i},L(X_{i}))+QL_{1-\alpha/2}(Y_{i},U(X_{i}))\right]/2.$
  • Average Interval Score Loss (AISL) [110]: This score balances interval width with coverage penalties (with similar properties as the quantile loss explained in Section B.3), defined as
    $\text{AISL}(L,U)=\frac{1}{N}\sum_{i=1}^{N}\bigg[$ $(U_{i}-L_{i})+\frac{2}{\alpha}(L_{i}-Y_{i})\,\mathds{1}\{Y_{i}<L_{i}\}+\frac{2}{\alpha}(Y_{i}-U_{i})\,\mathds{1}\{Y_{i}>U_{i}\}\bigg],$
    where $\mathds{1}\{\cdot\}$ is the indicator function.
Appendix E CLEAR with DE and SQR: Results on Real-World Data
This section presents the comprehensive experimental results of CLEAR when combined with deep learning-based uncertainty estimators: deep ensembles (DE) for epistemic uncertainty and simultaneous quantile regression (SQR) for aleatoric uncertainty. The experiments further validate CLEAR’s generality beyond the PCS and CQR methods that are presented in the body of our paper. The results demonstrate that CLEAR remains effective across different modeling paradigms. We evaluate both the neural baselines individually and their integration through CLEAR, comparing against conformalized versions to assess the added value of our dual-parameter calibration approach. Results are reported across all 17 datasets with 95% nominal coverage, using up to 10 random seeds for robustness.
Table 6: DE-SQR PICP at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.94 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
airfoil 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.98 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01
allstate 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.91 $\pm$ 0.01 0.95 $\pm$ 0.01 0.94 $\pm$ 0.01
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
computer 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.94 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
concrete 0.94 $\pm$ 0.02 0.95 $\pm$ 0.02 0.98 $\pm$ 0.01 0.96 $\pm$ 0.02 0.95 $\pm$ 0.03
elevator 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.94 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
energy_efficiency 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.99 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.03
insurance 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02
kin8nm 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.98 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
miami_housing 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00
naval_propulsion 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 1.00 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
parkinsons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
powerplant 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
qsar 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.93 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.96 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
Table 7: DE-SQR NIW at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor. Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values are the minimum (best) for that dataset, while underlined values indicate the second-best result.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 0.201 $\pm$ 0.017 0.284 $\pm$ 0.025 0.196 $\pm$ 0.014 0.285 $\pm$ 0.025 0.203 $\pm$ 0.015
airfoil 0.185 $\pm$ 0.023 0.223 $\pm$ 0.023 0.349 $\pm$ 0.025 0.226 $\pm$ 0.023 0.311 $\pm$ 0.025
allstate 0.333 $\pm$ 0.061 0.380 $\pm$ 0.073 0.283 $\pm$ 0.043 0.383 $\pm$ 0.075 0.324 $\pm$ 0.058
ca_housing 0.357 $\pm$ 9.1e-03 0.502 $\pm$ 0.020 0.378 $\pm$ 8.0e-03 0.502 $\pm$ 0.020 0.374 $\pm$ 4.0e-03
computer 0.087 $\pm$ 3.2e-03 0.121 $\pm$ 6.3e-03 0.104 $\pm$ 8.9e-03 0.122 $\pm$ 6.1e-03 0.106 $\pm$ 5.7e-03
concrete 0.264 $\pm$ 0.033 0.365 $\pm$ 0.068 0.441 $\pm$ 0.041 0.369 $\pm$ 0.067 0.392 $\pm$ 0.069
elevator 0.117 $\pm$ 6.6e-03 0.174 $\pm$ 0.011 0.114 $\pm$ 0.011 0.174 $\pm$ 0.011 0.117 $\pm$ 8.8e-03
energy_efficiency 0.085 $\pm$ 0.012 0.115 $\pm$ 0.020 0.480 $\pm$ 0.059 0.120 $\pm$ 0.020 0.368 $\pm$ 0.036
insurance 0.393 $\pm$ 0.044 0.523 $\pm$ 0.135 0.397 $\pm$ 0.073 0.558 $\pm$ 0.141 0.406 $\pm$ 0.085
kin8nm 0.197 $\pm$ 8.2e-03 0.260 $\pm$ 9.5e-03 0.246 $\pm$ 8.7e-03 0.261 $\pm$ 9.6e-03 0.208 $\pm$ 8.8e-03
miami_housing 0.093 $\pm$ 2.9e-03 0.131 $\pm$ 5.6e-03 0.101 $\pm$ 2.8e-03 0.131 $\pm$ 5.7e-03 0.100 $\pm$ 1.8e-03
naval_propulsion 1.5e-03 $\pm$ 4.2e-04 5.4e-03 $\pm$ 1.1e-03 3.7e-03 $\pm$ 1.3e-04 5.4e-03 $\pm$ 1.1e-03 1.7e-03 $\pm$ 1.5e-04
parkinsons 0.378 $\pm$ 0.014 0.476 $\pm$ 0.027 0.419 $\pm$ 0.021 0.479 $\pm$ 0.027 0.403 $\pm$ 0.017
powerplant 0.186 $\pm$ 6.7e-03 0.266 $\pm$ 0.012 0.204 $\pm$ 6.8e-03 0.267 $\pm$ 0.013 0.202 $\pm$ 7.3e-03
qsar 0.410 $\pm$ 0.136 0.508 $\pm$ 0.178 0.402 $\pm$ 0.135 0.509 $\pm$ 0.179 0.456 $\pm$ 0.153
sulfur 0.105 $\pm$ 8.5e-03 0.199 $\pm$ 0.015 0.120 $\pm$ 7.4e-03 0.200 $\pm$ 0.015 0.120 $\pm$ 8.3e-03
superconductor 0.208 $\pm$ 0.024 0.295 $\pm$ 0.038 0.247 $\pm$ 0.027 0.295 $\pm$ 0.038 0.245 $\pm$ 0.028
Table 8: DE-SQR Quantile Loss at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor. Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values are the minimum (best) for that dataset, while underlined values indicate the second-best result.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 9.6e-06 $\pm$ 3.1e-07 1.3e-05 $\pm$ 3.5e-07 9.6e-06 $\pm$ 4.0e-07 1.3e-05 $\pm$ 3.5e-07 9.5e-06 $\pm$ 3.7e-07
airfoil 0.097 $\pm$ 6.8e-03 0.117 $\pm$ 0.010 0.162 $\pm$ 0.017 0.117 $\pm$ 0.010 0.153 $\pm$ 0.018
allstate 1.4e+02 $\pm$ 9.329 1.5e+02 $\pm$ 8.112 1.5e+02 $\pm$ 9.210 1.5e+02 $\pm$ 8.106 1.4e+02 $\pm$ 8.368
ca_housing 3.0e+03 $\pm$ 63.719 3.9e+03 $\pm$ 1.2e+02 3.0e+03 $\pm$ 54.590 3.9e+03 $\pm$ 1.2e+02 3.0e+03 $\pm$ 53.118
computer 0.146 $\pm$ 5.5e-03 0.183 $\pm$ 6.8e-03 0.162 $\pm$ 8.3e-03 0.183 $\pm$ 6.8e-03 0.162 $\pm$ 7.9e-03
concrete 0.348 $\pm$ 0.049 0.427 $\pm$ 0.056 0.467 $\pm$ 0.031 0.428 $\pm$ 0.057 0.458 $\pm$ 0.051
elevator 1.2e-04 $\pm$ 1.9e-06 1.6e-04 $\pm$ 3.5e-06 1.1e-04 $\pm$ 2.5e-06 1.6e-04 $\pm$ 3.5e-06 1.1e-04 $\pm$ 2.4e-06
energy_efficiency 0.052 $\pm$ 9.5e-03 0.067 $\pm$ 0.012 0.221 $\pm$ 0.023 0.067 $\pm$ 0.012 0.188 $\pm$ 0.020
insurance 3.9e+02 $\pm$ 42.155 5.3e+02 $\pm$ 97.010 3.9e+02 $\pm$ 37.867 5.4e+02 $\pm$ 93.793 3.9e+02 $\pm$ 45.372
kin8nm 4.1e-03 $\pm$ 4.4e-05 5.2e-03 $\pm$ 8.2e-05 4.5e-03 $\pm$ 5.7e-05 5.2e-03 $\pm$ 8.2e-05 4.2e-03 $\pm$ 7.1e-05
miami_housing 4.5e+03 $\pm$ 2.1e+02 5.4e+03 $\pm$ 1.7e+02 4.4e+03 $\pm$ 1.2e+02 5.4e+03 $\pm$ 1.7e+02 4.4e+03 $\pm$ 1.2e+02
naval_propulsion 3.6e-05 $\pm$ 9.4e-06 1.3e-04 $\pm$ 2.7e-05 8.2e-05 $\pm$ 2.9e-06 1.3e-04 $\pm$ 2.7e-05 4.4e-05 $\pm$ 4.3e-06
parkinsons 0.298 $\pm$ 0.014 0.343 $\pm$ 9.8e-03 0.291 $\pm$ 5.7e-03 0.343 $\pm$ 9.8e-03 0.289 $\pm$ 5.8e-03
powerplant 0.227 $\pm$ 0.013 0.301 $\pm$ 0.014 0.234 $\pm$ 0.012 0.301 $\pm$ 0.014 0.234 $\pm$ 0.013
qsar 0.055 $\pm$ 3.2e-03 0.064 $\pm$ 3.5e-03 0.061 $\pm$ 3.7e-03 0.064 $\pm$ 3.5e-03 0.060 $\pm$ 3.6e-03
sulfur 1.7e-03 $\pm$ 1.0e-04 2.7e-03 $\pm$ 1.9e-04 1.8e-03 $\pm$ 1.3e-04 2.7e-03 $\pm$ 1.9e-04 1.8e-03 $\pm$ 1.3e-04
superconductor 0.545 $\pm$ 0.024 0.707 $\pm$ 0.024 0.560 $\pm$ 0.017 0.707 $\pm$ 0.024 0.559 $\pm$ 0.018
Table 9: DE-SQR NCIW at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor. Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values are the minimum (best) for that dataset, while underlined values indicate the second-best result.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 0.205 $\pm$ 0.017 0.292 $\pm$ 0.027 0.205 $\pm$ 0.017 0.291 $\pm$ 0.026 0.204 $\pm$ 0.017
airfoil 0.173 $\pm$ 0.016 0.220 $\pm$ 0.025 0.290 $\pm$ 0.026 0.217 $\pm$ 0.025 0.293 $\pm$ 0.026
allstate 0.328 $\pm$ 0.048 0.384 $\pm$ 0.062 0.342 $\pm$ 0.051 0.381 $\pm$ 0.060 0.330 $\pm$ 0.045
ca_housing 0.354 $\pm$ 8.1e-03 0.501 $\pm$ 0.021 0.372 $\pm$ 7.1e-03 0.501 $\pm$ 0.021 0.373 $\pm$ 7.8e-03
computer 0.088 $\pm$ 1.7e-03 0.120 $\pm$ 5.0e-03 0.106 $\pm$ 6.5e-03 0.120 $\pm$ 5.1e-03 0.106 $\pm$ 7.1e-03
concrete 0.274 $\pm$ 0.026 0.345 $\pm$ 0.053 0.371 $\pm$ 0.036 0.339 $\pm$ 0.051 0.376 $\pm$ 0.040
elevator 0.117 $\pm$ 7.0e-03 0.172 $\pm$ 0.010 0.117 $\pm$ 9.0e-03 0.172 $\pm$ 0.010 0.116 $\pm$ 9.4e-03
energy_efficiency 0.079 $\pm$ 7.4e-03 0.114 $\pm$ 0.017 0.351 $\pm$ 0.049 0.106 $\pm$ 0.014 0.346 $\pm$ 0.046
insurance 0.364 $\pm$ 0.060 0.465 $\pm$ 0.125 0.351 $\pm$ 0.064 0.460 $\pm$ 0.122 0.350 $\pm$ 0.061
kin8nm 0.194 $\pm$ 6.6e-03 0.255 $\pm$ 5.6e-03 0.205 $\pm$ 6.3e-03 0.254 $\pm$ 5.3e-03 0.208 $\pm$ 6.3e-03
miami_housing 0.092 $\pm$ 3.5e-03 0.129 $\pm$ 7.8e-03 0.100 $\pm$ 2.6e-03 0.129 $\pm$ 7.7e-03 0.100 $\pm$ 2.9e-03
naval_propulsion 1.5e-03 $\pm$ 4.2e-04 5.4e-03 $\pm$ 1.2e-03 1.7e-03 $\pm$ 2.0e-04 5.4e-03 $\pm$ 1.2e-03 1.6e-03 $\pm$ 1.7e-04
parkinsons 0.382 $\pm$ 0.015 0.473 $\pm$ 0.025 0.398 $\pm$ 0.013 0.470 $\pm$ 0.022 0.398 $\pm$ 0.013
powerplant 0.190 $\pm$ 6.0e-03 0.272 $\pm$ 9.5e-03 0.202 $\pm$ 8.1e-03 0.272 $\pm$ 9.5e-03 0.202 $\pm$ 8.1e-03
qsar 0.407 $\pm$ 0.134 0.477 $\pm$ 0.150 0.450 $\pm$ 0.151 0.476 $\pm$ 0.150 0.441 $\pm$ 0.146
sulfur 0.104 $\pm$ 6.2e-03 0.198 $\pm$ 0.016 0.120 $\pm$ 7.1e-03 0.197 $\pm$ 0.015 0.120 $\pm$ 7.2e-03
superconductor 0.209 $\pm$ 0.024 0.303 $\pm$ 0.036 0.242 $\pm$ 0.027 0.302 $\pm$ 0.036 0.245 $\pm$ 0.027
Table 10: DE-SQR Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds. CLEAR consists of the deep ensemble (DE) and the simultaneous quantile regressor (SQR). DE-conformal is the conformalized (calibrated) deep ensemble and SQR-conformal is the conformalized simultaneous quantile regressor. Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values are the minimum (best) for that dataset, while underlined values indicate the second-best result.
Dataset CLEAR DE SQR DE-conformal SQR-conformal
ailerons 7.7e-04 $\pm$ 2.5e-05 1.0e-03 $\pm$ 2.8e-05 7.7e-04 $\pm$ 3.2e-05 1.0e-03 $\pm$ 2.8e-05 7.6e-04 $\pm$ 2.9e-05
airfoil 7.761 $\pm$ 0.546 9.368 $\pm$ 0.826 12.937 $\pm$ 1.386 9.370 $\pm$ 0.813 12.279 $\pm$ 1.418
allstate 1.1e+04 $\pm$ 7.5e+02 1.2e+04 $\pm$ 6.5e+02 1.2e+04 $\pm$ 7.4e+02 1.2e+04 $\pm$ 6.5e+02 1.2e+04 $\pm$ 6.7e+02
ca_housing 2.4e+05 $\pm$ 5.1e+03 3.1e+05 $\pm$ 9.2e+03 2.4e+05 $\pm$ 4.4e+03 3.1e+05 $\pm$ 9.2e+03 2.4e+05 $\pm$ 4.2e+03
computer 11.669 $\pm$ 0.436 14.633 $\pm$ 0.543 12.938 $\pm$ 0.662 14.634 $\pm$ 0.545 12.926 $\pm$ 0.634
concrete 27.850 $\pm$ 3.946 34.179 $\pm$ 4.507 37.343 $\pm$ 2.464 34.203 $\pm$ 4.564 36.678 $\pm$ 4.051
elevator 9.2e-03 $\pm$ 1.5e-04 0.013 $\pm$ 2.8e-04 9.0e-03 $\pm$ 2.0e-04 0.013 $\pm$ 2.8e-04 9.0e-03 $\pm$ 1.9e-04
energy_efficiency 4.154 $\pm$ 0.758 5.369 $\pm$ 0.965 17.719 $\pm$ 1.871 5.387 $\pm$ 0.974 15.020 $\pm$ 1.595
insurance 3.1e+04 $\pm$ 3.4e+03 4.2e+04 $\pm$ 7.8e+03 3.1e+04 $\pm$ 3.0e+03 4.3e+04 $\pm$ 7.5e+03 3.1e+04 $\pm$ 3.6e+03
kin8nm 0.327 $\pm$ 3.5e-03 0.417 $\pm$ 6.5e-03 0.358 $\pm$ 4.6e-03 0.417 $\pm$ 6.6e-03 0.340 $\pm$ 5.7e-03
miami_housing 3.6e+05 $\pm$ 1.7e+04 4.3e+05 $\pm$ 1.4e+04 3.5e+05 $\pm$ 9.5e+03 4.3e+05 $\pm$ 1.4e+04 3.5e+05 $\pm$ 9.5e+03
naval_propulsion 2.9e-03 $\pm$ 7.5e-04 0.011 $\pm$ 2.2e-03 6.5e-03 $\pm$ 2.3e-04 0.011 $\pm$ 2.2e-03 3.5e-03 $\pm$ 3.5e-04
parkinsons 23.860 $\pm$ 1.124 27.437 $\pm$ 0.787 23.297 $\pm$ 0.456 27.449 $\pm$ 0.784 23.159 $\pm$ 0.461
powerplant 18.145 $\pm$ 1.005 24.110 $\pm$ 1.141 18.714 $\pm$ 0.990 24.106 $\pm$ 1.140 18.706 $\pm$ 1.002
qsar 4.419 $\pm$ 0.255 5.152 $\pm$ 0.278 4.875 $\pm$ 0.294 5.154 $\pm$ 0.278 4.800 $\pm$ 0.286
sulfur 0.134 $\pm$ 8.1e-03 0.213 $\pm$ 0.015 0.144 $\pm$ 0.010 0.213 $\pm$ 0.015 0.144 $\pm$ 0.010
superconductor 43.610 $\pm$ 1.937 56.594 $\pm$ 1.943 44.763 $\pm$ 1.388 56.591 $\pm$ 1.943 44.731 $\pm$ 1.406
Appendix F CLEAR with PCS and CQR: Results on Real-World Data
This appendix presents the detailed quantitative results from the benchmark experiments conducted on the 17 real-world datasets, complementing the summary findings in the main paper (Section 4.2). As a reminder, we evaluate three distinct configurations (variants) of our proposed CLEAR method alongside the baseline approaches (PCS, ALEATORIC, ALEATORIC-R). The differences among these 3 variants are considered base models. For each variant, CLEAR is applied on the already trained PCS and uncalibrated ALEATORIC-R. To recap these variants:
  • Variant (a): Employs a quantile PCS approach using a diverse pool of quantile estimators (QRF, QXGB, Expectile GAM [^6] Strictly speaking, Expectile GAM estimates expectiles rather than quantiles. While it is not a consistent estimator for quantiles in theory, mixing expectiles and quantiles may still be practical in applications.) for estimating the conditional medians based on the bootstraps. The top-performing model type ($k=1$) on the validation set is selected. 1) Based on the slected model type, we compute the epistemic component $\hat{q}^{\text{epi}}$ (via emprical quantiles over the estimated medians from the $b=m$ bootstraps) and the median $\hat{f}$ (as emprical median over the estimated medians). 2) Using the same selected quantile model type, the aleatoric component $\hat{q}^{\text{ale}}$ is derived from a bootstrapped ALEATORIC-R model.
  • Variant (b): Similar to (a), but restricts the model pool for both PCS and the paired ALEATORIC-R exclusively to QXGB.
  • Variant (c): Uses a standard PCS approach with mean-based regression models (e.g., Random Forest, XGBoost) for the epistemic component (via empirical quantiles over the estimated medians) and median prediction $\hat{f}$ (as empirical median over the estimated means), selecting the top model ($k=1$). The aleatoric component is derived from a bootstrapped ALEATORIC-R model using QRF.
The subsequent subsections present the comprehensive results for each variant (a, b, c). All experiments were run across 10 different random seeds [^7] Due to a bug in pygam’s Expectile GAM [39], for the naval_propulsion data in variant (c), one of the runs failed; hence, there are 9 runs available instead of 10., and the results on the test set are presented. The subsequent tables report the average performance metric across these 10 seeds, along with the standard deviation ($\pm\sigma$), to indicate the variability associated with data splitting. For each variant, we provide tables detailing the metrics presented in Section D.5 for all methods across all datasets at a 95% nominal coverage level. Lower values are preferable for all metrics except PICP, which should ideally be close to the target of 0.95. Additionally, summary plots (Figures 3, 7 and 8) visualize the relative NCIW and Quantile Loss performance normalized against the respective CLEAR variant baseline. Additionally, only for variant (c), we provide Table 27 with the values of $\lambda$ and $\gamma_{1}$ to provide some insights into the aleatoric/epistemic allocations of CLEAR. Finally, Figure 9 (Section F.4) provides a direct comparison of the NCIW and Quantile Loss between the three CLEAR variants themselves, using variant (a) as the reference baseline.
Section F.4 shows that the primary configuration of CLEAR presented in the main text, variant (a), leverages a diverse set of quantile models for robust epistemic uncertainty estimation and generally delivers the most consistent and superior performance. As detailed in Figure 9 in the appendix, CLEAR in variant (a) tends to marginally outperform the model in variant (b), which is restricted to QXGB, and variant (c), which utilizes standard mean-based PCS models. For instance, in variant (a), QXGB was selected in approximately 64% of cases, QRF in 24%, and ExpectileGAM in the remainder of cases. In variant (c), XGBoost was chosen about 70% of the time, with other models (excluding Ridge) sharing the rest. The importance of this dynamic model selection is evident in datasets like naval_propulsion, where variants (b) and (c) can struggle due to their fixed model choices (QXGB and QRF for the aleatoric part, respectively), whereas variant (a) can adapt by selecting, for example, ExpectileGAM. This highlights the stability advantage of CLEAR (a) and the benefit of PCS’s dynamic model choice, especially when the aleatoric component can also adapt. Nevertheless, the CLEAR framework overall demonstrates considerable robustness even with these alternative model choices. For instance, variant (b) still provides NCIW reductions of approximately 17.3% and 4.9% over its corresponding PCS and ALEATORIC-R baselines, respectively (Figure 7). A similar advantage is observed for variant (c) (Figure 8), which reduces NCIW by about 7.6% against ALEATORIC-R, the best-performing variant of CQR that uses our novel residual-based technique.
Detailed motivation behind variant (a).
Overall, variant (a) archives the best results (see Figure 9). Therefore, we recommend to use CLEAR variant (a). The experiments in variant (a) are fair, since each of the competing methods uses the same base model, which is selected per data set based on the RMSE on the validation data set. See Section F.1 for the results.
Detailed motivation behind variant (b).
Variant (b) is the simplest to understand, easiest to implement, and computationally cheapest variant and provides maximal fairness, making it particularly scientifically sound. Each method simply uses QXGB as base model without any model selection step. Strictly speaking, variant (b) is the only variant where ALEATORIC and ALEATORIC-R are fully conformal, since variant (b) uses the calibration data set only for calibration, while variant (a) and (c) reuse the validation set used for model selection as calibration data set. However, in practice, we observe this re-usage does not hurt calibration in any significant way (Table 12 and Table 22). See Section F.2 for the results of variant (b).
Detailed motivation behind variant (c).
Variant (c) uses the same set of base models as suggested by the authors of PCS uncertainty quantification as [30]. They conducted extensive experiments showing that PCS uncertainty quantification substantially outperforms popular conformal baselines such as split conformal regression [111], Studentized conformal regression [111], and Majority Vote [112] (by more than 20% on average in terms of interval width). Our experiments showing that CLEAR (a) outperforms CLEAR (c) (Figure 9) and CLEAR (c) outperforms PCS uncertainty quantification (c) (Section F.3), together with the experiments by [30], strongly suggests that CLEAR clearly outperforms these popular conformal baselines. See Section F.3 for the results of variant (c).
Overall, CLEAR shows the strongest performance across all variants (a), (b) and (c), demonstrating CLEAR’s stability across different settings, data sets, and metrics.
Table 11: Dataset statistics where $d$ represents the number of variables, $n$ represents the number of observations, followed by the minimum, maximum, and range values for $y$.
Dataset $n$ $d$ $y_{min}$ $y_{max}$ $y_{range}$
ailerons 13,750 40 -0.0036 0.00e+00 0.0036
airfoil 1,503 5 104.2040 140.1580 35.9540
allstate 5,000 1037 200.0000 3.31e+04 3.29e+04
ca_housing 20,640 8 1.50e+04 5.00e+05 4.85e+05
computer 8,192 21 0.00e+00 99.0000 99.0000
concrete 1,030 8 2.3300 81.7500 79.4200
elevator 16,599 18 0.0120 0.0780 0.0660
energy_efficiency 768 10 6.0100 43.1000 37.0900
insurance 1,338 8 1121.8739 6.26e+04 6.15e+04
kin8nm 8,192 8 0.0632 1.4585 1.3953
miami_housing 13,932 28 7.20e+04 2.65e+06 2.58e+06
naval_propulsion 11,934 24 0.0690 1.8320 1.7630
parkinsons 5,875 18 7.0000 54.9920 47.9920
powerplant 9,568 4 420.2600 495.7600 75.5000
qsar 5,742 500 -6.2400 11.0000 17.2400
sulfur 10,081 5 0.00e+00 1.0000 1.0000
superconductor 21,263 79 3.25e-04 185.0000 184.9997
F.1 Variant (a)
Table 12: Variant (a) PICP at 95% prediction intervals, aggregated across 10 seeds.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
airfoil 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01
allstate 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
computer 0.95 $\pm$ 0.01 0.97 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01
concrete 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01 0.94 $\pm$ 0.03 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
elevator 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
energy_efficiency 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
insurance 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.95 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01
kin8nm 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
miami_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
naval_propulsion 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
parkinsons 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
powerplant 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
qsar 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
Table 13: Variant (a) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.193 $\pm$ 0.016 0.208 $\pm$ 0.017 0.195 $\pm$ 0.017 0.215 $\pm$ 0.016 0.220 $\pm$ 0.021 0.193 $\pm$ 0.016 0.193 $\pm$ 0.016
airfoil 0.207 $\pm$ 0.012 0.367 $\pm$ 0.018 0.215 $\pm$ 0.018 0.215 $\pm$ 0.012 0.246 $\pm$ 0.024 0.205 $\pm$ 0.012 0.206 $\pm$ 0.012
allstate 0.260 $\pm$ 0.037 0.266 $\pm$ 0.045 0.259 $\pm$ 0.045 0.259 $\pm$ 0.044 0.327 $\pm$ 0.056 0.246 $\pm$ 0.042 0.247 $\pm$ 0.042
ca_housing 0.339 $\pm$ 0.008 0.760 $\pm$ 0.006 0.350 $\pm$ 0.006 0.354 $\pm$ 0.011 0.434 $\pm$ 0.020 0.333 $\pm$ 0.006 0.331 $\pm$ 0.007
computer 0.091 $\pm$ 0.008 0.099 $\pm$ 0.001 0.091 $\pm$ 0.010 18.269 $\pm$ 11.847 0.115 $\pm$ 0.007 0.091 $\pm$ 0.010 0.091 $\pm$ 0.007
concrete 0.249 $\pm$ 0.025 0.486 $\pm$ 0.028 0.265 $\pm$ 0.029 0.249 $\pm$ 0.017 0.263 $\pm$ 0.028 0.251 $\pm$ 0.022 0.253 $\pm$ 0.023
elevator 0.156 $\pm$ 0.009 0.144 $\pm$ 0.007 0.149 $\pm$ 0.007 0.178 $\pm$ 0.012 0.144 $\pm$ 0.008 0.156 $\pm$ 0.009 0.160 $\pm$ 0.008
energy_efficiency 0.047 $\pm$ 0.005 0.212 $\pm$ 0.015 0.050 $\pm$ 0.007 0.062 $\pm$ 0.006 0.062 $\pm$ 0.013 0.051 $\pm$ 0.004 0.053 $\pm$ 0.004
insurance 0.309 $\pm$ 0.032 0.332 $\pm$ 0.049 0.329 $\pm$ 0.056 0.585 $\pm$ 0.167 0.478 $\pm$ 0.070 0.331 $\pm$ 0.077 0.295 $\pm$ 0.044
kin8nm 0.360 $\pm$ 0.012 0.490 $\pm$ 0.020 0.374 $\pm$ 0.014 0.373 $\pm$ 0.014 0.397 $\pm$ 0.017 0.359 $\pm$ 0.012 0.360 $\pm$ 0.011
miami_housing 0.085 $\pm$ 0.002 0.105 $\pm$ 0.001 0.086 $\pm$ 0.004 0.098 $\pm$ 0.003 0.126 $\pm$ 0.006 0.084 $\pm$ 0.002 0.085 $\pm$ 0.002
naval_propulsion 6.1e-04 $\pm$ 6.3e-06 6.2e-04 $\pm$ 6.5e-06 6.0e-04 $\pm$ 6.4e-06 7.2e-04 $\pm$ 1.3e-05 6.0e-04 $\pm$ 7.0e-06 6.2e-04 $\pm$ 5.7e-06 6.1e-04 $\pm$ 5.6e-06
parkinsons 0.227 $\pm$ 0.010 0.321 $\pm$ 0.013 0.249 $\pm$ 0.014 0.312 $\pm$ 0.021 0.303 $\pm$ 0.022 0.228 $\pm$ 0.008 0.225 $\pm$ 0.008
powerplant 0.170 $\pm$ 0.007 0.196 $\pm$ 0.009 0.180 $\pm$ 0.009 0.178 $\pm$ 0.007 0.183 $\pm$ 0.006 0.173 $\pm$ 0.007 0.172 $\pm$ 0.007
qsar 0.363 $\pm$ 0.121 0.486 $\pm$ 0.160 0.386 $\pm$ 0.126 0.388 $\pm$ 0.131 0.420 $\pm$ 0.139 0.362 $\pm$ 0.121 0.362 $\pm$ 0.121
sulfur 0.109 $\pm$ 0.010 0.112 $\pm$ 0.010 0.107 $\pm$ 0.010 0.131 $\pm$ 0.017 0.122 $\pm$ 0.015 0.108 $\pm$ 0.011 0.108 $\pm$ 0.011
superconductor 0.196 $\pm$ 0.022 0.233 $\pm$ 0.025 0.196 $\pm$ 0.023 0.248 $\pm$ 0.028 0.329 $\pm$ 0.042 0.194 $\pm$ 0.022 0.194 $\pm$ 0.021
Table 14: Variant (a) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 9.2e-06 $\pm$ 2.3e-07 1.0e-05 $\pm$ 2.8e-07 9.7e-06 $\pm$ 2.8e-07 1.0e-05 $\pm$ 2.2e-07 1.1e-05 $\pm$ 3.7e-07 9.2e-06 $\pm$ 2.3e-07 9.2e-06 $\pm$ 2.3e-07
airfoil 0.109 $\pm$ 0.011 0.183 $\pm$ 0.013 0.123 $\pm$ 0.016 0.117 $\pm$ 0.012 0.147 $\pm$ 0.019 0.109 $\pm$ 0.011 0.109 $\pm$ 0.011
allstate 1.1e+02 $\pm$ 7.835 1.3e+02 $\pm$ 8.197 1.3e+02 $\pm$ 7.925 1.2e+02 $\pm$ 7.770 1.7e+02 $\pm$ 13.696 1.1e+02 $\pm$ 7.563 1.1e+02 $\pm$ 7.631
ca_housing 2.8e+03 $\pm$ 59.700 4.8e+03 $\pm$ 25.753 3.0e+03 $\pm$ 73.283 3.2e+03 $\pm$ 93.396 4.0e+03 $\pm$ 1.2e+02 2.8e+03 $\pm$ 66.414 2.8e+03 $\pm$ 70.545
computer 0.147 $\pm$ 0.022 0.150 $\pm$ 0.012 0.160 $\pm$ 0.029 22.643 $\pm$ 14.662 0.192 $\pm$ 0.010 0.146 $\pm$ 0.020 0.147 $\pm$ 0.023
concrete 0.338 $\pm$ 0.040 0.546 $\pm$ 0.061 0.359 $\pm$ 0.053 0.327 $\pm$ 0.036 0.376 $\pm$ 0.058 0.330 $\pm$ 0.037 0.331 $\pm$ 0.037
elevator 1.5e-04 $\pm$ 2.0e-06 1.4e-04 $\pm$ 2.9e-06 1.5e-04 $\pm$ 2.4e-06 1.7e-04 $\pm$ 5.6e-06 1.6e-04 $\pm$ 5.1e-06 1.5e-04 $\pm$ 2.4e-06 1.5e-04 $\pm$ 2.0e-06
energy_efficiency 0.031 $\pm$ 0.004 0.102 $\pm$ 0.008 0.033 $\pm$ 0.005 0.038 $\pm$ 0.003 0.041 $\pm$ 0.005 0.033 $\pm$ 0.003 0.034 $\pm$ 0.003
insurance 3.5e+02 $\pm$ 61.616 3.6e+02 $\pm$ 39.928 3.6e+02 $\pm$ 54.512 5.8e+02 $\pm$ 1.0e+02 4.4e+02 $\pm$ 38.440 3.6e+02 $\pm$ 68.243 3.5e+02 $\pm$ 62.851
kin8nm 7.2e-03 $\pm$ 1.7e-04 9.5e-03 $\pm$ 1.6e-04 7.7e-03 $\pm$ 2.7e-04 7.6e-03 $\pm$ 9.7e-05 8.3e-03 $\pm$ 2.5e-04 7.2e-03 $\pm$ 1.6e-04 7.2e-03 $\pm$ 1.7e-04
miami_housing 3.9e+03 $\pm$ 1.9e+02 5.1e+03 $\pm$ 3.1e+02 5.2e+03 $\pm$ 3.4e+02 4.2e+03 $\pm$ 1.6e+02 8.1e+03 $\pm$ 3.6e+02 3.9e+03 $\pm$ 1.9e+02 3.9e+03 $\pm$ 1.8e+02
naval_propulsion 1.5e-05 $\pm$ 2.1e-07 1.5e-05 $\pm$ 2.6e-07 1.5e-05 $\pm$ 2.6e-07 1.7e-05 $\pm$ 2.7e-07 1.6e-05 $\pm$ 3.5e-07 1.5e-05 $\pm$ 2.1e-07 1.5e-05 $\pm$ 2.3e-07
parkinsons 0.178 $\pm$ 0.010 0.202 $\pm$ 0.007 0.215 $\pm$ 0.012 0.224 $\pm$ 0.012 0.251 $\pm$ 0.014 0.185 $\pm$ 0.011 0.178 $\pm$ 0.010
powerplant 0.212 $\pm$ 0.012 0.228 $\pm$ 0.011 0.220 $\pm$ 0.011 0.221 $\pm$ 0.012 0.231 $\pm$ 0.010 0.213 $\pm$ 0.012 0.213 $\pm$ 0.012
qsar 0.049 $\pm$ 0.003 0.060 $\pm$ 0.003 0.052 $\pm$ 0.003 0.053 $\pm$ 0.003 0.057 $\pm$ 0.003 0.049 $\pm$ 0.003 0.049 $\pm$ 0.003
sulfur 2.0e-03 $\pm$ 1.4e-04 2.1e-03 $\pm$ 1.3e-04 2.3e-03 $\pm$ 1.1e-04 2.3e-03 $\pm$ 1.9e-04 3.5e-03 $\pm$ 2.0e-04 2.0e-03 $\pm$ 1.4e-04 2.0e-03 $\pm$ 1.4e-04
superconductor 0.490 $\pm$ 0.018 0.511 $\pm$ 0.014 0.538 $\pm$ 0.023 0.608 $\pm$ 0.023 0.861 $\pm$ 0.036 0.490 $\pm$ 0.018 0.491 $\pm$ 0.018
Table 15: Variant (a) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.195 $\pm$ 0.017 0.209 $\pm$ 0.019 0.196 $\pm$ 0.017 0.217 $\pm$ 0.021 0.220 $\pm$ 0.020 0.195 $\pm$ 0.017 0.195 $\pm$ 0.017
airfoil 0.203 $\pm$ 0.006 0.356 $\pm$ 0.016 0.216 $\pm$ 0.012 0.211 $\pm$ 0.007 0.246 $\pm$ 0.019 0.200 $\pm$ 0.007 0.200 $\pm$ 0.006
allstate 0.252 $\pm$ 0.037 0.264 $\pm$ 0.042 0.262 $\pm$ 0.042 0.266 $\pm$ 0.051 0.328 $\pm$ 0.051 0.245 $\pm$ 0.042 0.245 $\pm$ 0.041
ca_housing 0.336 $\pm$ 0.008 0.760 $\pm$ 0.009 0.348 $\pm$ 0.010 0.354 $\pm$ 0.012 0.440 $\pm$ 0.016 0.331 $\pm$ 0.007 0.328 $\pm$ 0.007
computer 0.091 $\pm$ 0.008 0.099 $\pm$ 0.001 0.091 $\pm$ 0.011 0.112 $\pm$ 0.028 0.113 $\pm$ 0.008 0.092 $\pm$ 0.009 0.091 $\pm$ 0.007
concrete 0.246 $\pm$ 0.033 0.421 $\pm$ 0.079 0.259 $\pm$ 0.035 0.241 $\pm$ 0.025 0.264 $\pm$ 0.037 0.238 $\pm$ 0.027 0.240 $\pm$ 0.026
elevator 0.155 $\pm$ 0.010 0.143 $\pm$ 0.008 0.148 $\pm$ 0.008 0.177 $\pm$ 0.011 0.145 $\pm$ 0.010 0.155 $\pm$ 0.010 0.159 $\pm$ 0.010
energy_efficiency 0.046 $\pm$ 0.005 0.196 $\pm$ 0.032 0.050 $\pm$ 0.005 0.057 $\pm$ 0.004 0.060 $\pm$ 0.006 0.050 $\pm$ 0.005 0.051 $\pm$ 0.004
insurance 0.303 $\pm$ 0.029 0.284 $\pm$ 0.049 0.308 $\pm$ 0.043 0.530 $\pm$ 0.178 0.454 $\pm$ 0.102 0.282 $\pm$ 0.049 0.271 $\pm$ 0.046
kin8nm 0.354 $\pm$ 0.008 0.482 $\pm$ 0.014 0.371 $\pm$ 0.009 0.371 $\pm$ 0.009 0.396 $\pm$ 0.013 0.355 $\pm$ 0.008 0.355 $\pm$ 0.008
miami_housing 0.084 $\pm$ 0.003 0.106 $\pm$ 0.004 0.085 $\pm$ 0.003 0.098 $\pm$ 0.002 0.124 $\pm$ 0.006 0.083 $\pm$ 0.002 0.084 $\pm$ 0.002
naval_propulsion 6.1e-04 $\pm$ 7.9e-06 6.1e-04 $\pm$ 7.4e-06 6.0e-04 $\pm$ 5.5e-06 7.1e-04 $\pm$ 1.5e-05 5.9e-04 $\pm$ 6.1e-06 6.1e-04 $\pm$ 9.3e-06 6.1e-04 $\pm$ 6.4e-06
parkinsons 0.225 $\pm$ 0.012 0.321 $\pm$ 0.013 0.247 $\pm$ 0.007 0.316 $\pm$ 0.032 0.302 $\pm$ 0.013 0.227 $\pm$ 0.007 0.224 $\pm$ 0.008
powerplant 0.173 $\pm$ 0.007 0.197 $\pm$ 0.009 0.182 $\pm$ 0.007 0.182 $\pm$ 0.007 0.187 $\pm$ 0.008 0.176 $\pm$ 0.007 0.174 $\pm$ 0.007
qsar 0.360 $\pm$ 0.119 0.479 $\pm$ 0.156 0.389 $\pm$ 0.130 0.389 $\pm$ 0.134 0.421 $\pm$ 0.142 0.360 $\pm$ 0.120 0.360 $\pm$ 0.120
sulfur 0.109 $\pm$ 0.010 0.111 $\pm$ 0.008 0.106 $\pm$ 0.009 0.128 $\pm$ 0.013 0.129 $\pm$ 0.011 0.108 $\pm$ 0.011 0.107 $\pm$ 0.010
superconductor 0.195 $\pm$ 0.021 0.231 $\pm$ 0.025 0.194 $\pm$ 0.021 0.247 $\pm$ 0.028 0.314 $\pm$ 0.035 0.194 $\pm$ 0.021 0.194 $\pm$ 0.021
Table 16: Variant (a) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 7.4e-04 $\pm$ 1.8e-05 8.1e-04 $\pm$ 2.2e-05 7.8e-04 $\pm$ 2.2e-05 8.0e-04 $\pm$ 1.8e-05 9.1e-04 $\pm$ 2.9e-05 7.4e-04 $\pm$ 1.8e-05 7.4e-04 $\pm$ 1.8e-05
airfoil 8.717 $\pm$ 0.870 14.669 $\pm$ 1.007 9.843 $\pm$ 1.249 9.397 $\pm$ 0.977 11.761 $\pm$ 1.548 8.727 $\pm$ 0.902 8.730 $\pm$ 0.892
allstate 9.1e+03 $\pm$ 6.3e+02 1.1e+04 $\pm$ 6.6e+02 1.0e+04 $\pm$ 6.3e+02 9.9e+03 $\pm$ 6.2e+02 1.3e+04 $\pm$ 1.1e+03 9.2e+03 $\pm$ 6.1e+02 9.2e+03 $\pm$ 6.1e+02
ca_housing 2.2e+05 $\pm$ 4.8e+03 3.9e+05 $\pm$ 2.1e+03 2.4e+05 $\pm$ 5.9e+03 2.6e+05 $\pm$ 7.5e+03 3.2e+05 $\pm$ 9.7e+03 2.2e+05 $\pm$ 5.3e+03 2.2e+05 $\pm$ 5.6e+03
computer 11.741 $\pm$ 1.723 12.023 $\pm$ 0.924 12.761 $\pm$ 2.285 1.8e+03 $\pm$ 1.2e+03 15.360 $\pm$ 0.785 11.667 $\pm$ 1.598 11.761 $\pm$ 1.841
concrete 27.020 $\pm$ 3.202 43.690 $\pm$ 4.842 28.684 $\pm$ 4.251 26.165 $\pm$ 2.857 30.054 $\pm$ 4.640 26.365 $\pm$ 2.976 26.445 $\pm$ 2.996
elevator 0.012 $\pm$ 0.000 0.011 $\pm$ 0.000 0.012 $\pm$ 0.000 0.014 $\pm$ 0.000 0.013 $\pm$ 0.000 0.012 $\pm$ 0.000 0.012 $\pm$ 0.000
energy_efficiency 2.465 $\pm$ 0.347 8.199 $\pm$ 0.672 2.671 $\pm$ 0.416 3.058 $\pm$ 0.217 3.253 $\pm$ 0.388 2.667 $\pm$ 0.280 2.721 $\pm$ 0.270
insurance 2.8e+04 $\pm$ 4.9e+03 2.9e+04 $\pm$ 3.2e+03 2.8e+04 $\pm$ 4.4e+03 4.7e+04 $\pm$ 8.1e+03 3.6e+04 $\pm$ 3.1e+03 2.9e+04 $\pm$ 5.5e+03 2.8e+04 $\pm$ 5.0e+03
kin8nm 0.577 $\pm$ 0.014 0.760 $\pm$ 0.013 0.616 $\pm$ 0.021 0.607 $\pm$ 0.008 0.666 $\pm$ 0.020 0.577 $\pm$ 0.013 0.577 $\pm$ 0.013
miami_housing 3.1e+05 $\pm$ 1.5e+04 4.1e+05 $\pm$ 2.5e+04 4.2e+05 $\pm$ 2.7e+04 3.3e+05 $\pm$ 1.3e+04 6.5e+05 $\pm$ 2.9e+04 3.1e+05 $\pm$ 1.5e+04 3.1e+05 $\pm$ 1.5e+04
naval_propulsion 1.2e-03 $\pm$ 1.7e-05 1.2e-03 $\pm$ 2.1e-05 1.2e-03 $\pm$ 2.1e-05 1.4e-03 $\pm$ 2.2e-05 1.3e-03 $\pm$ 2.8e-05 1.2e-03 $\pm$ 1.7e-05 1.2e-03 $\pm$ 1.8e-05
parkinsons 14.221 $\pm$ 0.792 16.142 $\pm$ 0.582 17.188 $\pm$ 0.985 17.947 $\pm$ 0.957 20.088 $\pm$ 1.130 14.762 $\pm$ 0.882 14.259 $\pm$ 0.786
powerplant 16.993 $\pm$ 0.937 18.211 $\pm$ 0.897 17.577 $\pm$ 0.884 17.715 $\pm$ 0.965 18.501 $\pm$ 0.815 17.064 $\pm$ 0.951 17.027 $\pm$ 0.954
qsar 3.951 $\pm$ 0.230 4.795 $\pm$ 0.219 4.187 $\pm$ 0.231 4.249 $\pm$ 0.217 4.558 $\pm$ 0.216 3.950 $\pm$ 0.225 3.953 $\pm$ 0.224
sulfur 0.161 $\pm$ 0.011 0.172 $\pm$ 0.011 0.187 $\pm$ 0.008 0.186 $\pm$ 0.015 0.280 $\pm$ 0.016 0.161 $\pm$ 0.011 0.161 $\pm$ 0.011
superconductor 39.217 $\pm$ 1.419 40.851 $\pm$ 1.157 43.049 $\pm$ 1.848 48.630 $\pm$ 1.867 68.891 $\pm$ 2.882 39.182 $\pm$ 1.438 39.296 $\pm$ 1.444
F.2 Variant (b)
Figure 7: Quantile loss and NCIW performance of different methods (CLEAR, PCS, ALEATORIC, ALEATORIC-R) for variant (b), over 10 seeds normalized relative to CLEAR (baseline = 1.0). Lower values indicate better performance. The inset boxplot shows the % improvement relative to the CLEAR baseline $\pm 1\sigma$. Values inside each subplot represent the mean improvement across all datasets.
Table 17: Variant (b) PICP at 95% prediction intervals, aggregated across 10 seeds.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
airfoil 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01
allstate 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
computer 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
concrete 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01 0.94 $\pm$ 0.03 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
elevator 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
energy_efficiency 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
insurance 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.95 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
kin8nm 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
miami_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
naval_propulsion 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 1.00 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00
parkinsons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
powerplant 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
qsar 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
Table 18: Variant (b) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.193 $\pm$ 0.016 0.208 $\pm$ 0.017 0.195 $\pm$ 0.017 0.215 $\pm$ 0.016 0.220 $\pm$ 0.021 0.193 $\pm$ 0.016 0.193 $\pm$ 0.016
airfoil 0.207 $\pm$ 0.012 0.367 $\pm$ 0.018 0.215 $\pm$ 0.018 0.215 $\pm$ 0.012 0.246 $\pm$ 0.024 0.205 $\pm$ 0.012 0.206 $\pm$ 0.012
allstate 0.256 $\pm$ 0.045 0.265 $\pm$ 0.047 0.260 $\pm$ 0.047 0.263 $\pm$ 0.053 0.325 $\pm$ 0.055 0.243 $\pm$ 0.044 0.245 $\pm$ 0.046
ca_housing 0.339 $\pm$ 0.008 0.760 $\pm$ 0.006 0.350 $\pm$ 0.006 0.354 $\pm$ 0.011 0.434 $\pm$ 0.020 0.333 $\pm$ 0.006 0.331 $\pm$ 0.007
computer 0.099 $\pm$ 0.006 0.143 $\pm$ 0.029 0.104 $\pm$ 0.005 0.142 $\pm$ 0.009 0.113 $\pm$ 0.006 0.101 $\pm$ 0.005 0.104 $\pm$ 0.005
concrete 0.249 $\pm$ 0.025 0.486 $\pm$ 0.028 0.265 $\pm$ 0.029 0.249 $\pm$ 0.017 0.263 $\pm$ 0.028 0.251 $\pm$ 0.022 0.253 $\pm$ 0.023
elevator 0.137 $\pm$ 0.007 0.171 $\pm$ 0.010 0.146 $\pm$ 0.008 0.193 $\pm$ 0.013 0.169 $\pm$ 0.010 0.136 $\pm$ 0.008 0.137 $\pm$ 0.008
energy_efficiency 0.047 $\pm$ 0.005 0.212 $\pm$ 0.015 0.050 $\pm$ 0.007 0.062 $\pm$ 0.006 0.062 $\pm$ 0.013 0.051 $\pm$ 0.004 0.053 $\pm$ 0.004
insurance 0.329 $\pm$ 0.045 0.407 $\pm$ 0.064 0.381 $\pm$ 0.055 0.434 $\pm$ 0.179 0.478 $\pm$ 0.070 0.339 $\pm$ 0.104 0.306 $\pm$ 0.058
kin8nm 0.360 $\pm$ 0.012 0.490 $\pm$ 0.020 0.374 $\pm$ 0.014 0.373 $\pm$ 0.014 0.397 $\pm$ 0.017 0.359 $\pm$ 0.012 0.360 $\pm$ 0.011
miami_housing 0.085 $\pm$ 0.002 0.105 $\pm$ 0.001 0.086 $\pm$ 0.004 0.098 $\pm$ 0.003 0.126 $\pm$ 0.006 0.084 $\pm$ 0.002 0.085 $\pm$ 0.002
naval_propulsion 1.9e-03 $\pm$ 7.3e-05 0.222 $\pm$ 0.003 1.9e-03 $\pm$ 9.7e-05 8.1e-03 $\pm$ 4.1e-04 2.4e-03 $\pm$ 1.4e-04 2.1e-03 $\pm$ 8.1e-05 2.8e-03 $\pm$ 1.3e-04
parkinsons 0.281 $\pm$ 0.009 0.483 $\pm$ 0.006 0.302 $\pm$ 0.012 0.325 $\pm$ 0.011 0.345 $\pm$ 0.011 0.282 $\pm$ 0.009 0.281 $\pm$ 0.009
powerplant 0.170 $\pm$ 0.007 0.196 $\pm$ 0.009 0.180 $\pm$ 0.009 0.178 $\pm$ 0.007 0.183 $\pm$ 0.006 0.173 $\pm$ 0.007 0.172 $\pm$ 0.007
qsar 0.366 $\pm$ 0.123 0.486 $\pm$ 0.160 0.387 $\pm$ 0.127 0.393 $\pm$ 0.134 0.421 $\pm$ 0.139 0.362 $\pm$ 0.121 0.362 $\pm$ 0.121
sulfur 0.108 $\pm$ 0.009 0.110 $\pm$ 0.008 0.107 $\pm$ 0.011 0.129 $\pm$ 0.014 0.122 $\pm$ 0.015 0.106 $\pm$ 0.009 0.106 $\pm$ 0.009
superconductor 0.219 $\pm$ 0.023 0.309 $\pm$ 0.033 0.228 $\pm$ 0.026 0.243 $\pm$ 0.027 0.354 $\pm$ 0.042 0.223 $\pm$ 0.025 0.216 $\pm$ 0.024
Table 19: Variant (b) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 9.2e-06 $\pm$ 2.3e-07 1.0e-05 $\pm$ 2.8e-07 9.7e-06 $\pm$ 2.8e-07 1.0e-05 $\pm$ 2.2e-07 1.1e-05 $\pm$ 3.7e-07 9.2e-06 $\pm$ 2.3e-07 9.2e-06 $\pm$ 2.3e-07
airfoil 0.109 $\pm$ 0.011 0.183 $\pm$ 0.013 0.123 $\pm$ 0.016 0.117 $\pm$ 0.012 0.147 $\pm$ 0.019 0.109 $\pm$ 0.011 0.109 $\pm$ 0.011
allstate 1.1e+02 $\pm$ 7.318 1.4e+02 $\pm$ 8.321 1.3e+02 $\pm$ 8.046 1.3e+02 $\pm$ 8.557 1.7e+02 $\pm$ 10.782 1.2e+02 $\pm$ 7.533 1.2e+02 $\pm$ 7.762
ca_housing 2.8e+03 $\pm$ 59.700 4.8e+03 $\pm$ 25.753 3.0e+03 $\pm$ 73.283 3.2e+03 $\pm$ 93.396 4.0e+03 $\pm$ 1.2e+02 2.8e+03 $\pm$ 66.414 2.8e+03 $\pm$ 70.545
computer 0.159 $\pm$ 0.007 0.214 $\pm$ 0.041 0.206 $\pm$ 0.015 0.207 $\pm$ 0.009 0.234 $\pm$ 0.019 0.159 $\pm$ 0.007 0.161 $\pm$ 0.008
concrete 0.338 $\pm$ 0.040 0.546 $\pm$ 0.061 0.359 $\pm$ 0.053 0.327 $\pm$ 0.036 0.376 $\pm$ 0.058 0.330 $\pm$ 0.037 0.331 $\pm$ 0.037
elevator 1.4e-04 $\pm$ 2.0e-06 1.8e-04 $\pm$ 3.0e-06 1.6e-04 $\pm$ 3.1e-06 1.9e-04 $\pm$ 6.2e-06 2.1e-04 $\pm$ 5.5e-06 1.4e-04 $\pm$ 2.0e-06 1.4e-04 $\pm$ 2.1e-06
energy_efficiency 0.031 $\pm$ 0.004 0.102 $\pm$ 0.008 0.033 $\pm$ 0.005 0.038 $\pm$ 0.003 0.041 $\pm$ 0.005 0.033 $\pm$ 0.003 0.034 $\pm$ 0.003
insurance 3.2e+02 $\pm$ 31.593 3.9e+02 $\pm$ 36.904 3.7e+02 $\pm$ 35.107 4.9e+02 $\pm$ 88.647 4.4e+02 $\pm$ 34.159 3.7e+02 $\pm$ 62.380 3.4e+02 $\pm$ 40.487
kin8nm 7.2e-03 $\pm$ 1.7e-04 9.5e-03 $\pm$ 1.6e-04 7.7e-03 $\pm$ 2.7e-04 7.6e-03 $\pm$ 9.7e-05 8.3e-03 $\pm$ 2.5e-04 7.2e-03 $\pm$ 1.6e-04 7.2e-03 $\pm$ 1.7e-04
miami_housing 3.9e+03 $\pm$ 1.9e+02 5.1e+03 $\pm$ 3.1e+02 5.2e+03 $\pm$ 3.4e+02 4.2e+03 $\pm$ 1.6e+02 8.1e+03 $\pm$ 3.6e+02 3.9e+03 $\pm$ 1.9e+02 3.9e+03 $\pm$ 1.8e+02
naval_propulsion 5.4e-05 $\pm$ 1.9e-06 4.9e-03 $\pm$ 7.0e-05 6.4e-05 $\pm$ 2.7e-06 1.8e-04 $\pm$ 9.1e-06 8.8e-05 $\pm$ 5.9e-06 6.0e-05 $\pm$ 1.8e-06 7.5e-05 $\pm$ 2.7e-06
parkinsons 0.209 $\pm$ 0.010 0.319 $\pm$ 0.009 0.228 $\pm$ 0.014 0.237 $\pm$ 0.008 0.265 $\pm$ 0.013 0.209 $\pm$ 0.010 0.209 $\pm$ 0.009
powerplant 0.212 $\pm$ 0.012 0.228 $\pm$ 0.011 0.220 $\pm$ 0.011 0.221 $\pm$ 0.012 0.231 $\pm$ 0.010 0.213 $\pm$ 0.012 0.213 $\pm$ 0.012
qsar 0.049 $\pm$ 0.003 0.060 $\pm$ 0.003 0.052 $\pm$ 0.003 0.053 $\pm$ 0.003 0.057 $\pm$ 0.003 0.049 $\pm$ 0.003 0.049 $\pm$ 0.003
sulfur 1.9e-03 $\pm$ 9.6e-05 2.2e-03 $\pm$ 1.1e-04 2.3e-03 $\pm$ 1.4e-04 2.2e-03 $\pm$ 8.6e-05 3.5e-03 $\pm$ 2.4e-04 1.9e-03 $\pm$ 8.2e-05 1.9e-03 $\pm$ 8.6e-05
superconductor 0.523 $\pm$ 0.018 0.648 $\pm$ 0.020 0.573 $\pm$ 0.024 0.611 $\pm$ 0.025 0.892 $\pm$ 0.035 0.522 $\pm$ 0.018 0.523 $\pm$ 0.020
Table 20: Variant (b) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.195 $\pm$ 0.017 0.209 $\pm$ 0.019 0.196 $\pm$ 0.017 0.217 $\pm$ 0.021 0.220 $\pm$ 0.020 0.195 $\pm$ 0.017 0.195 $\pm$ 0.017
airfoil 0.203 $\pm$ 0.006 0.356 $\pm$ 0.016 0.216 $\pm$ 0.012 0.211 $\pm$ 0.007 0.246 $\pm$ 0.019 0.200 $\pm$ 0.007 0.200 $\pm$ 0.006
allstate 0.250 $\pm$ 0.043 0.265 $\pm$ 0.043 0.266 $\pm$ 0.044 0.270 $\pm$ 0.052 0.329 $\pm$ 0.053 0.245 $\pm$ 0.045 0.244 $\pm$ 0.044
ca_housing 0.336 $\pm$ 0.008 0.760 $\pm$ 0.009 0.348 $\pm$ 0.010 0.354 $\pm$ 0.012 0.440 $\pm$ 0.016 0.331 $\pm$ 0.007 0.328 $\pm$ 0.007
computer 0.098 $\pm$ 0.005 0.144 $\pm$ 0.034 0.103 $\pm$ 0.003 0.142 $\pm$ 0.008 0.111 $\pm$ 0.005 0.101 $\pm$ 0.004 0.105 $\pm$ 0.004
concrete 0.246 $\pm$ 0.033 0.421 $\pm$ 0.079 0.259 $\pm$ 0.035 0.241 $\pm$ 0.025 0.264 $\pm$ 0.037 0.238 $\pm$ 0.027 0.240 $\pm$ 0.026
elevator 0.137 $\pm$ 0.008 0.173 $\pm$ 0.011 0.147 $\pm$ 0.009 0.192 $\pm$ 0.012 0.168 $\pm$ 0.012 0.137 $\pm$ 0.007 0.137 $\pm$ 0.007
energy_efficiency 0.046 $\pm$ 0.005 0.196 $\pm$ 0.032 0.050 $\pm$ 0.005 0.057 $\pm$ 0.004 0.060 $\pm$ 0.006 0.050 $\pm$ 0.005 0.051 $\pm$ 0.004
insurance 0.320 $\pm$ 0.059 0.344 $\pm$ 0.069 0.339 $\pm$ 0.066 0.346 $\pm$ 0.076 0.455 $\pm$ 0.090 0.292 $\pm$ 0.058 0.275 $\pm$ 0.054
kin8nm 0.354 $\pm$ 0.008 0.482 $\pm$ 0.014 0.371 $\pm$ 0.009 0.371 $\pm$ 0.009 0.396 $\pm$ 0.013 0.355 $\pm$ 0.008 0.355 $\pm$ 0.008
miami_housing 0.084 $\pm$ 0.003 0.106 $\pm$ 0.004 0.085 $\pm$ 0.003 0.098 $\pm$ 0.002 0.124 $\pm$ 0.006 0.083 $\pm$ 0.002 0.084 $\pm$ 0.002
naval_propulsion 1.8e-03 $\pm$ 5.6e-05 0.195 $\pm$ 0.011 1.9e-03 $\pm$ 8.0e-05 3.2e-03 $\pm$ 1.2e-04 2.4e-03 $\pm$ 1.4e-04 2.1e-03 $\pm$ 7.2e-05 2.8e-03 $\pm$ 1.1e-04
parkinsons 0.284 $\pm$ 0.010 0.473 $\pm$ 0.016 0.309 $\pm$ 0.018 0.324 $\pm$ 0.008 0.349 $\pm$ 0.019 0.285 $\pm$ 0.011 0.285 $\pm$ 0.011
powerplant 0.173 $\pm$ 0.007 0.197 $\pm$ 0.009 0.182 $\pm$ 0.007 0.182 $\pm$ 0.007 0.187 $\pm$ 0.008 0.176 $\pm$ 0.007 0.174 $\pm$ 0.007
qsar 0.363 $\pm$ 0.121 0.480 $\pm$ 0.157 0.390 $\pm$ 0.130 0.393 $\pm$ 0.131 0.423 $\pm$ 0.142 0.362 $\pm$ 0.120 0.362 $\pm$ 0.120
sulfur 0.106 $\pm$ 0.007 0.109 $\pm$ 0.006 0.105 $\pm$ 0.006 0.125 $\pm$ 0.009 0.127 $\pm$ 0.008 0.105 $\pm$ 0.006 0.104 $\pm$ 0.006
superconductor 0.218 $\pm$ 0.022 0.308 $\pm$ 0.032 0.226 $\pm$ 0.022 0.241 $\pm$ 0.028 0.347 $\pm$ 0.036 0.223 $\pm$ 0.024 0.216 $\pm$ 0.024
Table 21: Variant (b) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 7.4e-04 $\pm$ 1.8e-05 8.1e-04 $\pm$ 2.2e-05 7.8e-04 $\pm$ 2.2e-05 8.0e-04 $\pm$ 1.8e-05 9.1e-04 $\pm$ 2.9e-05 7.4e-04 $\pm$ 1.8e-05 7.4e-04 $\pm$ 1.8e-05
airfoil 8.717 $\pm$ 0.870 14.669 $\pm$ 1.007 9.843 $\pm$ 1.249 9.397 $\pm$ 0.977 11.761 $\pm$ 1.548 8.727 $\pm$ 0.902 8.730 $\pm$ 0.892
allstate 9.2e+03 $\pm$ 5.9e+02 1.1e+04 $\pm$ 6.7e+02 1.0e+04 $\pm$ 6.4e+02 1.0e+04 $\pm$ 6.8e+02 1.3e+04 $\pm$ 8.6e+02 9.3e+03 $\pm$ 6.0e+02 9.3e+03 $\pm$ 6.2e+02
ca_housing 2.2e+05 $\pm$ 4.8e+03 3.9e+05 $\pm$ 2.1e+03 2.4e+05 $\pm$ 5.9e+03 2.6e+05 $\pm$ 7.5e+03 3.2e+05 $\pm$ 9.7e+03 2.2e+05 $\pm$ 5.3e+03 2.2e+05 $\pm$ 5.6e+03
computer 12.707 $\pm$ 0.575 17.102 $\pm$ 3.298 16.503 $\pm$ 1.199 16.547 $\pm$ 0.692 18.684 $\pm$ 1.517 12.707 $\pm$ 0.591 12.897 $\pm$ 0.602
concrete 27.020 $\pm$ 3.202 43.690 $\pm$ 4.842 28.684 $\pm$ 4.251 26.165 $\pm$ 2.857 30.054 $\pm$ 4.640 26.365 $\pm$ 2.976 26.445 $\pm$ 2.996
elevator 0.011 $\pm$ 0.000 0.014 $\pm$ 0.000 0.013 $\pm$ 0.000 0.015 $\pm$ 0.000 0.017 $\pm$ 0.000 0.012 $\pm$ 0.000 0.012 $\pm$ 0.000
energy_efficiency 2.465 $\pm$ 0.347 8.199 $\pm$ 0.672 2.671 $\pm$ 0.416 3.058 $\pm$ 0.217 3.253 $\pm$ 0.388 2.667 $\pm$ 0.280 2.721 $\pm$ 0.270
insurance 2.6e+04 $\pm$ 2.5e+03 3.1e+04 $\pm$ 3.0e+03 3.0e+04 $\pm$ 2.8e+03 3.9e+04 $\pm$ 7.1e+03 3.5e+04 $\pm$ 2.7e+03 3.0e+04 $\pm$ 5.0e+03 2.8e+04 $\pm$ 3.2e+03
kin8nm 0.577 $\pm$ 0.014 0.760 $\pm$ 0.013 0.616 $\pm$ 0.021 0.607 $\pm$ 0.008 0.666 $\pm$ 0.020 0.577 $\pm$ 0.013 0.577 $\pm$ 0.013
miami_housing 3.1e+05 $\pm$ 1.5e+04 4.1e+05 $\pm$ 2.5e+04 4.2e+05 $\pm$ 2.7e+04 3.3e+05 $\pm$ 1.3e+04 6.5e+05 $\pm$ 2.9e+04 3.1e+05 $\pm$ 1.5e+04 3.1e+05 $\pm$ 1.5e+04
naval_propulsion 4.4e-03 $\pm$ 1.5e-04 0.391 $\pm$ 0.006 5.1e-03 $\pm$ 2.2e-04 0.014 $\pm$ 0.001 7.1e-03 $\pm$ 4.7e-04 4.8e-03 $\pm$ 1.4e-04 6.0e-03 $\pm$ 2.2e-04
parkinsons 16.697 $\pm$ 0.794 25.497 $\pm$ 0.692 18.270 $\pm$ 1.122 18.931 $\pm$ 0.600 21.220 $\pm$ 1.011 16.688 $\pm$ 0.771 16.680 $\pm$ 0.750
powerplant 16.993 $\pm$ 0.937 18.211 $\pm$ 0.897 17.577 $\pm$ 0.884 17.715 $\pm$ 0.965 18.501 $\pm$ 0.815 17.064 $\pm$ 0.951 17.027 $\pm$ 0.954
qsar 3.956 $\pm$ 0.218 4.802 $\pm$ 0.219 4.193 $\pm$ 0.227 4.270 $\pm$ 0.220 4.562 $\pm$ 0.217 3.946 $\pm$ 0.222 3.946 $\pm$ 0.223
sulfur 0.155 $\pm$ 0.008 0.174 $\pm$ 0.009 0.186 $\pm$ 0.012 0.180 $\pm$ 0.007 0.277 $\pm$ 0.020 0.156 $\pm$ 0.007 0.155 $\pm$ 0.007
superconductor 41.801 $\pm$ 1.407 51.826 $\pm$ 1.568 45.860 $\pm$ 1.883 48.899 $\pm$ 2.037 71.323 $\pm$ 2.803 41.747 $\pm$ 1.460 41.821 $\pm$ 1.571
F.3 Variant (c)
Figure 8: Quantile loss and NCIW performance of different methods (CLEAR, PCS, ALEATORIC, ALEATORIC-R) for variant (c), over 10 seeds normalized relative to CLEAR (baseline = 1.0). Lower values indicate better performance. The inset boxplot shows the % improvement relative to the CLEAR baseline $\pm 1\sigma$. Values inside each subplot represent the mean improvement across all datasets.
Table 22: Variant (c) PICP at 95% prediction intervals, aggregated across 10 seeds.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
airfoil 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
allstate 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
computer 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
concrete 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
elevator 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
energy_efficiency 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.95 $\pm$ 0.01 0.97 $\pm$ 0.01 0.96 $\pm$ 0.01
insurance 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
kin8nm 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
miami_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
naval_propulsion 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
parkinsons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
powerplant 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
qsar 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
Table 23: Variant (c) NIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.208 $\pm$ 0.017 0.227 $\pm$ 0.017 0.202 $\pm$ 0.018 0.237 $\pm$ 0.024 0.229 $\pm$ 0.020 0.208 $\pm$ 0.016 0.208 $\pm$ 0.017
airfoil 0.199 $\pm$ 0.013 0.286 $\pm$ 0.016 0.223 $\pm$ 0.024 0.204 $\pm$ 0.019 0.214 $\pm$ 0.024 0.205 $\pm$ 0.018 0.202 $\pm$ 0.014
allstate 0.279 $\pm$ 0.052 0.317 $\pm$ 0.065 0.277 $\pm$ 0.047 0.299 $\pm$ 0.065 0.315 $\pm$ 0.059 0.274 $\pm$ 0.053 0.272 $\pm$ 0.051
ca_housing 0.315 $\pm$ 0.010 0.425 $\pm$ 0.005 0.345 $\pm$ 0.013 0.332 $\pm$ 0.011 0.400 $\pm$ 0.016 0.312 $\pm$ 0.008 0.311 $\pm$ 0.008
computer 0.080 $\pm$ 0.004 0.102 $\pm$ 0.001 0.089 $\pm$ 0.004 0.084 $\pm$ 0.004 0.091 $\pm$ 0.006 0.079 $\pm$ 0.004 0.078 $\pm$ 0.004
concrete 0.278 $\pm$ 0.029 0.315 $\pm$ 0.019 0.284 $\pm$ 0.027 0.273 $\pm$ 0.031 0.276 $\pm$ 0.028 0.279 $\pm$ 0.029 0.281 $\pm$ 0.029
elevator 0.127 $\pm$ 0.008 0.192 $\pm$ 0.010 0.137 $\pm$ 0.008 0.152 $\pm$ 0.008 0.144 $\pm$ 0.010 0.127 $\pm$ 0.007 0.127 $\pm$ 0.007
energy_efficiency 0.043 $\pm$ 0.007 0.138 $\pm$ 0.007 0.051 $\pm$ 0.005 0.045 $\pm$ 0.007 0.049 $\pm$ 0.005 0.047 $\pm$ 0.007 0.046 $\pm$ 0.007
insurance 0.397 $\pm$ 0.091 0.421 $\pm$ 0.049 0.459 $\pm$ 0.050 0.395 $\pm$ 0.096 0.423 $\pm$ 0.083 0.440 $\pm$ 0.092 0.475 $\pm$ 0.109
kin8nm 0.325 $\pm$ 0.012 0.415 $\pm$ 0.018 0.330 $\pm$ 0.014 0.329 $\pm$ 0.011 0.344 $\pm$ 0.012 0.328 $\pm$ 0.014 0.326 $\pm$ 0.013
miami_housing 0.088 $\pm$ 0.001 0.112 $\pm$ 0.002 0.113 $\pm$ 0.008 0.093 $\pm$ 0.002 0.123 $\pm$ 0.008 0.078 $\pm$ 0.002 0.078 $\pm$ 0.002
naval_propulsion 1.6e-03 $\pm$ 3.2e-05 4.1e-03 $\pm$ 8.9e-05 1.6e-03 $\pm$ 4.3e-05 1.7e-03 $\pm$ 5.4e-05 1.8e-03 $\pm$ 3.9e-05 1.6e-03 $\pm$ 1.9e-05 1.6e-03 $\pm$ 3.0e-05
parkinsons 0.250 $\pm$ 0.007 0.371 $\pm$ 0.011 0.281 $\pm$ 0.011 0.260 $\pm$ 0.008 0.285 $\pm$ 0.010 0.251 $\pm$ 0.008 0.248 $\pm$ 0.007
powerplant 0.157 $\pm$ 0.007 0.175 $\pm$ 0.008 0.161 $\pm$ 0.008 0.169 $\pm$ 0.006 0.164 $\pm$ 0.007 0.158 $\pm$ 0.007 0.158 $\pm$ 0.007
qsar 0.344 $\pm$ 0.117 0.464 $\pm$ 0.153 0.353 $\pm$ 0.117 0.399 $\pm$ 0.138 0.386 $\pm$ 0.131 0.352 $\pm$ 0.120 0.346 $\pm$ 0.118
sulfur 0.111 $\pm$ 0.009 0.124 $\pm$ 0.010 0.109 $\pm$ 0.014 0.124 $\pm$ 0.009 0.114 $\pm$ 0.013 0.105 $\pm$ 0.009 0.104 $\pm$ 0.009
superconductor 0.195 $\pm$ 0.020 0.250 $\pm$ 0.025 0.210 $\pm$ 0.029 0.235 $\pm$ 0.027 0.308 $\pm$ 0.035 0.193 $\pm$ 0.020 0.193 $\pm$ 0.020
Table 24: Variant (c) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 9.6e-06 $\pm$ 2.9e-07 1.0e-05 $\pm$ 2.7e-07 9.9e-06 $\pm$ 3.6e-07 1.1e-05 $\pm$ 2.7e-07 1.2e-05 $\pm$ 3.7e-07 9.6e-06 $\pm$ 2.9e-07 9.6e-06 $\pm$ 2.9e-07
airfoil 0.107 $\pm$ 0.014 0.140 $\pm$ 0.007 0.129 $\pm$ 0.019 0.110 $\pm$ 0.013 0.130 $\pm$ 0.019 0.108 $\pm$ 0.014 0.108 $\pm$ 0.014
allstate 1.2e+02 $\pm$ 6.874 1.2e+02 $\pm$ 4.960 1.2e+02 $\pm$ 19.401 1.4e+02 $\pm$ 9.737 1.6e+02 $\pm$ 10.679 1.2e+02 $\pm$ 7.379 1.2e+02 $\pm$ 7.410
ca_housing 2.8e+03 $\pm$ 92.560 3.2e+03 $\pm$ 71.387 3.2e+03 $\pm$ 1.0e+02 2.9e+03 $\pm$ 87.120 3.6e+03 $\pm$ 1.2e+02 2.8e+03 $\pm$ 88.458 2.8e+03 $\pm$ 87.820
computer 0.137 $\pm$ 0.011 0.152 $\pm$ 0.013 0.160 $\pm$ 0.012 0.140 $\pm$ 0.012 0.164 $\pm$ 0.013 0.137 $\pm$ 0.011 0.138 $\pm$ 0.011
concrete 0.337 $\pm$ 0.032 0.374 $\pm$ 0.044 0.381 $\pm$ 0.059 0.336 $\pm$ 0.032 0.384 $\pm$ 0.056 0.336 $\pm$ 0.032 0.337 $\pm$ 0.032
elevator 1.3e-04 $\pm$ 3.2e-06 1.7e-04 $\pm$ 2.2e-06 1.5e-04 $\pm$ 3.3e-06 1.4e-04 $\pm$ 3.7e-06 1.6e-04 $\pm$ 3.1e-06 1.3e-04 $\pm$ 2.9e-06 1.3e-04 $\pm$ 2.9e-06
energy_efficiency 0.029 $\pm$ 0.007 0.068 $\pm$ 0.005 0.034 $\pm$ 0.010 0.030 $\pm$ 0.007 0.034 $\pm$ 0.010 0.030 $\pm$ 0.007 0.030 $\pm$ 0.007
insurance 4.2e+02 $\pm$ 80.269 3.9e+02 $\pm$ 33.576 4.0e+02 $\pm$ 53.283 4.4e+02 $\pm$ 65.348 4.1e+02 $\pm$ 41.712 4.3e+02 $\pm$ 77.643 4.5e+02 $\pm$ 67.994
kin8nm 6.7e-03 $\pm$ 1.8e-04 8.4e-03 $\pm$ 2.6e-04 6.8e-03 $\pm$ 1.9e-04 6.9e-03 $\pm$ 2.2e-04 7.4e-03 $\pm$ 2.8e-04 6.7e-03 $\pm$ 1.6e-04 6.7e-03 $\pm$ 1.7e-04
miami_housing 3.7e+03 $\pm$ 1.4e+02 4.2e+03 $\pm$ 1.2e+02 7.3e+03 $\pm$ 4.2e+02 3.7e+03 $\pm$ 1.2e+02 7.7e+03 $\pm$ 4.2e+02 4.0e+03 $\pm$ 2.4e+02 4.0e+03 $\pm$ 2.3e+02
naval_propulsion 4.0e-05 $\pm$ 1.0e-06 9.5e-05 $\pm$ 2.1e-06 4.9e-05 $\pm$ 1.6e-06 4.2e-05 $\pm$ 1.2e-06 5.7e-05 $\pm$ 1.2e-06 4.1e-05 $\pm$ 1.2e-06 4.5e-05 $\pm$ 1.6e-06
parkinsons 0.189 $\pm$ 0.006 0.249 $\pm$ 0.007 0.216 $\pm$ 0.008 0.191 $\pm$ 0.006 0.220 $\pm$ 0.008 0.189 $\pm$ 0.006 0.190 $\pm$ 0.007
powerplant 0.203 $\pm$ 0.012 0.213 $\pm$ 0.014 0.208 $\pm$ 0.012 0.211 $\pm$ 0.013 0.211 $\pm$ 0.012 0.203 $\pm$ 0.012 0.203 $\pm$ 0.012
qsar 0.048 $\pm$ 0.002 0.057 $\pm$ 0.003 0.051 $\pm$ 0.002 0.054 $\pm$ 0.002 0.055 $\pm$ 0.003 0.049 $\pm$ 0.002 0.049 $\pm$ 0.002
sulfur 1.8e-03 $\pm$ 6.3e-05 1.9e-03 $\pm$ 1.2e-04 2.7e-03 $\pm$ 1.6e-04 1.8e-03 $\pm$ 6.2e-05 2.9e-03 $\pm$ 1.7e-04 1.8e-03 $\pm$ 6.8e-05 1.8e-03 $\pm$ 7.1e-05
superconductor 0.489 $\pm$ 0.020 0.553 $\pm$ 0.018 0.568 $\pm$ 0.042 0.568 $\pm$ 0.025 0.792 $\pm$ 0.035 0.490 $\pm$ 0.020 0.489 $\pm$ 0.021
Table 25: Variant (c) NCIW at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 0.210 $\pm$ 0.020 0.231 $\pm$ 0.023 0.204 $\pm$ 0.021 0.239 $\pm$ 0.021 0.230 $\pm$ 0.021 0.209 $\pm$ 0.020 0.210 $\pm$ 0.019
airfoil 0.188 $\pm$ 0.014 0.258 $\pm$ 0.034 0.204 $\pm$ 0.017 0.199 $\pm$ 0.013 0.204 $\pm$ 0.014 0.190 $\pm$ 0.012 0.188 $\pm$ 0.010
allstate 0.277 $\pm$ 0.056 0.316 $\pm$ 0.065 0.281 $\pm$ 0.041 0.305 $\pm$ 0.070 0.319 $\pm$ 0.051 0.271 $\pm$ 0.054 0.270 $\pm$ 0.053
ca_housing 0.311 $\pm$ 0.010 0.419 $\pm$ 0.014 0.343 $\pm$ 0.010 0.329 $\pm$ 0.009 0.401 $\pm$ 0.018 0.309 $\pm$ 0.009 0.309 $\pm$ 0.009
computer 0.081 $\pm$ 0.004 0.104 $\pm$ 0.007 0.090 $\pm$ 0.004 0.084 $\pm$ 0.005 0.092 $\pm$ 0.005 0.079 $\pm$ 0.004 0.080 $\pm$ 0.004
concrete 0.259 $\pm$ 0.032 0.303 $\pm$ 0.054 0.263 $\pm$ 0.029 0.257 $\pm$ 0.031 0.265 $\pm$ 0.027 0.258 $\pm$ 0.033 0.259 $\pm$ 0.032
elevator 0.126 $\pm$ 0.008 0.192 $\pm$ 0.008 0.138 $\pm$ 0.008 0.151 $\pm$ 0.011 0.144 $\pm$ 0.010 0.127 $\pm$ 0.007 0.127 $\pm$ 0.007
energy_efficiency 0.041 $\pm$ 0.006 0.116 $\pm$ 0.015 0.048 $\pm$ 0.005 0.043 $\pm$ 0.005 0.049 $\pm$ 0.005 0.042 $\pm$ 0.006 0.042 $\pm$ 0.006
insurance 0.345 $\pm$ 0.078 0.413 $\pm$ 0.043 0.424 $\pm$ 0.074 0.343 $\pm$ 0.087 0.385 $\pm$ 0.112 0.367 $\pm$ 0.100 0.407 $\pm$ 0.137
kin8nm 0.323 $\pm$ 0.008 0.420 $\pm$ 0.018 0.327 $\pm$ 0.008 0.331 $\pm$ 0.010 0.341 $\pm$ 0.012 0.325 $\pm$ 0.009 0.324 $\pm$ 0.009
miami_housing 0.088 $\pm$ 0.002 0.113 $\pm$ 0.004 0.111 $\pm$ 0.007 0.094 $\pm$ 0.003 0.122 $\pm$ 0.007 0.077 $\pm$ 0.003 0.076 $\pm$ 0.003
naval_propulsion 1.6e-03 $\pm$ 5.0e-05 4.0e-03 $\pm$ 2.4e-04 1.6e-03 $\pm$ 4.1e-05 1.7e-03 $\pm$ 4.7e-05 1.8e-03 $\pm$ 5.8e-05 1.6e-03 $\pm$ 5.2e-05 1.6e-03 $\pm$ 6.2e-05
parkinsons 0.250 $\pm$ 0.012 0.366 $\pm$ 0.017 0.279 $\pm$ 0.013 0.259 $\pm$ 0.009 0.284 $\pm$ 0.013 0.251 $\pm$ 0.011 0.250 $\pm$ 0.011
powerplant 0.161 $\pm$ 0.007 0.176 $\pm$ 0.009 0.163 $\pm$ 0.008 0.171 $\pm$ 0.008 0.166 $\pm$ 0.007 0.161 $\pm$ 0.007 0.161 $\pm$ 0.007
qsar 0.348 $\pm$ 0.118 0.465 $\pm$ 0.156 0.361 $\pm$ 0.124 0.407 $\pm$ 0.142 0.395 $\pm$ 0.134 0.359 $\pm$ 0.124 0.354 $\pm$ 0.122
sulfur 0.109 $\pm$ 0.008 0.124 $\pm$ 0.008 0.108 $\pm$ 0.006 0.124 $\pm$ 0.008 0.115 $\pm$ 0.007 0.104 $\pm$ 0.007 0.102 $\pm$ 0.007
superconductor 0.192 $\pm$ 0.019 0.251 $\pm$ 0.025 0.206 $\pm$ 0.027 0.232 $\pm$ 0.025 0.297 $\pm$ 0.029 0.191 $\pm$ 0.019 0.191 $\pm$ 0.019
Table 26: Variant (c) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR ALEATORIC ALEATORIC-R PCS-EPISTEMIC Naive $\gamma_{1}=1$ $\lambda=1$
ailerons 7.7e-04 $\pm$ 2.3e-05 8.2e-04 $\pm$ 2.2e-05 7.9e-04 $\pm$ 2.9e-05 8.5e-04 $\pm$ 2.2e-05 9.4e-04 $\pm$ 3.0e-05 7.7e-04 $\pm$ 2.3e-05 7.7e-04 $\pm$ 2.3e-05
airfoil 8.569 $\pm$ 1.099 11.217 $\pm$ 0.549 10.341 $\pm$ 1.548 8.799 $\pm$ 1.034 10.361 $\pm$ 1.499 8.675 $\pm$ 1.116 8.624 $\pm$ 1.105
allstate 9.2e+03 $\pm$ 5.5e+02 9.6e+03 $\pm$ 4.0e+02 1.0e+04 $\pm$ 1.6e+03 1.1e+04 $\pm$ 7.8e+02 1.3e+04 $\pm$ 8.5e+02 9.2e+03 $\pm$ 5.9e+02 9.3e+03 $\pm$ 5.9e+02
ca_housing 2.2e+05 $\pm$ 7.4e+03 2.5e+05 $\pm$ 5.7e+03 2.5e+05 $\pm$ 8.1e+03 2.3e+05 $\pm$ 7.0e+03 2.9e+05 $\pm$ 9.7e+03 2.2e+05 $\pm$ 7.1e+03 2.2e+05 $\pm$ 7.0e+03
computer 10.978 $\pm$ 0.911 12.134 $\pm$ 1.001 12.827 $\pm$ 0.945 11.204 $\pm$ 0.996 13.130 $\pm$ 1.051 10.986 $\pm$ 0.873 11.030 $\pm$ 0.866
concrete 26.928 $\pm$ 2.529 29.955 $\pm$ 3.544 30.449 $\pm$ 4.693 26.855 $\pm$ 2.535 30.720 $\pm$ 4.505 26.882 $\pm$ 2.568 26.962 $\pm$ 2.581
elevator 0.010 $\pm$ 0.000 0.014 $\pm$ 0.000 0.012 $\pm$ 0.000 0.011 $\pm$ 0.000 0.013 $\pm$ 0.000 0.010 $\pm$ 0.000 0.010 $\pm$ 0.000
energy_efficiency 2.348 $\pm$ 0.554 5.407 $\pm$ 0.418 2.759 $\pm$ 0.822 2.397 $\pm$ 0.573 2.756 $\pm$ 0.822 2.392 $\pm$ 0.564 2.395 $\pm$ 0.588
insurance 3.4e+04 $\pm$ 6.4e+03 3.1e+04 $\pm$ 2.7e+03 3.2e+04 $\pm$ 4.3e+03 3.5e+04 $\pm$ 5.2e+03 3.3e+04 $\pm$ 3.3e+03 3.4e+04 $\pm$ 6.2e+03 3.6e+04 $\pm$ 5.4e+03
kin8nm 0.534 $\pm$ 0.014 0.673 $\pm$ 0.021 0.544 $\pm$ 0.015 0.552 $\pm$ 0.017 0.591 $\pm$ 0.022 0.536 $\pm$ 0.013 0.535 $\pm$ 0.013
miami_housing 2.9e+05 $\pm$ 1.1e+04 3.3e+05 $\pm$ 9.5e+03 5.9e+05 $\pm$ 3.4e+04 3.0e+05 $\pm$ 9.6e+03 6.1e+05 $\pm$ 3.4e+04 3.2e+05 $\pm$ 1.9e+04 3.2e+05 $\pm$ 1.9e+04
naval_propulsion 3.2e-03 $\pm$ 8.4e-05 7.6e-03 $\pm$ 1.7e-04 3.9e-03 $\pm$ 1.3e-04 3.4e-03 $\pm$ 9.6e-05 4.6e-03 $\pm$ 9.6e-05 3.2e-03 $\pm$ 9.3e-05 3.6e-03 $\pm$ 1.3e-04
parkinsons 15.149 $\pm$ 0.518 19.899 $\pm$ 0.576 17.316 $\pm$ 0.634 15.312 $\pm$ 0.459 17.596 $\pm$ 0.621 15.150 $\pm$ 0.499 15.161 $\pm$ 0.537
powerplant 16.243 $\pm$ 0.985 17.062 $\pm$ 1.128 16.673 $\pm$ 0.987 16.846 $\pm$ 1.029 16.874 $\pm$ 0.975 16.231 $\pm$ 0.984 16.231 $\pm$ 0.983
qsar 3.877 $\pm$ 0.176 4.521 $\pm$ 0.225 4.086 $\pm$ 0.193 4.334 $\pm$ 0.125 4.390 $\pm$ 0.205 3.927 $\pm$ 0.146 3.892 $\pm$ 0.151
sulfur 0.143 $\pm$ 0.005 0.156 $\pm$ 0.010 0.213 $\pm$ 0.013 0.146 $\pm$ 0.005 0.228 $\pm$ 0.013 0.143 $\pm$ 0.005 0.145 $\pm$ 0.006
superconductor 39.104 $\pm$ 1.604 44.230 $\pm$ 1.428 45.479 $\pm$ 3.358 45.456 $\pm$ 2.039 63.347 $\pm$ 2.824 39.224 $\pm$ 1.613 39.119 $\pm$ 1.642
Table 27: Variant (c) CLEAR calibration parameters $\lambda$ and $\gamma_{1}$ for 95% prediction intervals across 10 seeds. Using all available variables. Showing median [min:max] values.
Dataset $\lambda$ $\gamma_{1}$
ailerons 0.89 [0.17:4.47] 0.98 [0.87:1.58]
airfoil 0.76 [0.23:5.25] 1.70 [0.25:4.30]
allstate 0.21 [0.00:14.06] 1.14 [0.16:1.93]
ca_housing 1.90 [1.14:5.08] 0.70 [0.33:0.93]
computer 2.21 [1.63:8.87] 0.74 [0.21:0.96]
concrete 1.33 [0.06:100.00] 1.65 [0.01:16.73]
elevator 0.94 [0.67:1.43] 1.04 [0.82:1.22]
energy_efficiency 0.12 [0.01:100.00] 7.96 [0.02:24.25]
insurance 5.07 [0.14:100.00] 0.38 [0.02:27.45]
kin8nm 1.87 [1.64:3.82] 0.76 [0.58:0.83]
miami_housing 7.06 [4.84:19.45] 0.23 [0.09:0.33]
naval_propulsion 16.16 [12.48:21.17] 0.63 [0.53:0.76]
parkinsons 1.16 [0.72:8.42] 1.26 [0.22:1.84]
powerplant 1.05 [0.73:3.26] 1.05 [0.51:1.36]
qsar 0.51 [0.32:1.70] 1.45 [0.96:2.37]
sulfur 2.41 [1.53:7.16] 0.71 [0.29:0.91]
superconductor 0.91 [0.34:1.39] 1.07 [0.80:1.34]
F.4 Comparing Variants of CLEAR
Figure 9: Quantile loss and NCIW performance of different variants of CLEAR (a, b and c) over 10 seeds normalized relative to CLEAR (a) (baseline = 1.0). Lower values indicate better performance. The inset boxplot shows the % improvement relative to the CLEAR (a) baseline $\pm 1\sigma$. Values inside each subplot represent the mean improvement across all datasets.
F.5 Runtime
The grid search for finding the optimal parameter is extremely fast and negligible compared to the baselines, despite using a grid with 4000 points, which is finer and larger than necessary. In Tables 28, 29 and 30, we provide the average training time for each variant on a given real-world dataset, computed on a machine with an Intel® Core i9-13900KF CPU (with maximum 20+ threads used in parallel). For experiments involving SQR and DEs, we had access to an NVIDIA® GeForce RTX 4090 GPU. All results are computed using 32 GB of memory. The times are provided in seconds and have been averaged over 10 seeds. For CLEAR, the required computation is the calibration time (computation of $\lambda$ and $\gamma_{1}$, denoted as Grid Search in the tables). Only the runtimes for the methods included in the paper are provided. The experiments were run exactly as described in the paper (particularly for the 100 bootstraps).
As stated previously, CLEAR is highly modular, allowing the modules and their components (such as the choice of base models within the PCS module) to be replaced or modified to adhere to computational budget limitations. For example, if the dataset is very large, base models that scale well with the training data size can be used (such as deep learning models, trained via stochastic gradient descent). Alternatively, if training each model is expensive, one might want to use fewer than 100 different bootstraps per dataset, or maybe even use techniques (such as Monte-Carlo Dropout) to obtain an ensemble from a single trained model [17, 26, 48, 49, 19, 50, 30].
Note that ALEATORIC-R cannot be computed without first computing PCS, as it utilizes the residuals from PCS. In practice, the total computational costs of ALEATORIC-R would be equal to the costs of PCS plus the additional costs of the residual approach. Thus, the minimal total computational costs necessary to obtain CLEAR results are equal to the minimal total computational costs required to obtain ALEATORIC-R results (up to a few seconds of the grid search).
Remark F.1 (Parallelization and Scalability).
Every step in the entire CLEAR pipeline is parallelizable and distributable. The most expensive part is fitting multiple models to multiple bootstraps of the data for PCS. However, these models can be trained perfectly in parallel on different distributed nodes, as they do not have to communicate with each other. Similar parallelization (distributed) is also possible for ALEATORIC-R, where multiple models are fitted independently. The computational costs of CLEAR’s calibration step (Grid-Search), albeit negligible, could be further reduced by distributing the exploration of the grid across multiple servers. The computations necessary for each grid point are fully vectorized and are highly suited for GPUs (in the case of a large calibration dataset). In other words, regardless of whether the user has access to many weak CPU servers, a powerful GPU, or multiple powerful GPU servers, the calibration step can efficiently utilize all these different infrastructures (when tuning the implementation accordingly). For all these reasons, CLEAR is highly scalable.
Table 28: Variant (a) average runtime (over 10 seeds) in seconds for base components and the CLEAR’s grid search. The grid search includes any other overhead for CLEAR.
Dataset PCS ALEATORIC-R Grid-Search Total
ailerons 13.46 6.80 0.27 20.53
airfoil 3.10 1.92 0.17 5.19
allstate 1124.47 22.93 0.19 1147.59
ca_housing 8.51 7.79 0.54 16.84
computer 61.92 76.09 0.16 138.16
concrete 1.10 2.16 0.17 3.43
elevator 18.79 122.16 0.60 141.54
energy_efficiency 0.81 1.65 0.16 2.62
insurance 3.48 17.23 0.11 20.82
kin8nm 3.25 5.43 0.21 8.89
miami_housing 7.31 7.66 0.72 15.68
naval_propulsion 19.83 81.32 0.29 101.44
parkinsons 39.53 60.99 0.14 100.66
powerplant 3.08 4.92 0.21 8.22
qsar 348.59 10.00 0.20 358.78
sulfur 21.55 16.56 0.21 38.32
superconductor 561.84 795.45 12.70 1369.99
Total 2240.60 1241.06 17.06 3498.71
Table 29: Variant (b) average runtime (over 10 seeds) in seconds for base components and the CLEAR’s grid search. The grid search includes any other overhead for CLEAR.
Dataset PCS ALEATORIC-R Grid-Search Total
ailerons 4.51 6.89 0.26 11.66
airfoil 0.78 1.64 0.16 2.59
allstate 27.73 14.22 0.20 42.15
ca_housing 5.42 7.69 0.39 13.50
computer 3.46 5.97 0.21 9.65
concrete 0.86 1.82 0.17 2.85
elevator 4.43 6.71 0.79 11.93
energy_efficiency 0.68 1.40 0.17 2.25
insurance 0.80 1.65 0.17 2.62
kin8nm 2.61 4.96 0.21 7.78
miami_housing 4.61 7.70 0.70 13.01
naval_propulsion 3.72 6.38 0.55 10.64
parkinsons 2.71 5.37 0.20 8.28
powerplant 2.73 4.32 0.22 7.26
qsar 5.00 9.67 0.20 14.86
sulfur 2.89 4.64 0.23 7.76
superconductor 13.81 22.74 1.00 37.55
Total 86.77 113.76 5.81 206.34
Table 30: Variant (c) average runtime (over 10 seeds) in seconds for base components and the CLEAR’s grid search. The grid search includes any other overhead for CLEAR.
Dataset PCS ALEATORIC-R Grid-Search Total
ailerons 10.20 104.32 0.65 115.17
airfoil 0.88 17.82 0.11 18.81
allstate 651.02 125.77 0.14 776.93
ca_housing 3.63 130.87 4.16 138.66
computer 4.55 65.28 0.16 69.98
concrete 1.16 17.49 0.11 18.76
elevator 3.25 98.73 1.97 103.95
energy_efficiency 0.95 15.42 0.10 16.47
insurance 0.90 19.80 0.11 20.81
kin8nm 16.89 69.17 0.16 86.22
miami_housing 5.34 123.32 0.92 129.58
naval_propulsion 1.14 115.41 0.74 117.28
parkinsons 3.71 53.43 0.14 57.27
powerplant 2.37 60.99 0.17 63.52
qsar 21.48 93.84 0.14 115.46
sulfur 2.43 67.41 0.17 70.01
superconductor 60.63 644.13 8.10 712.86
Total 790.52 1823.18 18.05 2631.75
Appendix G Conformalized CLEAR with PCS and CQR: Results on Real-World Data
This section presents results from our conformalized experimental configuration, where we split the 20% validation set into separate 10% validation and 10% calibration sets. This approach provides stronger finite-sample distribution-free marginal coverage guarantees following conformal prediction principles, as the calibration set remains completely unseen during model selection and hyperparameter optimization. While the conformalized approach may sacrifice some performance due to reduced data availability for validation, it offers theoretical rigor by ensuring the conformal calibration step operates on the held-out data. Similar to the standard results, CLEAR adapts to this more stringent experimental setting while maintaining its advantages over baseline methods across all three variants (a), (b), and (c).
G.1 Variant (a) Conformalized
Table 31: Conformalized Variant (a) PICP at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00
airfoil 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01
allstate 0.96 $\pm$ 0.01 0.94 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
computer 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01
concrete 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.94 $\pm$ 0.02 0.95 $\pm$ 0.02
elevator 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00
energy_efficiency 0.97 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02
insurance 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02
kin8nm 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
miami_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
naval_propulsion 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
parkinsons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
powerplant 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
qsar 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00
Table 32: Conformalized Variant (a) NIW at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 0.193 $\pm$ 0.016 0.217 $\pm$ 0.017 0.195 $\pm$ 0.018 0.193 $\pm$ 0.016
airfoil 0.215 $\pm$ 0.019 0.217 $\pm$ 0.022 0.227 $\pm$ 0.031 0.207 $\pm$ 0.012
allstate 0.267 $\pm$ 0.041 0.263 $\pm$ 0.060 0.259 $\pm$ 0.046 0.260 $\pm$ 0.037
ca_housing 0.339 $\pm$ 0.006 0.355 $\pm$ 0.012 0.349 $\pm$ 0.008 0.339 $\pm$ 0.008
computer 0.091 $\pm$ 0.006 8.020 $\pm$ 11.958 0.094 $\pm$ 0.016 0.091 $\pm$ 0.008
concrete 0.272 $\pm$ 0.030 0.260 $\pm$ 0.050 0.266 $\pm$ 0.038 0.249 $\pm$ 0.025
elevator 0.157 $\pm$ 0.009 0.177 $\pm$ 0.014 0.149 $\pm$ 0.008 0.156 $\pm$ 0.009
energy_efficiency 0.053 $\pm$ 0.009 0.058 $\pm$ 0.009 0.053 $\pm$ 0.009 0.047 $\pm$ 0.005
insurance 0.338 $\pm$ 0.063 0.535 $\pm$ 0.315 0.379 $\pm$ 0.089 0.309 $\pm$ 0.032
kin8nm 0.362 $\pm$ 0.014 0.373 $\pm$ 0.013 0.370 $\pm$ 0.011 0.360 $\pm$ 0.012
miami_housing 0.085 $\pm$ 0.003 0.097 $\pm$ 0.002 0.087 $\pm$ 0.005 0.085 $\pm$ 0.002
naval_propulsion 6.2e-04 $\pm$ 5.9e-06 7.1e-04 $\pm$ 1.5e-05 6.0e-04 $\pm$ 9.6e-06 6.1e-04 $\pm$ 6.3e-06
parkinsons 0.227 $\pm$ 0.017 0.315 $\pm$ 0.034 0.253 $\pm$ 0.020 0.227 $\pm$ 0.010
powerplant 0.171 $\pm$ 0.010 0.179 $\pm$ 0.008 0.183 $\pm$ 0.008 0.170 $\pm$ 0.007
qsar 0.369 $\pm$ 0.123 0.388 $\pm$ 0.132 0.381 $\pm$ 0.125 0.363 $\pm$ 0.121
sulfur 0.112 $\pm$ 0.009 0.136 $\pm$ 0.014 0.108 $\pm$ 0.010 0.109 $\pm$ 0.010
superconductor 0.197 $\pm$ 0.022 0.249 $\pm$ 0.032 0.195 $\pm$ 0.023 0.196 $\pm$ 0.022
Table 33: Conformalized Variant (a) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 9.2e-06 $\pm$ 2.3e-07 1.0e-05 $\pm$ 2.1e-07 9.7e-06 $\pm$ 2.7e-07 9.2e-06 $\pm$ 2.3e-07
airfoil 0.110 $\pm$ 0.013 0.119 $\pm$ 0.012 0.126 $\pm$ 0.015 0.109 $\pm$ 0.011
allstate 1.1e+02 $\pm$ 8.530 1.3e+02 $\pm$ 8.381 1.3e+02 $\pm$ 7.926 1.1e+02 $\pm$ 7.835
ca_housing 2.8e+03 $\pm$ 56.920 3.2e+03 $\pm$ 91.466 3.0e+03 $\pm$ 73.272 2.8e+03 $\pm$ 59.700
computer 0.142 $\pm$ 0.013 9.961 $\pm$ 14.797 0.157 $\pm$ 0.021 0.147 $\pm$ 0.022
concrete 0.342 $\pm$ 0.039 0.337 $\pm$ 0.047 0.355 $\pm$ 0.047 0.338 $\pm$ 0.040
elevator 1.5e-04 $\pm$ 2.0e-06 1.7e-04 $\pm$ 6.3e-06 1.5e-04 $\pm$ 2.5e-06 1.5e-04 $\pm$ 2.0e-06
energy_efficiency 0.033 $\pm$ 0.005 0.038 $\pm$ 0.003 0.034 $\pm$ 0.005 0.031 $\pm$ 0.004
insurance 3.5e+02 $\pm$ 55.800 5.8e+02 $\pm$ 1.7e+02 3.8e+02 $\pm$ 49.983 3.5e+02 $\pm$ 61.616
kin8nm 7.2e-03 $\pm$ 1.7e-04 7.6e-03 $\pm$ 1.1e-04 7.7e-03 $\pm$ 2.7e-04 7.2e-03 $\pm$ 1.7e-04
miami_housing 3.9e+03 $\pm$ 2.1e+02 4.2e+03 $\pm$ 1.7e+02 5.2e+03 $\pm$ 3.4e+02 3.9e+03 $\pm$ 1.9e+02
naval_propulsion 1.5e-05 $\pm$ 1.9e-07 1.7e-05 $\pm$ 2.9e-07 1.5e-05 $\pm$ 2.7e-07 1.5e-05 $\pm$ 2.1e-07
parkinsons 0.179 $\pm$ 0.008 0.226 $\pm$ 0.016 0.216 $\pm$ 0.012 0.178 $\pm$ 0.010
powerplant 0.213 $\pm$ 0.011 0.222 $\pm$ 0.011 0.220 $\pm$ 0.011 0.212 $\pm$ 0.012
qsar 0.050 $\pm$ 0.003 0.053 $\pm$ 0.003 0.052 $\pm$ 0.003 0.049 $\pm$ 0.003
sulfur 2.0e-03 $\pm$ 1.7e-04 2.4e-03 $\pm$ 2.2e-04 2.3e-03 $\pm$ 1.6e-04 2.0e-03 $\pm$ 1.4e-04
superconductor 0.490 $\pm$ 0.018 0.609 $\pm$ 0.021 0.538 $\pm$ 0.023 0.490 $\pm$ 0.018
Table 34: Conformalized Variant (a) NCIW at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 0.195 $\pm$ 0.018 0.217 $\pm$ 0.021 0.196 $\pm$ 0.017 0.195 $\pm$ 0.017
airfoil 0.203 $\pm$ 0.007 0.211 $\pm$ 0.007 0.216 $\pm$ 0.011 0.203 $\pm$ 0.006
allstate 0.255 $\pm$ 0.040 0.272 $\pm$ 0.057 0.263 $\pm$ 0.043 0.252 $\pm$ 0.037
ca_housing 0.336 $\pm$ 0.006 0.354 $\pm$ 0.012 0.348 $\pm$ 0.011 0.336 $\pm$ 0.008
computer 0.091 $\pm$ 0.008 0.108 $\pm$ 0.016 0.090 $\pm$ 0.008 0.091 $\pm$ 0.008
concrete 0.262 $\pm$ 0.038 0.259 $\pm$ 0.047 0.264 $\pm$ 0.040 0.246 $\pm$ 0.033
elevator 0.155 $\pm$ 0.010 0.177 $\pm$ 0.011 0.148 $\pm$ 0.008 0.155 $\pm$ 0.010
energy_efficiency 0.046 $\pm$ 0.006 0.057 $\pm$ 0.004 0.050 $\pm$ 0.005 0.046 $\pm$ 0.005
insurance 0.299 $\pm$ 0.037 0.480 $\pm$ 0.181 0.312 $\pm$ 0.036 0.303 $\pm$ 0.029
kin8nm 0.355 $\pm$ 0.008 0.371 $\pm$ 0.009 0.371 $\pm$ 0.009 0.354 $\pm$ 0.008
miami_housing 0.084 $\pm$ 0.004 0.098 $\pm$ 0.002 0.085 $\pm$ 0.003 0.084 $\pm$ 0.003
naval_propulsion 6.1e-04 $\pm$ 7.3e-06 7.1e-04 $\pm$ 1.5e-05 6.0e-04 $\pm$ 5.4e-06 6.1e-04 $\pm$ 7.9e-06
parkinsons 0.226 $\pm$ 0.012 0.316 $\pm$ 0.032 0.248 $\pm$ 0.007 0.225 $\pm$ 0.012
powerplant 0.173 $\pm$ 0.008 0.182 $\pm$ 0.007 0.182 $\pm$ 0.007 0.173 $\pm$ 0.007
qsar 0.361 $\pm$ 0.120 0.389 $\pm$ 0.134 0.389 $\pm$ 0.131 0.360 $\pm$ 0.119
sulfur 0.110 $\pm$ 0.007 0.127 $\pm$ 0.008 0.105 $\pm$ 0.005 0.109 $\pm$ 0.010
superconductor 0.195 $\pm$ 0.021 0.247 $\pm$ 0.028 0.194 $\pm$ 0.021 0.195 $\pm$ 0.021
Table 35: Conformalized Variant (a) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 7.4e-04 $\pm$ 1.8e-05 8.1e-04 $\pm$ 1.7e-05 7.8e-04 $\pm$ 2.2e-05 7.4e-04 $\pm$ 1.8e-05
airfoil 8.794 $\pm$ 1.047 9.523 $\pm$ 0.959 10.075 $\pm$ 1.227 8.717 $\pm$ 0.870
allstate 9.2e+03 $\pm$ 6.8e+02 1.0e+04 $\pm$ 6.7e+02 1.0e+04 $\pm$ 6.3e+02 9.1e+03 $\pm$ 6.3e+02
ca_housing 2.2e+05 $\pm$ 4.6e+03 2.6e+05 $\pm$ 7.3e+03 2.4e+05 $\pm$ 5.9e+03 2.2e+05 $\pm$ 4.8e+03
computer 11.388 $\pm$ 1.051 8.0e+02 $\pm$ 1.2e+03 12.535 $\pm$ 1.698 11.741 $\pm$ 1.723
concrete 27.347 $\pm$ 3.094 26.963 $\pm$ 3.739 28.402 $\pm$ 3.793 27.020 $\pm$ 3.202
elevator 0.012 $\pm$ 0.000 0.014 $\pm$ 0.001 0.012 $\pm$ 0.000 0.012 $\pm$ 0.000
energy_efficiency 2.609 $\pm$ 0.417 3.019 $\pm$ 0.237 2.681 $\pm$ 0.431 2.465 $\pm$ 0.347
insurance 2.8e+04 $\pm$ 4.5e+03 4.6e+04 $\pm$ 1.3e+04 3.0e+04 $\pm$ 4.0e+03 2.8e+04 $\pm$ 4.9e+03
kin8nm 0.578 $\pm$ 0.014 0.607 $\pm$ 0.009 0.615 $\pm$ 0.022 0.577 $\pm$ 0.014
miami_housing 3.2e+05 $\pm$ 1.7e+04 3.3e+05 $\pm$ 1.4e+04 4.2e+05 $\pm$ 2.7e+04 3.1e+05 $\pm$ 1.5e+04
naval_propulsion 1.2e-03 $\pm$ 1.6e-05 1.4e-03 $\pm$ 2.3e-05 1.2e-03 $\pm$ 2.1e-05 1.2e-03 $\pm$ 1.7e-05
parkinsons 14.332 $\pm$ 0.644 18.091 $\pm$ 1.263 17.266 $\pm$ 0.988 14.221 $\pm$ 0.792
powerplant 17.066 $\pm$ 0.919 17.724 $\pm$ 0.919 17.567 $\pm$ 0.899 16.993 $\pm$ 0.937
qsar 3.967 $\pm$ 0.222 4.251 $\pm$ 0.212 4.194 $\pm$ 0.234 3.951 $\pm$ 0.230
sulfur 0.161 $\pm$ 0.013 0.189 $\pm$ 0.018 0.182 $\pm$ 0.013 0.161 $\pm$ 0.011
superconductor 39.234 $\pm$ 1.418 48.702 $\pm$ 1.695 43.062 $\pm$ 1.835 39.217 $\pm$ 1.419
G.2 Variant (b) Conformalized
Table 36: Conformalized Variant (b) PICP at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00
airfoil 0.96 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01
allstate 0.95 $\pm$ 0.01 0.94 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
computer 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
concrete 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02 0.95 $\pm$ 0.02
elevator 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00
energy_efficiency 0.97 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02
insurance 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.01
kin8nm 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
miami_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
naval_propulsion 0.95 $\pm$ 0.01 1.00 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
parkinsons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
powerplant 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
qsar 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00
Table 37: Conformalized Variant (b) NIW at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 0.193 $\pm$ 0.016 0.217 $\pm$ 0.017 0.195 $\pm$ 0.018 0.193 $\pm$ 0.016
airfoil 0.215 $\pm$ 0.019 0.217 $\pm$ 0.022 0.227 $\pm$ 0.031 0.207 $\pm$ 0.012
allstate 0.257 $\pm$ 0.044 0.263 $\pm$ 0.057 0.258 $\pm$ 0.046 0.256 $\pm$ 0.045
ca_housing 0.339 $\pm$ 0.006 0.355 $\pm$ 0.012 0.349 $\pm$ 0.008 0.339 $\pm$ 0.008
computer 0.099 $\pm$ 0.006 0.142 $\pm$ 0.010 0.104 $\pm$ 0.005 0.099 $\pm$ 0.006
concrete 0.268 $\pm$ 0.023 0.247 $\pm$ 0.040 0.271 $\pm$ 0.044 0.249 $\pm$ 0.025
elevator 0.137 $\pm$ 0.009 0.194 $\pm$ 0.012 0.145 $\pm$ 0.008 0.137 $\pm$ 0.007
energy_efficiency 0.053 $\pm$ 0.009 0.058 $\pm$ 0.009 0.053 $\pm$ 0.009 0.047 $\pm$ 0.005
insurance 0.357 $\pm$ 0.070 0.461 $\pm$ 0.292 0.408 $\pm$ 0.102 0.329 $\pm$ 0.045
kin8nm 0.362 $\pm$ 0.014 0.373 $\pm$ 0.013 0.370 $\pm$ 0.011 0.360 $\pm$ 0.012
miami_housing 0.085 $\pm$ 0.003 0.097 $\pm$ 0.002 0.087 $\pm$ 0.005 0.085 $\pm$ 0.002
naval_propulsion 1.9e-03 $\pm$ 9.1e-05 8.1e-03 $\pm$ 4.1e-04 1.9e-03 $\pm$ 1.0e-04 1.9e-03 $\pm$ 7.3e-05
parkinsons 0.279 $\pm$ 0.014 0.325 $\pm$ 0.016 0.305 $\pm$ 0.012 0.281 $\pm$ 0.009
powerplant 0.171 $\pm$ 0.010 0.179 $\pm$ 0.008 0.183 $\pm$ 0.008 0.170 $\pm$ 0.007
qsar 0.371 $\pm$ 0.126 0.395 $\pm$ 0.134 0.386 $\pm$ 0.127 0.366 $\pm$ 0.123
sulfur 0.111 $\pm$ 0.010 0.129 $\pm$ 0.016 0.106 $\pm$ 0.011 0.108 $\pm$ 0.009
superconductor 0.220 $\pm$ 0.025 0.240 $\pm$ 0.027 0.227 $\pm$ 0.025 0.219 $\pm$ 0.023
Table 38: Conformalized Variant (b) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 9.2e-06 $\pm$ 2.3e-07 1.0e-05 $\pm$ 2.1e-07 9.7e-06 $\pm$ 2.7e-07 9.2e-06 $\pm$ 2.3e-07
airfoil 0.110 $\pm$ 0.013 0.119 $\pm$ 0.012 0.126 $\pm$ 0.015 0.109 $\pm$ 0.011
allstate 1.2e+02 $\pm$ 7.662 1.3e+02 $\pm$ 8.290 1.3e+02 $\pm$ 7.774 1.1e+02 $\pm$ 7.318
ca_housing 2.8e+03 $\pm$ 56.920 3.2e+03 $\pm$ 91.466 3.0e+03 $\pm$ 73.272 2.8e+03 $\pm$ 59.700
computer 0.159 $\pm$ 0.007 0.207 $\pm$ 0.009 0.206 $\pm$ 0.015 0.159 $\pm$ 0.007
concrete 0.340 $\pm$ 0.037 0.327 $\pm$ 0.039 0.356 $\pm$ 0.048 0.338 $\pm$ 0.040
elevator 1.4e-04 $\pm$ 2.4e-06 1.9e-04 $\pm$ 6.7e-06 1.6e-04 $\pm$ 3.0e-06 1.4e-04 $\pm$ 2.0e-06
energy_efficiency 0.033 $\pm$ 0.005 0.038 $\pm$ 0.003 0.034 $\pm$ 0.005 0.031 $\pm$ 0.004
insurance 3.3e+02 $\pm$ 28.437 5.2e+02 $\pm$ 1.4e+02 3.8e+02 $\pm$ 43.883 3.2e+02 $\pm$ 31.593
kin8nm 7.2e-03 $\pm$ 1.7e-04 7.6e-03 $\pm$ 1.1e-04 7.7e-03 $\pm$ 2.7e-04 7.2e-03 $\pm$ 1.7e-04
miami_housing 3.9e+03 $\pm$ 2.1e+02 4.2e+03 $\pm$ 1.7e+02 5.2e+03 $\pm$ 3.4e+02 3.9e+03 $\pm$ 1.9e+02
naval_propulsion 5.5e-05 $\pm$ 1.9e-06 1.8e-04 $\pm$ 9.1e-06 6.4e-05 $\pm$ 2.7e-06 5.4e-05 $\pm$ 1.9e-06
parkinsons 0.209 $\pm$ 0.008 0.237 $\pm$ 0.008 0.228 $\pm$ 0.013 0.209 $\pm$ 0.010
powerplant 0.213 $\pm$ 0.011 0.222 $\pm$ 0.011 0.220 $\pm$ 0.011 0.212 $\pm$ 0.012
qsar 0.050 $\pm$ 0.003 0.053 $\pm$ 0.003 0.052 $\pm$ 0.003 0.049 $\pm$ 0.003
sulfur 2.0e-03 $\pm$ 1.1e-04 2.3e-03 $\pm$ 8.5e-05 2.3e-03 $\pm$ 1.5e-04 1.9e-03 $\pm$ 9.6e-05
superconductor 0.523 $\pm$ 0.018 0.611 $\pm$ 0.026 0.573 $\pm$ 0.024 0.523 $\pm$ 0.018
Table 39: Conformalized Variant (b) NCIW at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 0.195 $\pm$ 0.018 0.217 $\pm$ 0.021 0.196 $\pm$ 0.017 0.195 $\pm$ 0.017
airfoil 0.203 $\pm$ 0.007 0.211 $\pm$ 0.007 0.216 $\pm$ 0.011 0.203 $\pm$ 0.006
allstate 0.248 $\pm$ 0.041 0.270 $\pm$ 0.052 0.266 $\pm$ 0.044 0.250 $\pm$ 0.043
ca_housing 0.336 $\pm$ 0.006 0.354 $\pm$ 0.012 0.348 $\pm$ 0.011 0.336 $\pm$ 0.008
computer 0.098 $\pm$ 0.005 0.142 $\pm$ 0.008 0.103 $\pm$ 0.003 0.098 $\pm$ 0.005
concrete 0.248 $\pm$ 0.032 0.241 $\pm$ 0.025 0.261 $\pm$ 0.039 0.246 $\pm$ 0.033
elevator 0.137 $\pm$ 0.008 0.192 $\pm$ 0.012 0.147 $\pm$ 0.009 0.137 $\pm$ 0.008
energy_efficiency 0.046 $\pm$ 0.006 0.057 $\pm$ 0.004 0.050 $\pm$ 0.005 0.046 $\pm$ 0.005
insurance 0.314 $\pm$ 0.052 0.346 $\pm$ 0.076 0.335 $\pm$ 0.059 0.320 $\pm$ 0.059
kin8nm 0.355 $\pm$ 0.008 0.371 $\pm$ 0.009 0.371 $\pm$ 0.009 0.354 $\pm$ 0.008
miami_housing 0.084 $\pm$ 0.004 0.098 $\pm$ 0.002 0.085 $\pm$ 0.003 0.084 $\pm$ 0.003
naval_propulsion 1.8e-03 $\pm$ 4.7e-05 3.2e-03 $\pm$ 1.2e-04 1.9e-03 $\pm$ 8.4e-05 1.8e-03 $\pm$ 5.6e-05
parkinsons 0.284 $\pm$ 0.011 0.324 $\pm$ 0.008 0.309 $\pm$ 0.018 0.284 $\pm$ 0.010
powerplant 0.173 $\pm$ 0.008 0.182 $\pm$ 0.007 0.182 $\pm$ 0.007 0.173 $\pm$ 0.007
qsar 0.363 $\pm$ 0.122 0.393 $\pm$ 0.131 0.390 $\pm$ 0.131 0.363 $\pm$ 0.121
sulfur 0.108 $\pm$ 0.008 0.125 $\pm$ 0.009 0.105 $\pm$ 0.006 0.106 $\pm$ 0.007
superconductor 0.218 $\pm$ 0.023 0.241 $\pm$ 0.028 0.226 $\pm$ 0.022 0.218 $\pm$ 0.022
Table 40: Conformalized Variant (b) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR
ailerons 7.4e-04 $\pm$ 1.8e-05 8.1e-04 $\pm$ 1.7e-05 7.8e-04 $\pm$ 2.2e-05 7.4e-04 $\pm$ 1.8e-05
airfoil 8.794 $\pm$ 1.047 9.523 $\pm$ 0.959 10.075 $\pm$ 1.227 8.717 $\pm$ 0.870
allstate 9.2e+03 $\pm$ 6.1e+02 1.0e+04 $\pm$ 6.6e+02 1.0e+04 $\pm$ 6.2e+02 9.2e+03 $\pm$ 5.9e+02
ca_housing 2.2e+05 $\pm$ 4.6e+03 2.6e+05 $\pm$ 7.3e+03 2.4e+05 $\pm$ 5.9e+03 2.2e+05 $\pm$ 4.8e+03
computer 12.704 $\pm$ 0.571 16.543 $\pm$ 0.706 16.505 $\pm$ 1.190 12.707 $\pm$ 0.575
concrete 27.168 $\pm$ 2.969 26.195 $\pm$ 3.142 28.509 $\pm$ 3.874 27.020 $\pm$ 3.202
elevator 0.011 $\pm$ 0.000 0.015 $\pm$ 0.001 0.013 $\pm$ 0.000 0.011 $\pm$ 0.000
energy_efficiency 2.609 $\pm$ 0.417 3.019 $\pm$ 0.237 2.681 $\pm$ 0.431 2.465 $\pm$ 0.347
insurance 2.6e+04 $\pm$ 2.3e+03 4.2e+04 $\pm$ 1.1e+04 3.1e+04 $\pm$ 3.5e+03 2.6e+04 $\pm$ 2.5e+03
kin8nm 0.578 $\pm$ 0.014 0.607 $\pm$ 0.009 0.615 $\pm$ 0.022 0.577 $\pm$ 0.014
miami_housing 3.2e+05 $\pm$ 1.7e+04 3.3e+05 $\pm$ 1.4e+04 4.2e+05 $\pm$ 2.7e+04 3.1e+05 $\pm$ 1.5e+04
naval_propulsion 4.4e-03 $\pm$ 1.5e-04 0.014 $\pm$ 0.001 5.1e-03 $\pm$ 2.2e-04 4.4e-03 $\pm$ 1.5e-04
parkinsons 16.712 $\pm$ 0.669 18.967 $\pm$ 0.645 18.250 $\pm$ 1.030 16.697 $\pm$ 0.794
powerplant 17.066 $\pm$ 0.919 17.724 $\pm$ 0.919 17.567 $\pm$ 0.899 16.993 $\pm$ 0.937
qsar 3.960 $\pm$ 0.207 4.277 $\pm$ 0.215 4.195 $\pm$ 0.229 3.956 $\pm$ 0.218
sulfur 0.157 $\pm$ 0.008 0.180 $\pm$ 0.007 0.186 $\pm$ 0.012 0.155 $\pm$ 0.008
superconductor 41.818 $\pm$ 1.464 48.902 $\pm$ 2.063 45.854 $\pm$ 1.910 41.801 $\pm$ 1.407
G.3 Variant (c) Conformalized
Table 41: Conformalized Variant (c) PICP at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR UACQR-P UACQR-S
ailerons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.00
airfoil 0.96 $\pm$ 0.02 0.96 $\pm$ 0.02 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01
allstate 0.95 $\pm$ 0.01 0.94 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
ca_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.96 $\pm$ 0.00
computer 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.97 $\pm$ 0.00
concrete 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.96 $\pm$ 0.01 0.96 $\pm$ 0.02 0.96 $\pm$ 0.02
elevator 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.97 $\pm$ 0.00
energy_efficiency 0.97 $\pm$ 0.01 0.95 $\pm$ 0.03 0.96 $\pm$ 0.01 0.95 $\pm$ 0.01 0.97 $\pm$ 0.02 0.96 $\pm$ 0.02
insurance 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.96 $\pm$ 0.02 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
kin8nm 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
miami_housing 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
naval_propulsion 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.99 $\pm$ 0.00
parkinsons 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.02 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.97 $\pm$ 0.01
powerplant 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01
qsar 0.95 $\pm$ 0.01 0.94 $\pm$ 0.01 0.94 $\pm$ 0.01 0.95 $\pm$ 0.01 0.96 $\pm$ 0.01 0.96 $\pm$ 0.01
sulfur 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01
superconductor 0.95 $\pm$ 0.01 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.00 0.95 $\pm$ 0.01
Table 42: Conformalized Variant (c) NIW at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR UACQR-P UACQR-S
ailerons 0.207 $\pm$ 0.017 0.239 $\pm$ 0.017 0.202 $\pm$ 0.019 0.208 $\pm$ 0.017 0.177 $\pm$ 0.014 0.207 $\pm$ 0.017
airfoil 0.210 $\pm$ 0.021 0.212 $\pm$ 0.020 0.230 $\pm$ 0.040 0.199 $\pm$ 0.013 0.365 $\pm$ 0.022 0.391 $\pm$ 0.029
allstate 0.282 $\pm$ 0.052 0.292 $\pm$ 0.061 0.279 $\pm$ 0.037 0.279 $\pm$ 0.052 0.285 $\pm$ 0.051 0.305 $\pm$ 0.055
ca_housing 0.317 $\pm$ 0.009 0.330 $\pm$ 0.013 0.344 $\pm$ 0.010 0.315 $\pm$ 0.010 0.366 $\pm$ 0.007 0.405 $\pm$ 0.007
computer 0.079 $\pm$ 0.003 0.083 $\pm$ 0.003 0.088 $\pm$ 0.004 0.080 $\pm$ 0.004 0.084 $\pm$ 0.003 0.101 $\pm$ 0.001
concrete 0.290 $\pm$ 0.037 0.273 $\pm$ 0.032 0.287 $\pm$ 0.037 0.278 $\pm$ 0.029 0.373 $\pm$ 0.025 0.403 $\pm$ 0.022
elevator 0.126 $\pm$ 0.008 0.153 $\pm$ 0.008 0.136 $\pm$ 0.008 0.127 $\pm$ 0.008 0.158 $\pm$ 0.011 0.200 $\pm$ 0.010
energy_efficiency 0.052 $\pm$ 0.007 0.043 $\pm$ 0.008 0.053 $\pm$ 0.007 0.043 $\pm$ 0.007 inf $\pm$ nan 0.162 $\pm$ 0.014
insurance 0.414 $\pm$ 0.094 0.376 $\pm$ 0.134 0.434 $\pm$ 0.086 0.397 $\pm$ 0.091 0.305 $\pm$ 0.046 0.311 $\pm$ 0.033
kin8nm 0.328 $\pm$ 0.015 0.327 $\pm$ 0.012 0.327 $\pm$ 0.015 0.325 $\pm$ 0.012 0.446 $\pm$ 0.018 0.460 $\pm$ 0.019
miami_housing 0.088 $\pm$ 0.003 0.096 $\pm$ 0.005 0.112 $\pm$ 0.019 0.088 $\pm$ 0.001 0.104 $\pm$ 0.003 0.113 $\pm$ 0.002
naval_propulsion 1.6e-03 $\pm$ 3.6e-05 1.7e-03 $\pm$ 6.5e-05 1.6e-03 $\pm$ 4.9e-05 1.6e-03 $\pm$ 3.2e-05 1.7e-03 $\pm$ 8.2e-05 3.6e-03 $\pm$ 8.5e-05
parkinsons 0.250 $\pm$ 0.015 0.262 $\pm$ 0.009 0.281 $\pm$ 0.015 0.250 $\pm$ 0.007 0.269 $\pm$ 0.020 0.313 $\pm$ 0.018
powerplant 0.158 $\pm$ 0.011 0.170 $\pm$ 0.007 0.162 $\pm$ 0.008 0.157 $\pm$ 0.007 0.193 $\pm$ 0.007 0.202 $\pm$ 0.009
qsar 0.354 $\pm$ 0.120 0.402 $\pm$ 0.143 0.349 $\pm$ 0.115 0.344 $\pm$ 0.117 0.410 $\pm$ 0.135 0.452 $\pm$ 0.152
sulfur 0.112 $\pm$ 0.011 0.126 $\pm$ 0.012 0.107 $\pm$ 0.016 0.111 $\pm$ 0.009 0.115 $\pm$ 0.012 0.126 $\pm$ 0.012
superconductor 0.196 $\pm$ 0.022 0.234 $\pm$ 0.026 0.208 $\pm$ 0.025 0.195 $\pm$ 0.020 0.220 $\pm$ 0.022 0.232 $\pm$ 0.026
Table 43: Conformalized Variant (c) Quantile Loss at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR UACQR-P UACQR-S
ailerons 9.6e-06 $\pm$ 2.9e-07 1.1e-05 $\pm$ 2.9e-07 9.8e-06 $\pm$ 3.2e-07 9.6e-06 $\pm$ 2.9e-07 1.0e-05 $\pm$ 0.0e+00 1.0e-05 $\pm$ 0.0e+00
airfoil 0.110 $\pm$ 0.015 0.111 $\pm$ 0.013 0.131 $\pm$ 0.019 0.107 $\pm$ 0.014 0.188 $\pm$ 0.013 0.200 $\pm$ 0.010
allstate 1.2e+02 $\pm$ 6.810 1.4e+02 $\pm$ 11.278 1.3e+02 $\pm$ 22.506 1.2e+02 $\pm$ 6.874 1.1e+02 $\pm$ 7.356 1.2e+02 $\pm$ 6.507
ca_housing 2.8e+03 $\pm$ 91.022 2.9e+03 $\pm$ 88.973 3.2e+03 $\pm$ 1.1e+02 2.8e+03 $\pm$ 92.560 3.0e+03 $\pm$ 58.564 3.0e+03 $\pm$ 57.081
computer 0.138 $\pm$ 0.012 0.140 $\pm$ 0.012 0.163 $\pm$ 0.013 0.137 $\pm$ 0.011 0.147 $\pm$ 0.006 0.151 $\pm$ 0.004
concrete 0.343 $\pm$ 0.033 0.334 $\pm$ 0.031 0.381 $\pm$ 0.058 0.337 $\pm$ 0.032 0.413 $\pm$ 0.037 0.425 $\pm$ 0.022
elevator 1.3e-04 $\pm$ 3.3e-06 1.4e-04 $\pm$ 3.9e-06 1.5e-04 $\pm$ 3.5e-06 1.3e-04 $\pm$ 3.2e-06 1.8e-04 $\pm$ 5.4e-06 1.9e-04 $\pm$ 4.9e-06
energy_efficiency 0.031 $\pm$ 0.006 0.030 $\pm$ 0.007 0.035 $\pm$ 0.010 0.029 $\pm$ 0.007 inf $\pm$ nan 0.079 $\pm$ 0.006
insurance 4.1e+02 $\pm$ 73.015 4.3e+02 $\pm$ 61.342 3.9e+02 $\pm$ 53.648 4.2e+02 $\pm$ 80.269 3.4e+02 $\pm$ 46.073 3.3e+02 $\pm$ 34.606
kin8nm 6.7e-03 $\pm$ 1.7e-04 6.9e-03 $\pm$ 2.2e-04 6.8e-03 $\pm$ 2.0e-04 6.7e-03 $\pm$ 1.8e-04 8.8e-03 $\pm$ 1.6e-04 9.0e-03 $\pm$ 1.6e-04
miami_housing 3.8e+03 $\pm$ 1.9e+02 3.8e+03 $\pm$ 2.2e+02 6.8e+03 $\pm$ 1.2e+03 3.7e+03 $\pm$ 1.4e+02 4.3e+03 $\pm$ 1.3e+02 4.5e+03 $\pm$ 1.3e+02
naval_propulsion 4.0e-05 $\pm$ 1.0e-06 4.2e-05 $\pm$ 1.3e-06 4.9e-05 $\pm$ 1.6e-06 4.0e-05 $\pm$ 1.0e-06 9.3e-05 $\pm$ 9.0e-06 8.1e-05 $\pm$ 3.0e-06
parkinsons 0.190 $\pm$ 0.007 0.192 $\pm$ 0.006 0.217 $\pm$ 0.008 0.189 $\pm$ 0.006 0.180 $\pm$ 0.009 0.200 $\pm$ 0.010
powerplant 0.204 $\pm$ 0.012 0.210 $\pm$ 0.013 0.209 $\pm$ 0.013 0.203 $\pm$ 0.012 0.226 $\pm$ 0.011 0.234 $\pm$ 0.010
qsar 0.049 $\pm$ 0.002 0.054 $\pm$ 0.001 0.051 $\pm$ 0.003 0.048 $\pm$ 0.002 0.051 $\pm$ 0.003 0.054 $\pm$ 0.002
sulfur 1.8e-03 $\pm$ 8.0e-05 1.9e-03 $\pm$ 8.7e-05 2.6e-03 $\pm$ 2.6e-04 1.8e-03 $\pm$ 6.3e-05 1.9e-03 $\pm$ 1.2e-04 2.0e-03 $\pm$ 1.5e-04
superconductor 0.488 $\pm$ 0.020 0.573 $\pm$ 0.026 0.563 $\pm$ 0.047 0.489 $\pm$ 0.020 0.507 $\pm$ 0.017 0.527 $\pm$ 0.016
Table 44: Conformalized Variant (c) NCIW at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR UACQR-P UACQR-S
ailerons 0.209 $\pm$ 0.021 0.242 $\pm$ 0.021 0.203 $\pm$ 0.021 0.210 $\pm$ 0.020 0.188 $\pm$ 0.030 0.207 $\pm$ 0.017
airfoil 0.189 $\pm$ 0.016 0.199 $\pm$ 0.013 0.204 $\pm$ 0.017 0.188 $\pm$ 0.014 0.317 $\pm$ 0.025 0.329 $\pm$ 0.033
allstate 0.274 $\pm$ 0.046 0.296 $\pm$ 0.058 0.285 $\pm$ 0.029 0.277 $\pm$ 0.056 0.279 $\pm$ 0.047 0.296 $\pm$ 0.049
ca_housing 0.312 $\pm$ 0.010 0.329 $\pm$ 0.009 0.342 $\pm$ 0.011 0.311 $\pm$ 0.010 0.368 $\pm$ 0.007 0.405 $\pm$ 0.007
computer 0.080 $\pm$ 0.003 0.083 $\pm$ 0.003 0.090 $\pm$ 0.004 0.081 $\pm$ 0.004 0.087 $\pm$ 0.005 0.101 $\pm$ 0.001
concrete 0.255 $\pm$ 0.032 0.257 $\pm$ 0.031 0.263 $\pm$ 0.028 0.259 $\pm$ 0.032 0.332 $\pm$ 0.056 0.344 $\pm$ 0.054
elevator 0.126 $\pm$ 0.007 0.151 $\pm$ 0.011 0.138 $\pm$ 0.008 0.126 $\pm$ 0.008 0.165 $\pm$ 0.015 0.200 $\pm$ 0.010
energy_efficiency 0.043 $\pm$ 0.005 0.043 $\pm$ 0.005 0.048 $\pm$ 0.005 0.041 $\pm$ 0.006 0.140 $\pm$ 0.014 0.138 $\pm$ 0.024
insurance 0.356 $\pm$ 0.071 0.358 $\pm$ 0.078 0.400 $\pm$ 0.091 0.345 $\pm$ 0.078 0.291 $\pm$ 0.043 0.301 $\pm$ 0.034
kin8nm 0.323 $\pm$ 0.008 0.331 $\pm$ 0.010 0.327 $\pm$ 0.008 0.323 $\pm$ 0.008 0.433 $\pm$ 0.009 0.445 $\pm$ 0.013
miami_housing 0.087 $\pm$ 0.003 0.096 $\pm$ 0.004 0.107 $\pm$ 0.013 0.088 $\pm$ 0.002 0.102 $\pm$ 0.004 0.110 $\pm$ 0.004
naval_propulsion 1.6e-03 $\pm$ 5.0e-05 1.7e-03 $\pm$ 4.7e-05 1.6e-03 $\pm$ 4.1e-05 1.6e-03 $\pm$ 5.0e-05 1.7e-03 $\pm$ 8.2e-05 3.6e-03 $\pm$ 8.5e-05
parkinsons 0.250 $\pm$ 0.010 0.259 $\pm$ 0.009 0.279 $\pm$ 0.013 0.250 $\pm$ 0.012 0.273 $\pm$ 0.019 0.313 $\pm$ 0.018
powerplant 0.162 $\pm$ 0.007 0.171 $\pm$ 0.008 0.163 $\pm$ 0.008 0.161 $\pm$ 0.007 0.191 $\pm$ 0.009 0.200 $\pm$ 0.009
qsar 0.350 $\pm$ 0.121 0.412 $\pm$ 0.144 0.360 $\pm$ 0.120 0.348 $\pm$ 0.118 0.409 $\pm$ 0.134 0.449 $\pm$ 0.150
sulfur 0.111 $\pm$ 0.010 0.124 $\pm$ 0.008 0.107 $\pm$ 0.005 0.109 $\pm$ 0.008 0.114 $\pm$ 0.010 0.126 $\pm$ 0.010
superconductor 0.194 $\pm$ 0.021 0.234 $\pm$ 0.025 0.204 $\pm$ 0.024 0.192 $\pm$ 0.019 0.220 $\pm$ 0.023 0.232 $\pm$ 0.025
Table 45: Conformalized Variant (c) Interval Score Loss at 95% prediction intervals, aggregated across 10 seeds. Methods with suffix ‘-c’ denote conformalized variants obtained using the validation set divided into two parts, one for validation and one for calibration.Values $\geq 100$ or <0.01<0.01 are presented in scientific notation with 1 decimal place. Bold values (desirable) are the minimum for that dataset and metric, while the underlined values indicate the second-best result. Red values are more than 33% worse than the best result.
Dataset CLEAR-c PCS-EPISTEMIC-c ALEATORIC-R-c CLEAR UACQR-P UACQR-S
ailerons 7.7e-04 $\pm$ 2.3e-05 8.6e-04 $\pm$ 2.3e-05 7.9e-04 $\pm$ 2.6e-05 7.7e-04 $\pm$ 2.3e-05 7.9e-04 $\pm$ 3.1e-05 7.9e-04 $\pm$ 2.3e-05
airfoil 8.780 $\pm$ 1.202 8.857 $\pm$ 1.035 10.504 $\pm$ 1.503 8.569 $\pm$ 1.099 15.052 $\pm$ 1.025 15.999 $\pm$ 0.776
allstate 9.4e+03 $\pm$ 5.4e+02 1.1e+04 $\pm$ 9.0e+02 1.0e+04 $\pm$ 1.8e+03 9.2e+03 $\pm$ 5.5e+02 9.0e+03 $\pm$ 5.9e+02 9.3e+03 $\pm$ 5.2e+02
ca_housing 2.2e+05 $\pm$ 7.3e+03 2.3e+05 $\pm$ 7.1e+03 2.5e+05 $\pm$ 8.6e+03 2.2e+05 $\pm$ 7.4e+03 2.4e+05 $\pm$ 4.7e+03 2.4e+05 $\pm$ 4.6e+03
computer 11.040 $\pm$ 0.933 11.164 $\pm$ 0.938 13.042 $\pm$ 1.009 10.978 $\pm$ 0.911 11.757 $\pm$ 0.463 12.103 $\pm$ 0.323
concrete 27.403 $\pm$ 2.641 26.754 $\pm$ 2.444 30.441 $\pm$ 4.624 26.928 $\pm$ 2.529 33.077 $\pm$ 2.941 33.990 $\pm$ 1.797
elevator 0.011 $\pm$ 0.000 0.012 $\pm$ 0.000 0.012 $\pm$ 0.000 0.010 $\pm$ 0.000 0.014 $\pm$ 0.000 0.015 $\pm$ 0.000
energy_efficiency 2.463 $\pm$ 0.488 2.422 $\pm$ 0.596 2.796 $\pm$ 0.804 2.348 $\pm$ 0.554 6.214 $\pm$ 0.408 6.354 $\pm$ 0.480
insurance 3.3e+04 $\pm$ 5.8e+03 3.4e+04 $\pm$ 4.9e+03 3.1e+04 $\pm$ 4.3e+03 3.4e+04 $\pm$ 6.4e+03 2.7e+04 $\pm$ 3.7e+03 2.7e+04 $\pm$ 2.8e+03
kin8nm 0.536 $\pm$ 0.014 0.552 $\pm$ 0.018 0.545 $\pm$ 0.016 0.534 $\pm$ 0.014 0.705 $\pm$ 0.013 0.723 $\pm$ 0.013
miami_housing 3.0e+05 $\pm$ 1.5e+04 3.1e+05 $\pm$ 1.7e+04 5.5e+05 $\pm$ 9.6e+04 2.9e+05 $\pm$ 1.1e+04 3.5e+05 $\pm$ 1.0e+04 3.6e+05 $\pm$ 1.0e+04
naval_propulsion 3.2e-03 $\pm$ 8.0e-05 3.4e-03 $\pm$ 1.0e-04 3.9e-03 $\pm$ 1.3e-04 3.2e-03 $\pm$ 8.4e-05 7.5e-03 $\pm$ 7.5e-04 6.6e-03 $\pm$ 1.3e-04
parkinsons 15.231 $\pm$ 0.540 15.327 $\pm$ 0.476 17.375 $\pm$ 0.656 15.149 $\pm$ 0.518 14.384 $\pm$ 0.700 15.978 $\pm$ 0.779
powerplant 16.287 $\pm$ 0.959 16.839 $\pm$ 1.025 16.687 $\pm$ 1.010 16.243 $\pm$ 0.985 18.078 $\pm$ 0.906 18.725 $\pm$ 0.822
qsar 3.903 $\pm$ 0.190 4.360 $\pm$ 0.105 4.094 $\pm$ 0.213 3.877 $\pm$ 0.176 4.102 $\pm$ 0.230 4.333 $\pm$ 0.182
sulfur 0.145 $\pm$ 0.006 0.148 $\pm$ 0.007 0.209 $\pm$ 0.021 0.143 $\pm$ 0.005 0.153 $\pm$ 0.010 0.158 $\pm$ 0.012
superconductor 39.077 $\pm$ 1.615 45.819 $\pm$ 2.070 45.024 $\pm$ 3.738 39.104 $\pm$ 1.604 40.545 $\pm$ 1.386 42.122 $\pm$ 1.257
Table 46: Conformalized Variant (c) CLEAR calibration parameters $\lambda$ and $\gamma_{1}$ for 95% prediction intervals across 10 seeds. Using all available variables. Showing median [min:max] values.
Dataset $\lambda$ $\gamma_{1}$
ailerons 1.01 [0.26:1.75] 1.01 [0.87:1.32]
airfoil 1.04 [0.10:100.00] 1.45 [0.02:6.72]
allstate 0.72 [0.00:100.00] 1.10 [0.02:1.43]
ca_housing 1.93 [1.11:13.99] 0.67 [0.14:0.95]
computer 2.85 [1.39:10.47] 0.60 [0.18:0.95]
concrete 1.18 [0.25:100.00] 1.79 [0.02:5.83]
elevator 0.73 [0.45:1.29] 1.13 [0.86:1.43]
energy_efficiency 0.22 [0.02:100.00] 6.87 [0.02:19.92]
insurance 4.77 [0.37:100.00] 0.55 [0.03:5.83]
kin8nm 2.21 [0.39:3.27] 0.74 [0.61:1.02]
miami_housing 4.61 [1.23:100.00] 0.36 [0.02:1.09]
naval_propulsion 14.44 [10.01:22.37] 0.69 [0.52:0.85]
parkinsons 1.11 [0.53:15.09] 1.34 [0.12:2.31]
powerplant 0.96 [0.57:3.30] 1.12 [0.51:1.41]
qsar 0.59 [0.22:2.58] 1.45 [0.78:2.51]
sulfur 2.98 [0.90:100.00] 0.63 [0.02:1.25]
superconductor 0.73 [0.22:1.38] 1.17 [0.85:1.33]
Appendix H Case Study: House Price Prediction with Varying Number of Predictors
We illustrate our method using the Ames Housing dataset, which contains data on 2,930 residential properties sold in Ames, Iowa, between 2006 and 2010. The target variable is the sale price of a house, and the full dataset includes around 80 predictor variables describing various aspects such as square footage, neighborhood, and building type. This dataset, originally collected by the Ames City Assessor’s Office and curated by [113], was explored in detail in Chapter 13 of [40].
The PCS framework [40] involves quantifying all sources of extended epistemic uncertainty stemming from the entire data-science cycle, such as uncertainty stemming from data processing. An pipeline of PCS uncertainty quantification applied to the Ames Housing data can be found in Chapter 13 in [40]. We follow the same steps [^8] Minor differences arise due to: (a) implementation differences in base models between R and Python, and (b) manual calibration of PCS intervals, which was not part of the original implementation. Additionally, both CQR and CLEAR were trained using only linear quantile regressors as the model selection step of PCS selected a linear model based on the RMSE on the validation dataset. for estimating $\hat{f}$ and estimating epistemic uncertainty bounds $\hat{q}_{0.05}^{\text{epi}},\hat{q}_{0.95}^{\text{epi}}$.
To investigate how the amount of available information affects predictive uncertainty, we consider two versions of the dataset:
  • Full: the original dataset with all $\sim$80 variables,
  • Top 2: a further reduced dataset containing only the top two features (as determined by either feature importance of a random forest, or by correlation with the outcome variable, as both approaches lead to the same choice).
These settings simulate scenarios where fewer variables are available, reflecting different levels of information accessibility. Reducing the number of predictors is expected to increase aleatoric uncertainty (due to missing key predictive information) and decrease epistemic uncertainty (due to a simpler model class and lower dimensionality).
In each case, we applied the CLEAR procedure as described in Section 2.4. This involved: (1) data cleaning and preprocessing (excluding irregular sales, imputing missing values, encoding categorical variables), resulting in $N_{1}=438$ cleaned datasets; (2) fitting a predictive model $\hat{f}$ and estimating epistemic uncertainty bounds $\hat{q}_{0.05}^{\text{epi}},\hat{q}_{0.95}^{\text{epi}}$; (3) estimating aleatoric uncertainty bounds $\hat{q}_{0.05}^{\text{ale}},\hat{q}_{0.95}^{\text{ale}}$ via quantile regression, as described in  Section 2.3; (4) estimating $\lambda$ and calibrating prediction intervals on the validation set.
Table 2 and the discussion in Section 4.3 summarize the results. In the reduced 2-variable setting, CLEAR selected $\lambda=0.6$, assigning more weight to aleatoric uncertainty. In contrast, in the full-feature setting, $\lambda=14.5$, emphasizing epistemic uncertainty. The corresponding calibrated epistemic-to-aleatoric ratios were computed as $\frac{1}{n}\sum_{i=1}^{n}\frac{\gamma_{1}\times\text{epistemic}(x_{i})}{\gamma_{2}\times\text{aleatoric}(x_{i})}$ in (1), reflecting the final width contributions from each component. They were 0.03 and 7.72, respectively. This ratio quantifies how much of the interval width is attributed to epistemic versus aleatoric uncertainty after calibration. The results demonstrate CLEAR’s ability to adaptively re-weight the two uncertainty sources in response to the underlying information regime, producing intervals that are both sharp and well-calibrated across heterogeneous settings.
Appendix I On the Role of Relative and Absolute Uncertainty in Coverage Guarantees
In predictive inference, a fundamental distinction arises between conditional and marginal coverage—closely related to what has been termed “relative vs. absolute uncertainty” [27] or “adaptive vs. calibrated uncertainty.” Conditional coverage requires that prediction intervals achieve the target coverage level at every individual input, thereby capturing local or relative uncertainty. Formally, a prediction interval $C(X_{n+1})$ satisfies conditional coverage at level $1-\alpha$ if
$\forall x\in\text{supp}(X):\quad\mathbb{P}\left[Y_{n+1}\in C(X_{n+1})\mid X_{n+1}=x\right]\geq 1-\alpha.$
By contrast, marginal coverage only guarantees coverage on average over the distribution of inputs:
$\mathbb{P}\left[Y_{n+1}\in C(X_{n+1})\right]\geq 1-\alpha.$
While conditional coverage implies marginal coverage, the reverse does not hold. This distinction is especially important in settings with heteroskedasticity, where the variability of $Y\mid X$ changes across the input space, and under distribution shift, where the test distribution of $X$ differs from the training distribution. Distribution shift—such as covariate shift or domain adaptation—can render marginal guarantees unreliable since they depend on the marginal $\mathbb{P}_{X}$. In contrast, conditional coverage ensures that prediction intervals remain valid even when $\mathbb{P}_{X}$ changes, provided the conditional distribution $\mathbb{P}(Y\mid X)$ remains stable. In what follows, we explore the implications of these distinctions and how they shape both the evaluation and design of uncertainty quantification methods.
I.1 Metrics: Relevance
To assess the quality of predictive uncertainty, various metrics capture different aspects of coverage and adaptivity:
  • Quantile Loss, CRPS, and AISL combine both relative and absolute components incentivizing conditional coverage. These metrics penalize both poor ranking and miscalibration, rewarding methods that adapt well to heteroskedasticity.
  • NCIW is invariant to the overall scale of the uncertainty but evaluates the ranking—whether a method assigns wider intervals to more uncertain points. In practice, a low NCIW often correlates with good relative uncertainty. However, a minimal NCIW theoretically encourages suboptimal conditional coverage, as it can lead to under-coverage in high-uncertainty regions and over-coverage in low-uncertainty regions. Consequently, good conditional coverage typically requires a slightly higher NCIW than the minimum. Similar to NCIW, Minimum Negative Log-Likelihood (NLLmin) [27] also assesses relative uncertainty by focusing on whether a method correctly ranks more versus less uncertain inputs, independent of the predicted scale.
  • PICP and NIW each provide only partial information: PICP measures calibration but is blind to adaptivity, while NIW captures the average scale of uncertainty.
These metrics reflect different priorities in uncertainty quantification. Choosing among them (or combining them) depends on whether the primary goal is calibration, adaptivity, or both.
I.2 Applications
Understanding whether a method captures relative or absolute uncertainty has practical implications across a range of applications. In active learning, the primary objective is to identify inputs $x$ for which the model is most uncertain, guiding efficient data acquisition. Here, only the ranking of uncertainty matters—selecting the point with the highest epistemic uncertainty. Methods that preserve good relative uncertainty, even if miscalibrated, often suffice. In Bayesian optimization, many acquisition functions (such as upper confidence bound or entropy search) depend more on relative than absolute uncertainty [114, 115]. Using upper bounds of the form $\hat{f}(x)+c$ with constant $c$ across all $x$ does not improve over exploiting $\hat{f}(x)$ alone, highlighting the centrality of uncertainty ranking over calibration in this setting. In human-in-the-loop automation, relative uncertainty can guide prioritization—for instance, flagging uncertain cases for expert review. While calibrated intervals may not always be necessary, correct ordering of confidence can improve decision efficiency and safety.
I.3 Methods
Different uncertainty quantification methods prioritize and estimate relative and absolute uncertainty to varying degrees:
  • NOMU [27] is explicitly designed to estimate only relative uncertainty. It does not attempt to calibrate the absolute scale, making it suitable for applications where ranking matters but calibrated intervals are unnecessary.
  • Deep ensembles [22] typically yield strong performance in capturing relative uncertainty, particularly through diversity in predictions across ensemble members. However, they often suffer from miscalibration in the absolute scale of uncertainty unless explicitly corrected.
  • CLEAR separately estimates relative epistemic and aleatoric uncertainty and combines them using a tunable parameter $\lambda$, which also refines the ranking of uncertainty—offering an alternative to standard calibration techniques. The absolute scale is then calibrated using a second parameter, $\gamma_{1}$, allowing for flexible control over both adaptivity and calibration.
These methods highlight the spectrum of approaches to uncertainty quantification, from ranking-only models to fully calibrated systems.
I.4 Calibration Techniques
The absolute scale of uncertainty is critical in applications where the expected risk or cost over a population matters (e.g., in probabilistic risk management, climate forecasting, or decision-making under uncertainty). However, as emphasized by pathological cases that achieve perfect PICP with no adaptivity, relative uncertainty remains essential for practicality. Several approaches can be taken to adjust the scale of uncertainty estimates while preserving different structures:
  • Multiplicative calibration scales all predictions by a constant factor, preserving the multiplicative structure of uncertainty. This is appropriate when the model’s ranking is reliable.
  • Additive calibration shifts all intervals uniformly, preserving additive differences but potentially distorting proportional uncertainty.
  • Isotonic calibration applies a nonparametric monotonic transformation that preserves the ranking of uncertainty, suitable when only order is trusted. E.g., [3] uses isotonic calibration for distributional uncertainty quantification.
  • CLEAR calibrates aleatoric and epistemic uncertainty separately, allowing independent yet coherent control of both components. This facilitates calibrated estimates of total predictive uncertainty while preserving the relative structure.
I.5 Achieving Conditional Coverage
Achieving conditional coverage requires addressing both components of predictive uncertainty:
  1. 1.
    Estimating relative uncertainty accurately—capturing how uncertainty varies across inputs.
  2. 2.
    Calibrating the absolute scale to ensure the desired coverage level holds at each input.
Even when conditional coverage is not required, improving relative uncertainty tends to reduce average interval width and improve marginal calibration under distribution shifts. Thus, adaptivity and calibration are not mutually exclusive but can reinforce one another. CLEAR is explicitly designed to target both forms of uncertainty—modeling and calibrating relative and absolute epistemic, aleatoric, and total predictive uncertainty. As such, it provides a flexible and principled framework for applications demanding both adaptivity and reliability.